### v2

# supReferences

Abstract
[…]
Mathieu Gerber
University of Bristol
1 Introduction
1.1 Context
Particle filters (PFs) are used for real-time inference in state-space models (SSMs) in various applications, including robotics (Thrun,, 2002; Hailu et al.,, 2024), self-driving cars (Berntorp and Di Cairano,, 2019; Hafez et al.,, 2024), assisted surgery (Kummert et al.,, 2021) and ballistics objects tracking (Kim et al.,, 2023).
Denoting by $t\in\mathbb{N}$ the time index and by $\hat{\eta}_{t}$ the conditional distribution of the state variable at time $t$ given the available observations $\{y_{s}\}_{s=1}^{t}$, a PF is a Monte Carlo algorithm that propagates over time a set of $N\in\mathbb{N}$ particles that can be used to estimate, in an online fashion, the sequence of the so-called filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$. For instance, in self-driving cars applications $\hat{\eta}_{t}$ is the conditional distribution of the position of the car at time $t$ given its GPS localisation (known to have a precision of a few meters) and its distance with respect to some nearby landmarks, under the assumed SSM.
The popularity of PFs stems from the fact that this class of algorithms can be deployed on a very large class of SSMs. In addition, it is widely believed that the quality of the PF estimate $\hat{\eta}_{t}^{N}$ of $\hat{\eta}_{t}$ does not deteriorate over time, a belief supported by several theoretical analyses showing that, for some class of functions $\mathfrak{F}$ and $p\in\mathbb{N}$, we have (see Caffarel et al.,, 2025, and references therein)
$\lim_{N\rightarrow\infty}\sup_{t\geq 1}\mathbb{E}\big{[}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|^{p}\big{]}=0,\quad\forall f\in\mathfrak{F}.$ (1)
However, this type of results only provides a guarantee for the error that arises when estimating $\hat{\eta}_{t}$ by $\hat{\eta}_{t}^{N}$ at a single time instant $t$, while in practice we are usually interested in estimating accurately $\hat{\eta}_{t}$ for several time instants $t\in\{1,\dots,T\}$; for instance, a PF used for positioning in a self-driving car needs to provide a precise localization of the vehicle during its whole life time. In this context, instead of (1) a more relevant theoretical guarantee for PFs would be that, in some sense, $\sup_{t\geq 1}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\rightarrow 0$ as $N\rightarrow\infty$ and for all functions $f$ belonging to some class of functions.
1.2 Contributions of the paper
[…]
For a given $\kappa\in(0,1)$ the PF estimate $(\hat{\eta}_{t}^{N})_{t\geq 1}$ of $(\hat{\eta}_{t})_{n\geq 1}$ is therefore necessarily such that we have $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$. Given the ever increasing use of PFs in life critical applications, notably due to the increasing prevalence of autonomous systems (Ogunsina et al.,, 2024) and the role that PFs play in these systems, it is then crucial to understand the behaviour of $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ as a function of the number of particle $N$ and of the time horizon $T$. This is particularly important for applications which require to accurately estimate a very long sequence of filtering distributions $\{\hat{\eta}_{t}\}_{t=1}^{T}$; for instance, according to Kalra and Paddock, (2016), in order to demonstrate the reliability of self-driving cars their onboard algorithms must be tested for hundreds of billions of miles.
The second contribution of this paper is to provide a first theoretical analysis of the probability $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$. Considering the same filtering problem as the one described above, we prove that there exists a finite constant $C_{\kappa}$ such that for any $q\in(0,1)$ and time horizon $T\geq 1$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq C_{\kappa}(\log(T)+\log(1/q))$. Consequently, to guarantee that we still have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ as $T$ increases, the number of particles only needs to grow at speed $\log(T)$ while, to decrease $q$, the number of particles should only grow at speed $\log(1/q)$. To the best of our knowledge Del Moral, (2013) provides the only previous attempt to study the ability of PFs to jointly estimate multiple filtering distributions. In particular, in Del Moral, (2013, Corollary 14.3.3, page 435) it is shown that, for some finite constant $C_{\kappa}$, for all $q\in(0,1)$ and $T\in\mathbb{N}$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\geq\kappa)\leq q$ for all $N\geq C_{\kappa}\log(T)\log(1/q)$ and for any function $f$ with oscillations bounded by one. This result is however obtained under strong assumptions on the SSM (which are not verified for the filtering problem that we consider in this work) and it only provides a theoretical guarantee for estimating the filtering expectations $\{\hat{\eta}_{t}(f)\}_{t=1}^{T}$ of a single function.
Despite these reassuring results concerning the time evolution of PF estimates, it is not entirely satisfactory to delegate (life-critical) filtering tasks to algorithms that we know will fail infinitely often. Following a previous discussion, to resolve this limitation of PFs some de-randomization is needed, and a possible approach to do so is to replace its underpinning independent $\mathcal{U}(0,1)$ random variables by randomized quasi-Monte (RQMC) point sets. This idea has been proposed and studied in Gerber and Chopin, (2015) with the aim of obtaining a filtering algorithm, called sequential quasi-Monte Carlo (SQMC), which has a faster convergence rate (as $N\rightarrow\infty$) than PFs. Letting $(\tilde{\eta}^{N}_{t})_{t\geq 1}$ denote the estimate of $(\hat{\eta}_{t})_{t\geq 1}$ obtained with SQMC, the third contribution of this paper is to show, for the one-dimensional filtering problem outlined above, $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\tilde{\eta}^{N}_{t}-\hat{\eta}_{t}\|=0$ with probability one.
To analyse the time evolution of PF and SQMC estimates we introduce a new approach for studying PF algorithms which builds on tools developed in the PF and in the quasi-Monte Carlo literature. Informally speaking, in a first step we analyse the behaviour of the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ for a given realization of the $\mathcal{U}(0,1)$ random numbers underpinning the algorithm. Then, in a second step, we leverage the known properties of sets of independent $\mathcal{U}(0,1)$ random numbers and of RQMC point sets to apply the result obtained in the first step to PFs and SQMC.
1.3 Notation and organization of the paper
In what follows all the random variables are defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. To introduce some notation let $s\in\mathbb{N}$. Then, we let $\mathcal{B}(\mathbb{R}^{s})$ denote the Borel $\sigma$-algebra on $\mathbb{R}^{s}$, $\mathcal{P}(\mathbb{R}^{s})$ denote the set of probability measures on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$, $\lambda_{s}$ denote the Lebesgue measure on $\mathbb{R}^{s}$ and
$\mathcal{B}_{s}=\Big{\{}B\subset\mathbb{R}^{s}:B=\prod_{i=1}^{s}(-\infty,a_{i}],\,(a_{1},\dots a_{s})\in\mathbb{R}^{s}\Big{\}}.$
In addition, for any signed-measure $\mu$ on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$ we let $\|\mu\|=\sup_{B\in\mathcal{B}_{s}}|\mu(B)|$ and, for a point set $\{u^{n}\}_{n=1}^{N}$ in $[0,1)^{s}$, we denote by $D^{*}_{s}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}$ its star discrepancy, that is
$D^{*}_{s}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}:=\sup_{b\in[0,1)^{s}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{[0,b)}(u^{n})-\lambda_{s}\big{(}[0,b)\big{)}\Big{|}.$
Next, for a probability distribution $\pi\in\mathcal{P}(\mathbb{R}^{s})$, a measurable function $f:\mathbb{R}^{s}\rightarrow\mathbb{R}$ and a (potentially un-normalized) kernel $K$ acting on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$, we let $\pi(f)=\int_{\mathbb{R}^{s}}f(x)\pi(\mathrm{d}x)$, $\pi K$ denote the measure such that $\pi K(A)=\int_{\mathbb{R}^{s}}K(x,A)\pi(\mathrm{d}x)$ for all $A\in\mathcal{B}(\mathbb{R}^{s})$, and $(K^{k})_{k\geq 1}$ denote the sequence of (potentially un-normalized) kernels acting on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$ defined by
$K^{k}(x,A)=\int_{\mathbb{R}^{s}}K(x,\mathrm{d}x^{\prime})K^{k-1}(x^{\prime},A),\quad A\in\mathcal{B}(\mathbb{R}^{s}),\quad k\geq 1$
with the convention that $K^{k-1}(x^{\prime},\mathrm{d}x)=\delta_{\{x^{\prime}\}}(\mathrm{d}x)$ for all $x^{\prime}\in\mathbb{R}^{s}$ when $k=1$. Finally, for any probability distribution $\pi\in\mathcal{P}(\mathbb{R})$ we let $F_{\pi}$ denote the cumulative density function of $\pi$ and $F^{-1}_{\pi}$ denote its (generalized-)inverse.
The rest of the paper is organized as follows. Section 2 introduces formally the filtering problem and algorithms discussed in this work, and the time evolution of the resulting PF estimates is studied in Section 3. The result outlined in the introductory section for SQMC is given in Section 4 and Section 5 concludes. All the proof are gathered in the appendix.
2 Model and filtering algorithms
2.1 The Model
We let $(Y_{t})_{t\geq 1}$ be a sequence of $\mathbb{R}$-valued random variables and throughout this work we consider the SSM which assumes that, for some constants $\rho\in\mathbb{R}$ and $(\sigma,c)\in(0,\infty)^{2}$, and some distribution $\eta_{1}\in\mathcal{P}(\mathbb{R})$, the sequence $(Y_{t})_{t\geq 1}$ is such that
$Y_{t}=X_{t}+c^{-1/2}Z_{t},\quad X_{t+1}=\rho X_{t}+\sigma W_{t+1},\quad\forall t\geq 1$ (2)
with $X_{1}\sim\eta_{1}$ and where the $Z_{t}$’s and $W_{t+1}$’s are independent $\mathcal{N}_{1}(0,1)$ random variables.
We now let $(y_{t})_{t\geq 1}$ be a sequence in $\mathbb{R}$ and, for all $t\geq 1$, we denote by $\hat{\eta}_{t}$ the filtering distribution of $X_{t}$, that is the conditional distribution of $X_{t}$ given that $Y_{s}=y_{s}$ for all $s\in\{1,\dots,t\}$ under the model (2). In addition, for all $t\geq 2$ we denote by $\eta_{t}$ the predictive distribution of $X_{t}$, that is the conditional distribution of $X_{t}$ given that $Y_{s}=y_{s}$ for all $s\in\{1,\dots,t-1\}$ under the model (2).
We assume henceforth that $y_{t}=0$ for all $t\geq 1$ and, for all $z\in\mathbb{R}$, we let $M(z,\mathrm{d}x)=\mathcal{N}_{1}(\rho z,\sigma^{2})$, $G(z)=e^{-cz^{2}/2}$ and $Q(z,\mathrm{d}x)=M(z,\mathrm{d}x)G(x)$. With this notation and assumption on $(y_{t})_{t\geq 1}$, the two sequences of distributions $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are defined by
$\hat{\eta}_{1}(\mathrm{d}x_{1})=\frac{G(x_{1})\eta_{1}(\mathrm{d}x_{1})}{\eta_{1}(G)}$ (3)
and by
$\eta_{t}(\mathrm{d}x_{t})=\hat{\eta}_{t-1}M(\mathrm{d}x_{t}),\quad\hat{\eta}_{t}(\mathrm{d}x_{t})=\frac{\hat{\eta}_{1}Q^{t-1}(\mathrm{d}x_{t})}{\hat{\eta}_{1}Q^{t-1}(\mathbb{R})},\quad t\geq 2.$ (4)
The following proposition shows that the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ admit a well-defined limit in $\mathcal{P}(\mathbb{R})$ and that, as $t\rightarrow\infty$, the impact of the initial distribution $\eta_{1}$ on $\hat{\eta}_{t}$ and on $\eta_{t}$ vanishes exponentially fast:
Proposition 1.
There exist constants $(C_{\star},\sigma^{2}_{\infty})\in(0,\infty)^{2}$ and $\epsilon_{\star}\in(0,1)$ such that
$\big{\|}\eta_{t+1}-\mathcal{N}_{1}(0,\rho\sigma^{2}_{\infty}+\sigma^{2})\big{\|}\leq\|\hat{\eta}_{t}-\mathcal{N}_{1}(0,\sigma^{2}_{\infty})\|\leq C_{\star}\epsilon_{\star}^{t},\quad\forall t\geq 1.$
Remark 1.
When $\eta_{1}$ is a Gaussian distribution the conclusion of the lemma can be obtained from the results in Del Moral and Horton, (2023).
We stress that, despite the simplicity of the filtering problem that we consider in this work, little is known about the finite $N$ behaviour of the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}$ of $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$. Notably, using Corollary 3 in Caffarel et al., (2025) only allows to establish that, for all functions $f$ in some class of functions (containing unbounded functions) and under the assumption that $|\rho|<1$, we have $\sup_{t\geq 1}\mathbb{E}\big{[}|\eta_{t}^{N}(f)-\eta_{t}(f)|]\leq CN^{-\beta/2}$ for some unknown constants $C$ and $\beta\in(0,1]$.
2.2 Particle filters and SQMC
In the specific case where $\eta_{1}=\mathcal{N}_{1}(\mu_{1},\sigma_{1}^{2})$ for some constants $\mu_{1}\in\mathbb{R}$ and $\sigma_{1}^{2}\in(0,\infty)$ the SSM defined in (2) is a linear Gaussian SSM and the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are sequences of Gaussian distributions whose means and variances can be computed using the Kalman filter. For other choices of initial distribution $\eta_{1}$ these two sequences are intractable and therefore need to be approximated.
Algorithm 1 describes a genetic filtering algorithm that can be used to perform this task. When in Algorithm 1 the $U_{t}^{n}$’s are independent $\mathcal{U}(0,1)^{2}$ random variables we recover the bootstrap PF with multinomial resampling, which is arguably the simplest PF algorithm. It is well-known that the variance of PF estimates can be reduced by replacing multinomial resampling by more advanced resampling schemes and, for instance, Algorithm 1 reduces to the (bootstrap) PF with ordered stratified resampling when the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables while the sets ($(\{U_{t,1}^{n}\}_{n=1}^{N})_{t\geq 1}$ used in the resampling steps are independent sets of $N$ dependent $\mathcal{U}(0,1)$ random variables (see Gerber et al.,, 2019, for more details on resampling algorithms for PFs).
Sequential quasi-Monte Carlo goes one step further than PF with stratified resampling in the de-randomization process of the vanilla PF (i.e. of the PF with multinomial resampling) by also introducing, at each time $t\geq 1$, some dependence in the $\mathcal{U}(0,1)$ random numbers $\{U_{t,2}^{n}\}_{n=1}^{N}$ used to perform the mutation step. In addition, for all $t\geq 1$ the $N$ dependent $\mathcal{U}(0,1)^{2}$ random variables $\{U_{t}^{n}\}^{N}_{n=1}$ used in SQMC spread evenly over $(0,1)^{2}$ with $\mathbb{P}$-probability one, that is, with $\mathbb{P}$-probability one these sets of random variables are such that $\sup_{t\geq 1}D^{*}_{2}(\{U_{t}^{n}\}^{N}_{n=1})\leq r_{N}$ for some sequence $r_{N}\rightarrow\infty$.
Formally, we recover SQMC when, in Algorithm 1, for all $t\geq 1$ the set $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(t,2)$-sequence in based $b$ (see Section 4.1 for a definition), and these random sets are independent. In this case, with $\mathbb{P}$-probability one we have $\sup_{t\geq 1}D^{*}_{2}(\{U_{t}^{n}\}_{n=1}^{N})=\mathcal{O}(N^{-1}\log(N)^{2})$. Moreover, there exists a finite universal constant $C$ such that, for all $t\geq 1$ and any square integrable function $f:[0,1]^{2}\rightarrow\mathbb{R}$, the variance of the estimate $\frac{1}{N}\sum_{n=1}^{N}f(U_{t}^{n})$ of $\int_{[0,1]^{2}}f(u)\mathrm{d}u$ is smaller than $C$ times the variance we would have obtained if $\{U_{t}^{n}\}_{n=1}^{N}$ was a collection of independent $\mathcal{U}(0,1)^{2}$ random numbers (see Gerber,, 2015, and references therein). This latter property is important as it ensures that if for a $p=2$ the result in (1) holds for the vanilla PF then it also holds for SQMC. On the other hand, for cost per iteration of SQMC is $\mathcal{O}(N\log(N))$ against $\mathcal{O}(N)$ for PFs.
1:let $\{U_{1}^{n}=(U_{1,1}^{n},U_{1,2}^{n})\}_{n=1}^{N}$ be a random point set in $(0,1)^{2}$
2:let $X_{1}^{n}=F^{-1}_{\eta_{1}}(U_{1,1}^{n})$ and $W_{1}^{n}=G(X_{1}^{n})/\sum_{m=1}^{N}G(X_{1}^{m})$
3:let $\eta_{1}^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{1}^{n}\}}$ and $\hat{\eta}_{1}^{N}=\sum_{n=1}^{N}W_{1}^{n}\delta_{\{X_{1}^{n}\}}$
4:for $t\geq 2$ do
5:     let $\{U_{t}^{n}=(U_{t,1}^{n},U_{t,2}^{n})\}_{n=1}^{N}$ be a random point set in $(0,1)^{2}$
6:     let $\hat{X}_{t-1}^{n}=F^{-1}_{\hat{\eta}_{t-1}^{N}}(U_{t,1}^{n})$ , $X_{t}^{n}=F^{-1}_{M(\hat{X}_{t-1}^{n},\mathrm{d}x)}(U_{t,2}^{n})$ and $W_{t}^{n}=G(X_{t}^{n})/\sum_{m=1}^{N}G(X_{t}^{m})$
7:     let $\eta_{t}^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}$ and $\hat{\eta}_{t}^{N}=\sum_{n=1}^{N}W_{t}^{n}\delta_{\{X_{t}^{n}\}}$
8:end for
Algorithm 1 Generic filtering algorithm
(Operations with index $n$ must be performed for all $n\in\{1,\dots,N\}$)
2.3 Time uniform bounds for deterministic filtering algorithms
In the previous subsection we focussed on stochastic versions of Algorithm 1, that is on the case where $U_{t}^{n}\sim\mathcal{U}(0,1)^{2}$ for all $n\in\{1,\dots,N\}$ and all $t\geq 1$. Alternatively, one may consider determinist versions of the algorithm in which the $U_{t}^{n}$’s are non-random. We do not necessarily advocate for the use of such filtering algorithms but, as it will be clear in what follows, studying Algorithm 1 based on deterministic elements of $(0,1)^{2}$ is a key step for analysing the time evolution of $\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ and of $\|\eta_{t}^{N}-\eta_{t}\|$ for both PF and SQMC. In particular, the following lemma will play a key role to study these two algorithms:
Lemma 1.
Let $N\geq 1$ and for all $t\geq 1$ let $\big{\{}u_{t}^{n}=(u_{t,1},u_{t,2}^{n})\big{\}}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ such that, for $i=1,2$, we have $u_{t,i}\neq a2^{-k}$ for all $k\in\{1,\dots,7\}$ and all $a\in\{0,\dots,2^{7}\}$, and such that $\sup_{t\in\{1,\dots,T\}}D^{*}(\{u_{t}^{n}\}_{n=1}^{N})<2^{-10}$ for some $T\in\mathbb{N}\cup\{\infty\}$. Consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}=\{u_{t}^{n}\}_{n=1}^{N}$ for all $t\geq 1$. Then, there exists a constant $\bar{C}\in(0,\infty)$, depending only on $(c,\sigma,\rho)$ and $\eta_{1}$, such that
$\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\bar{C}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\quad\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\bar{C}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}}$
with $\epsilon_{\star}\in(0,1)$ as in Proposition 1 and with $\delta_{N,T}=\sup_{t\in\{1,\dots,T\}}D^{*}(\{u_{t}^{n}\}_{n=1}^{N})$.
We stress that, in the lemma, we did not aim at optimizing the assumptions on ($(\{u_{t}^{n}\}_{n=1}^{N})_{\geq 1}$ but rather at obtaining a result that holds under weak assumptions on this sequence of sets. In particular, when $\{u_{t}^{n}\}_{n=1}^{N}$ is the realization of an RQMC point set the upper bound on $D^{*}(\{u_{t}^{n}\}_{n=1}^{N})$ imposed in the lemma can be drastically reduced (see Remark 5 in the Appendix).
In practice we are usually interested in estimating the expectation of some function $f$ under the filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$ and under the predictive distributions $(\eta_{t})_{t\geq 1}$. Building on the result of Lemma 1, the following lemma will be used below to address this problem using PFs and SQMC:
Lemma 2.
Consider the set up of Lemma 1 with $T\geq 2$. Then, there exist constants $(\varrho_{1},\varrho_{2})\in(0,\infty)^{2}$, depending only on $c$, $\sigma$, $\rho$ and $\eta_{1}$, such that for any function $f\in\mathcal{C}^{1}(\mathbb{R})$ verifying $\sup_{x\in\mathbb{R}}|f(x)|e^{-\varrho_{1}x^{2}}<\infty$ there is a constant $C_{f}\in(0,\infty)$, depending only on $f$, $c$, $\sigma$, $\rho$ and $\eta_{1}$, for which we have, for all $a\in(0,\infty]$ and all $\gamma\in(0,1)$, and with $\epsilon_{\star}$ and $\bar{C}$ as in Lemma 1,
$\sup_{t\in\{2,\dots,T\}}\bigg{(}\big{|}\hat{\eta}_{t}^{N}\big{(}f\mathbbm{1}_{[-a,a]}\big{)}-\hat{\eta}_{t}(f)\big{|}\vee\big{|}\eta_{t}^{N}\big{(}f\mathbbm{1}_{[-a,a]}\big{)}-\eta_{t}(f)\big{|}\Big{)}\leq\delta_{N,T}^{\gamma/2}\frac{\bar{C}F(a)}{1-\epsilon_{\star}^{1-\gamma}}+C_{f}e^{-\varrho_{2}a^{2}}$
where $F(a)=|f(a)|+|f(-a)|+\int_{-a}^{a}|f^{\prime}(x)|\mathrm{d}x$ with the convention that $F(a)=\lim_{z\rightarrow\infty}F(z)$ when $a=\infty$.
3 Time evolution of particle filter estimates
3.1 Almost sure behaviour of particle filter estimates
We consider the use of Algorithm 1 to approximate the sequences of distributions $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ defined in Section 2.1, where the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables. As explained in Section 2.2, this framework includes the bootstrap PF with both multinomial resampling and some more advanced resampling techniques. In fact, as it will be clear from the calculations that follow, the main result of this subsection (Proposition 2) holds for any resampling mechanism.
Let $N\geq 1$, $\mathcal{F}^{N}_{0}=\{0,\Omega\}$ be the trivial $\sigma$-algebra and for all $t\geq 1$ let
$\mathcal{F}^{N}_{t}=\sigma\Big{(}\big{\{}(X_{s}^{n},\hat{X}_{s}^{n}),\,n=1,\dots,N,\,\,s=1,\dots,t\Big{)}.$
Next, let $[a,b]\subset\mathbb{R}$ be an arbitrary non-empty interval and note that
$\inf_{x^{\prime}:\,\rho x^{\prime}\in[a,b]}M\big{(}x^{\prime},[a,b]^{c}\big{)}\geq 1-\Phi\big{(}2(b-a)/\sigma\big{)},\quad\inf_{x^{\prime}:\,\rho x^{\prime}\not\in[a,b]}M\big{(}x^{\prime},[a,b]^{c}\big{)}\geq\frac{1}{2}.$
Therefore, letting $\varrho=\min\big{\{}1-\Phi(2(b-a)/\sigma),1/2\big{\}}>0$, it follows that
$\mathbb{P}\big{(}\eta_{t}^{N}([a,b])=0\big{|}\mathcal{F}^{N}_{t-1}\big{)}\geq\varrho^{N}>0,\quad\forall t\geq 1,\quad\mathbb{P}-a.s.$ (5)
Then, using the second Borel-Cantelli lemma for dependent events given in Durrett, (2019, Theorem 4.3.4, page 225), the result in (5) implies that with $\mathbb{P}$-probability one we have $\eta_{t}^{N}([a,b])=0$ infinitely often. Since $N\geq 1$ and the interval $[a,b]$ are arbitrary, we obtain in particular that
$\mathbb{P}\big{(}\hat{\eta}_{t}^{N}\big{(}[-a,a]^{c}\big{)}=1,\,\,i.o.\big{)}=\mathbb{P}\big{(}\eta_{t}^{N}\big{(}[-a,a]^{c}\big{)}=1,\,\,i.o.\big{)}=1,\quad\forall N\geq 1,\quad\forall a\in(0,\infty).$ (6)
In words, for any number of particles $N\geq 1$ and interval $[-a,a]\subset\mathbb{R}$, with $\mathbb{P}$-probability one the set $\{X_{t}^{n}\}_{n=1}^{N}$ contains no elements in $[-a,a]$ for infinitely many time instants $t\geq 1$.
By combining (6) with Proposition 1 we readily obtain the first result announced in the introductory section:
Proposition 2.
Let $N\geq 1$ and consider Algorithm 1 where ($\big{(}\{U_{t,2}^{n}\}_{n=1}^{N}\big{)}_{t\geq 1}$ is a sequence of independent sets of independent $\mathcal{U}(0,1)$ random variables. Then,
$\mathbb{P}\big{(}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa,\,\,i.o.\big{)}=\mathbb{P}\big{(}\|\eta_{t}^{N}-\eta_{t}\|\geq\kappa,\,\,i.o.\big{)}=1,\quad\forall N\geq 1,\quad\forall\kappa\in(0,1).$
3.2 Discussion
As shown in the above calculations, the negative result given in (6) (and thus in Proposition 2) for PFs (i.e. for Algorithm 1 where the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables) arises for the following simple and intuitive reason. On the one hand, since for all $t\geq 1$ the uniform random variables $\{U_{t,2}^{n}\}_{n=1}^{N}$ used by a PF are independent, each mutation step of the algorithm moves all the particle outside any given interval $[a,b]\subset\mathbb{R}$ with a constant positive probability. On the other hand, because the different mutation steps of a PF rely on independent sets of uniform random variables this event will necessarily happen infinity often.
Interestingly, in the literature it has been realized that if, in Algorithm 1, the function $G$ is defined by $G(x)=\mathbbm{1}_{\mathbb{R}\setminus I}(x)$ for some bounded interval $I\subset\mathbb{R}$ then, for a PF, it may happen that at some time $t\geq 1$ all the particles $\{X_{t}^{n}\}_{n=1}^{N}$ fall into $I$, in which case the particle system is said to be extinct (Del Moral,, 2004, see, Section 7.2.2, page 219). Surprisingly, it has not been realized that this positive probability to have, at any time $t\geq 1$, all the particles located in a given interval implies that PFs are not suitable for estimating an infinite sequence of filtering distributions.
Above we establish (6) using only the properties of the mutation steps of Algorithm 1. This observation has two key implications. Firstly, considering more sophisticated resampling schemes than those allowed by Algorithm 1 will not prevent (6) to hold. Secondly, this latter result remains true if for all $t\geq 1$ we replace the function $G(x)=e^{-cx^{2}/2}$ by some arbitrary function $G_{t}:\mathbb{R}\rightarrow(0,\infty)$. Notably, for the SSM considered in this work (6) holds for any sequence of observations $(y_{t})_{t\geq 1}$. More generally, by looking at the calculations used to prove (6) it seems hard to think of a meaningful SSM for which (6) does not hold. In particular, we stress that (6) is not due to the fact that we consider an unbounded state-space. To illustrate this point let $\mathsf{X}=[-l,l]$ for some constant $l\in(0,\infty)$, and let $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ be as defined in (3)-(4) where, for all $x^{\prime}\in\mathbb{R}$, $M(x^{\prime},\mathrm{d}x)$ denote the $\mathcal{N}_{1}(\rho x^{\prime},\sigma^{2})$ distribution truncated on $\mathsf{X}$ and where $\eta_{1}$ is such that $\eta_{1}(\mathsf{X})=1$. Then, by repeating the above calculations, it is trivial to verify that if in Algorithm 1 the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables then
$\mathbb{P}\Big{(}\hat{\eta}_{t}^{N}\big{(}[-l,l]\setminus[-a,a]\big{)}=1,\,\,i.o.\Big{)}=\mathbb{P}\Big{(}\eta_{t}^{N}\big{(}[-l,l]\setminus[-a,a]\big{)}=1,\,\,i.o.\Big{)}=1,\quad\forall N\geq 1,\quad\forall a\in(0,l).$
3.3 Finite time horizon and probabilistic behaviour of particle filter estimates
As proved above, as time progress a PF will inevitably reach a time instant $\tau$ at which it will fail at estimating well the filtering and/or the predictive distribution of the SSM of interest. From a practical point of view it is of key interest to study the probability that such an undesirable event happens in a finite number $T$ of time steps.
Interestingly, Lemma 1, obtained for deterministic versions of Algorithm 1, can be used to address this problem. This is the case because if, in the algorithm, the $U_{t}^{n}$’s are independent $\mathcal{U}(0,1)^{2}$ random variables then $\sup_{t\in\{1,\dots,T\}}D^{*}_{2}(\{U_{t}^{n}\}_{n=1}^{N})\leq\sqrt{2/N}$ with high probability, provided that $N$ is sufficiently large. More precisely, using Aistleitner and Hofer, (2014, Theorem 1) we have the following key result:
Lemma 3.
Let $(s,T)\in\mathbb{N}^{2}$, $(\delta,q)\in(0,1)^{2}$ and, for all $N\geq 1$, let $\{\{U_{t}^{n}\}_{n=1}^{N}\}_{t=1}^{T}$ be $T$ independent sets of $N$ independent $\mathcal{U}(0,1)^{s}$ random variables. Then,
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}D^{*}_{s}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}\geq 1-q,\quad\forall N\geq N_{s,T,q}:=\frac{160s+33\log(T)-33\log(q)}{\delta^{2}}.$
By combining Lemma 1 and Lemma 3 we easily obtain the following result for PF with multinomial resampling:
Theorem 1.
Let $N\geq 1$ and consider Algorithm 1 where ($\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}_{t\geq 1}$ is a sequence of independent sets of $N$ independent $\mathcal{U}(0,1)^{2}$ random variables. Then, there exists a constant $\bar{C}_{1}\in[1,\infty)$ (independent of $N$) such that, for all $q\in(0,1)$ and all $T\in\mathbb{N}$, we have
$\mathbb{P}\Bigg{(}\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\bar{C}_{1}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T,q}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\,\,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\bar{C}_{1}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T,q}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}}\Bigg{)}\geq 1-q$
with $\epsilon_{\star}\in(0,1)$ as in Proposition 1 and with $\delta_{N,T,q}=N^{-1}\big{(}160+33\log(T)-33\log(q)\big{)}$.
Remark 2.
A direct application of Lemma 1 and of Lemma 3 would lead the conclusion of Theorem 1 to hold with $\delta_{N,T,q}$ replaced by $\delta_{N,T,q}^{1/2}$. In Theorem 1 we exploit a particular property of independent uniform random variables to improve this result (see the proof of the theorem for more details).
Remark 3.
In Theorem 1 the condition $C_{1}\geq 1$ comes from Lemma 3, and implies that the probabilistic bounds given in the theorem are meaningful (i.e. smaller than one) only when $N>N_{1,T,q}$, with $N_{1,T,q}$ as defined in Lemma 3.
A direct implication of Theorem 1 is that if, for some $\kappa\in(0,1)$, we have
$N\geq\Big{(}160+33\log(T)-33\log(q)\Big{)}\inf_{\gamma\in(0,1)}\kappa^{-\frac{2}{\gamma}}\bigg{(}\frac{\bar{C}_{1}}{1-\epsilon_{\star}^{1-\gamma}}\bigg{)}^{\frac{2}{\gamma}}$ (7)
then with probability at least $1-q$ we have $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ and $\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa$. This constitutes a reassuring result concerning the time evolution of PF estimates for two reasons. Firstly, it shows that, as the time horizon $T$ that we consider increases, the number of particles $N$ should grow at most at speed $\log(T)$ to ensure that the PF algorithm estimates the whole sets of distributions $\{\hat{\eta}_{t}\}_{t=1}^{T}$ and $\{\eta_{t}\}_{t=1}^{T}$ with a constant (probabilistic) error. Secondly, Theorem 1 establishes that $N$ should grow at most at speed $\log(1/q)$ if we want to increase the confidence level $1-q$.
The following proposition shows that these results concerning the dependence in $T$ and in $q$ of the number of particles needed to have both $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ and $\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa$ with probability at least $1-q$ are sharp:
Proposition 3.
Consider the set-up of Theorem 1. Then, for all $\kappa\in(0,1)$ there exist constants $\varrho_{\kappa}\in(0,1)$ and $T_{\kappa}\in\mathbb{N}$ such that, for all $q\in(0,1)$ and all $T\geq T_{\kappa}$, we have
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa\Big{)}<1-q,\quad\forall N<\frac{\log(T)-\log(q)}{\log(1/\varrho_{\kappa})}.$
By using Lemma 2 and Theorem 1 we can easily obtain probabilistic bounds for the joint estimation of all the filtering and predictive expectations $\{\hat{\eta}_{t}(f)\}_{t=1}^{T}$ and $\{\eta_{t}(f)\}_{t=1}^{T}$ for a large class of functions, as illustrated with the following proposition:
Proposition 4.
Consider the set-up of Theorem 1 and let $\gamma\in(0,1)$ and $\mathfrak{F}\subset\mathcal{C}^{1}(\mathbb{R})$. Then, there exists a constant $C_{\mathfrak{F}}\in(0,\infty)$ such that, for all $T\geq 2$, all $q\in(0,1)$ and all $N\geq 1$, we have
$\mathbb{P}\bigg{(}\sup_{t\in\{2,\dots,T\}}\sup_{f\in\mathfrak{F}}\Big{(}\big{|}\hat{\eta}_{t}^{N}\big{(}f\mathbbm{1}_{[-a_{N},a_{N},]}\big{)}-\hat{\eta}_{t}(f)\big{|}\vee\big{|}\eta_{t}^{N}\big{(}f\mathbbm{1}_{[-a_{N},a_{N}]}\big{)}-\eta_{t}(f)\big{|}\Big{)}\leq C_{\mathfrak{F}}\,\delta_{N,T,q}^{\gamma\beta/2}\bigg{)}\geq 1-q$
for
  • $\mathfrak{F}=\big{\{}f\in\mathcal{C}^{1}(\mathbb{R}):\,\int_{\mathbb{R}}|f^{\prime}(x)|\mathrm{d}x=1\big{\}}$, if $a_{N}=\infty$ and if $\beta=1$.
  • $\mathfrak{F}=\big{\{}f\in\mathcal{C}^{1}(\mathbb{R}):\,|f(x)|\leq 1+|x|\text{ and }|f^{\prime}(x)|\leq 1,\,\forall x\in\mathbb{R}\big{\}}$, if $a_{N}=(1\wedge\delta_{N,T,q})^{-\gamma\alpha/2}$ for some constant $\alpha\in(0,1)$ and if $\beta=1-\alpha$.
  • $\mathfrak{F}=\big{\{}f\in\mathcal{C}^{1}(\mathbb{R}):\,|f(x)|\leq 1+e^{\zeta|x|}\text{and }|f^{\prime}(x)|\leq e^{\zeta|x|},\,\,\forall x\in\mathbb{R}\big{\}}$ for some constant $\zeta\in(0,\infty)$, if $a_{N}=-\gamma\alpha/(2\zeta)\log(1\wedge\delta_{N,T,q})$ for some constant $\alpha\in(0,1)$ and if $\beta=1-\alpha$.
In the last two examples given in Proposition 4, estimating e.g. $\hat{\eta}_{t}(f)$ with the truncated estimator $\hat{\eta}_{t}^{N}(f\mathbbm{1}_{[-a_{N},a_{N}]})$ is unusual (the typical estimate being the empirical counterpart $\hat{\eta}_{t}^{N}(f)$) but has an intuitive explanation. In these two examples the function $f$ is unbounded and thus, informally speaking, even a small estimation error of $\hat{\eta}_{t}$ in a region $R\subset\mathbb{R}$ where $\inf_{x\in R}|f(x)|$ is large will have a significant impact on the overall estimation error $|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|$.
4 Almost sure behaviour of sequential quasi-Monte Carlo
4.1 Remainder on scrambled $(t,s)$-sequences
Let $s\geq 1$ and $b\geq 2$ be two integers,
$\left\{\prod_{j=1}^{s}\left[a_{j}b^{-d_{j}},(a_{j}+1)b^{-d_{j}}\right)\subseteq[0,1)^{s},\,a_{j},\,d_{j}\in\mathbb{N},\,a_{j}<b^{d_{j}},\,j=1,...,s\right\}$
be the set of all $b$-ary boxes and $t\in\mathbb{N}_{0}$.
Then, for an integer $m\geq t$ the point set $\{u^{n}\}_{n=1}^{b^{m}}$ is called a $(t,m,s)$-net in base $b$ if every $b$-ary box of volume $b^{t-m}$ contains exactly $b^{t}$ points, while a sequence $(u^{n})_{n\geq 1}$ of points in $[0,1)^{s}$ is called a $(t,s)$-sequence in base $b$ if, for any integers $a\geq 0$ and $m\geq t$, the point set $\{u^{n}\}_{n=ab^{m}+1}^{(a+1)b^{m}}$ is a $(t,m,s)$-net in base $b$. From these definitions it should be clear that, from a practical perspective, $(t,s)$-sequences in base $b$ with $t=0$ and $b=2$ are preferable but such sequences exist only for $s\in\{1,2\}$, respectively (Niederreiter,, 1992, Corollary 4.24, page 62). In the context of this work, the crucial property of a $(t,s)$-sequence $(u^{n})_{n\geq 1}$ in base $b$ is to be such that, for some constant $C_{s,t,b}\in(0,\infty)$ depending only on $s$,$t$ and $b$, we have (see Niederreiter,, 1992, Section 4.1)
$D^{*}_{s}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}\leq C_{s,t,b}\frac{\log(N)^{s}}{N},\quad\forall N\geq b^{t}.$ (8)
Finally, a sequence $(U^{n})_{n\geq 1}$ of $(0,1)^{s}$-valued random variables is a scrambled $(t,s)$-sequence in base $b$ if, $\mathbb{P}$-a.s., $(U^{n})_{n\geq 1}$ is a $(t,s)$-sequence in base $b$ and if $U^{n}\sim\mathcal{U}(0,1)^{s}$ for all $n\geq 1$. A scrambled $(t,s)$-sequence is obtained by randomizing, or scrambling, a $(t,s)$-sequence, and various scrambling methods have been proposed in the literature. The first $N$ points of a scrambled sequence can be generated in $\mathcal{O}(N\log(N))$ operations but the memory requirements of the algorithm needed to perform this task depend on the scrambling procedure that is used (Hong and Hickernell,, 2003). We refer to Chapter 3 in Dick and Pillichshammer, (2010) and to Owen, (2023) for a discussion on scrambling algorithms.
4.2 Main result
In the context of this work we only need point sets in $(0,1)^{2}$ and thus, following the discussion of the previous subsection, we limit our attention to scrambled $(0,2)$-sequences in base $b=2$. More precisely, in what follows we study Algorithm 1 in the case where $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$. In this situation we have, using Niederreiter, (1992, Theorem 4.7, page 54 and Theorem 4.14, page 59),
$\sup_{t\geq 1}D^{*}_{2}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta_{N},\quad\forall N\geq 1\quad\mathbb{P}-a.s.$
with
$\delta_{N}=\begin{cases}\frac{\log(N)+3}{2N},&\text{ if }N\in\{2^{k},\,k\in\mathbb{N}\}\\ \frac{\log(N)^{2}+11\log(2)\log(N)+18\log(2)^{2}}{N8\log(2)^{2}},&\text{otherwise}\end{cases},\quad\forall N\geq 1.$ (9)
A direct application of Lemma 1 then gives the following almost sure and time uniform guarantee for the resulting algorithm:
Theorem 2.
Let $N\geq 1$ and consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$. Then, there exists a constant $\bar{C}_{2}\in(0,\infty)$ (independent of $N$) such that, with $\delta_{N}$ as defined in (9) and with $\epsilon_{\star}\in(0,1)$ as in Proposition 1,
$\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\bar{C}_{2}\inf_{\gamma\in(0,1)}\frac{\delta_{N}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\quad\sup_{t\geq 1}\|\eta_{t}^{N}-\eta_{t}\|\leq\bar{C}_{2}\inf_{\gamma\in(0,1)}\frac{\delta_{N}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\quad\mathbb{P}-a.s.$
Remark 4.
Theorem 2 does not assume that the random point sets $\{U_{t}^{n}\}_{n=1}^{N}$’s are independent.
By using Lemma 2 and Theorem 2 we can easily obtain almost sure time uniform bounds for the SQMC estimates of the filtering and predictive expectations $(\hat{\eta}_{t}(f))_{t\geq 1}$ and $(\eta_{t}(f))_{t\geq 1}$ for a large class of functions. For instance, for SQMC the conclusion of Proposition 4 holds with $T=\infty$, $q=1$ and with $\delta_{N,T,q}$ replaced by $\delta_{N}$.
4.3 Discussion
To simplify the discussion we focus below on the estimation of the filtering distributions but the same comments hold for the estimation of the predictive distributions $(\eta_{t})_{t\geq 1}$.
Due to the presence of the $\log(N)$ term in the definition of $\delta_{N}$, the dependence in $N$ of bound given in Theorem 2 for $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ is slightly worse than that of the probabilistic bound obtained in Theorem 1 for PFs. On the other hand, in the context of this work, using RQMC within PF algorithms has the advantage to produce a non-linear filleting algorithm having time uniform and almost sure guarantees. In particular, it follows from Theorem 2 that if $N\in\{2^{k},\,k\in\mathbb{N}\}$ and, for some $\kappa\in(0,1)$, we have
$N\geq\frac{\log(N)+3}{2}\inf_{\gamma\in(0,1)}\kappa^{-\frac{2}{\gamma}}\bigg{(}\frac{\bar{C}_{2}}{1-\epsilon_{\star}^{1-\gamma}}\bigg{)}^{\frac{2}{\gamma}}$ (10)
then $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ with $\mathbb{P}$-probability one.
Unfortunately, we cannot compare the lower bound on $N$ given in (10) with the one obtained in (7), used to ensure that with probability at least $1-q$ the PF estimates $\{\hat{\eta}_{t}^{N}\}_{t=1}^{T}$ are such that $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$, the reason being that there is no reason for the constants $\bar{C}_{1}$ and $\bar{C}_{2}$ appearing in these bounds to be identical. Given the properties of scrambled $(t,s)$-sequences, reminded in Section 4.1, it is however reasonable to expect that, in the worst case, the constant $\bar{C}_{2}$ in (10) can only be slightly larger than the constant $\bar{C}_{1}$ appearing in (7). In any cases, if we want to compute estimates $\{\hat{\eta}_{t}^{N}\}_{t=1}^{T}$ of $\{\hat{\eta}_{t}\}_{t=1}^{T}$ such that $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ with probability at least $1-q$, then it follows from Theorem 1 and Theorem 2 that there exists a value $T_{\kappa,q}\in\mathbb{N}$ such that SQMC is computationally cheaper than PFs for performing this task for any $T\geq T_{\kappa,q}$. Based on (7) and (10), it is reasonable to expect that $T_{\kappa,q}$ is an increasing function of both $\kappa$ and $q$.
The conclusions drawn in this subsection are based on a toy filtering problem, and future research should aim at establishing if they remain valid when considering more realistic filtering tasks.
5 Concluding remarks
Given the popularity of PF algorithms in many applications it is important to understand the behaviour of the resulting joint estimation error, $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$. Surprisingly, with the exception of Del Moral, (2013, Corollary 14.3.3, page 435), this latter quantity (or a variant of it) does not seem to have been previously studied in the literature. The results we obtained in this paper are reassuring, but an important question that has yet to be studied is how the dimension of the filtering problem impacts the behaviour of $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.
Lemma 1, the key result that allows us to study $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ for PFs as well as the quantity $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ for SQMC, is obtained for a toy filtering problem but we expect that it can be generalized by following a similar proof technique. The proof of this lemma, which studies the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ for a given realization $(u_{i})_{i\geq 1}$ of the $\mathcal{U}(0,1)$ random numbers underpinning the algorithm, relies notably on three key building blocks, namely a result showing that the SSM forgets exponentially quickly its initial distribution $\eta_{1}$ as $t\rightarrow\infty$ (illustrated by Proposition 1), a bound for $\big{\|}\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}-\hat{\eta}_{t-1}^{N}M\big{\|}$ (Lemma 8 in the appendix) and a stability result for the particles generated by the algorithm (Lemma 9 in the appendix). The first of these three building blocks is standard, in the sense that in the literature on PFs all the existing bounds for $\sup_{t\geq 1}\mathbb{E}[|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|]$ have been established under the assumption that the model forgets exponentially quickly its initial distribution (see Caffarel et al.,, 2025, and references therein), an assumption that has been shown to hold for a large class of SSMs (see e.g. Del Moral et al.,, 2023). By contrast, to prove Lemma 1 we derive a bound for $\big{\|}\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}-\hat{\eta}_{t-1}^{N}M\big{\|}$ which applies only to some one-dimensional SSMs and a stability result for the particle system which is tailored to the specific filtering problem addressed in this work. We stress that this latter result is quite strong, as it shows (under a reasonable assumption on $(u_{i})_{i\geq 1}$) that there exists a bounded interval $I\subset\mathbb{R}$ that contains, at any time instant $t\geq 1$, at least a proportion $\varsigma>0$ of the particles. Our conjecture is that a result similar to that of Lemma 1 can be established for more realistic filtering problems where this strong stability property of the particle system cannot hold. For instance, we expect that this is the case for the SSM considered in this paper if, instead of assuming that $y_{t}=0$ for all $t\geq 1$, we assume that the sequence of observations $(y_{t})_{t\geq 1}$ is unbounded but such that $\sup_{t\geq 1}|y_{t+1}-y_{t}|<\infty$.
In this work, we focussed on the use of scrambled $(t,s)$-sequences within PF algorithms to improve the safety of the filtering procedure, that is to obtain filtering algorithms for which we have $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|=0$ with $\mathbb{P}$-probability one. Future research should aim at determining if other approaches can be used to derive safe filtering algorithms.
Appendix A Proofs
A.1 Additional notation
For a function $f\in\mathcal{C}^{1}(\mathbb{R})$ we let $V(f)=\int_{\mathbb{R}}|f^{\prime}(x)|\mathrm{d}x$ with the convention that $V(f)=0$ if $f^{\prime}(x)=0$ for all $x\in\mathbb{R}$, and, abusing notation, for a function $f:\mathbb{R}\rightarrow\mathbb{R}$ continuously differentiable on some non-empty interval $(a,b)$ we let $V(\mathbbm{1}_{(a,b]}f)=V(\mathbbm{1}_{(a,b)}f)=\int_{a}^{b}|f^{\prime}(x)|\mathrm{d}x$. For all $(z,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $M_{r,s^{2}}(z,\mathrm{d}x)=\mathcal{N}_{1}(rz,s^{2})$ and $G_{s}(z)=e^{-sz^{2}/2}$, and for any bounded function $H:\mathbb{R}\rightarrow(0,\infty)$ we let $\Psi_{H}:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be defined by
$\Psi_{H}(\pi)=\frac{H(x)\pi(\mathrm{d}x)}{\pi(H)},\quad\pi\in\mathcal{P}(\mathbb{R}).$
Finally, we let $\varphi(\cdot)$ and $\Phi(\cdot)$ denote the probability and cumulative density function of the $\mathcal{N}_{1}(0,1)$ distribution, respectively, and for all $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $h_{a,r,s}:\mathbb{R}\rightarrow[-1,1]$ be defined by
$h_{a,r,s}(x)=\Phi\bigg{(}\frac{a-rx}{s}\bigg{)}-\Phi\bigg{(}\frac{a}{s}\bigg{)},\quad x\in\mathbb{R}.$ (11)
A.2 Preliminary results
A.2.1 A simple but useful technical lemma
Lemma 4.
For all $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we have $V(h_{a,r,s})\leq 1$ and
$|h_{a,r,s}(x)|\leq\frac{|rx|}{s\sqrt{2\pi}},\quad\forall x\in\mathbb{R}.$
Proof.
Let $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$. If $r=0$ the result of the lemma is trivial and we therefore assume that $r\neq 0$ in what follows. Then,
$V(h_{a,r,s})=\int_{\mathbb{R}}|h^{\prime}_{a,r,s}(x)|\mathrm{d}x$ $=\int_{\mathbb{R}}\frac{|r|}{s}\varphi\Big{(}\frac{a-rx}{s}\bigg{)}\mathrm{d}x=\int_{\mathbb{R}}\frac{1}{s}\varphi\Big{(}\frac{a-z}{s}\Big{)}\mathrm{d}z=1$
showing the first part of the lemma. On the other hand, using the mean value theorem and noting that $|\Phi^{\prime}(x)|=|\varphi(x)|\leq 1/\sqrt{2\pi}$ for all $x\in\mathbb{R}$, we have
$|h_{a,r,s}(x)|$ $=\bigg{|}\Phi\bigg{(}\frac{a-rx}{s}\bigg{)}-\Phi\bigg{(}\frac{a}{s}\bigg{)}\bigg{|}\leq\frac{|rx|}{s\sqrt{2\pi}},\quad\forall x\in\mathbb{R}$
showing the second part of the lemma. The proof is complete. ∎
A.2.2 A first useful property of the model of Section 2.1
Lemma 5.
Let $c_{0}=\sigma_{0}=0$, $\rho_{0}=1$, and let $(c_{k})_{k\geq 1}$, $(\rho_{k})_{k\geq 1}$ and $(\sigma_{k})_{k\geq 1}$ be such that
$c_{k}=\frac{(c+c_{k-1})\rho^{2}}{1+(c+c_{k-1})\sigma^{2}},\quad\rho_{k}=\prod_{i=0}^{k-1}\frac{\rho}{1+(c+c_{i})\sigma^{2}},\quad\sigma_{k}^{2}=\sigma_{k-1}^{2}+\frac{\rho_{k-1}^{2}\sigma^{2}}{1+(c+c_{k-1})\sigma^{2}},\quad\forall k\geq 1.$
Then, for all $k\geq 1$, we have
$Q^{k}(x^{\prime},\mathrm{d}x)\propto G_{c_{k}}(x^{\prime})M_{\rho_{k},\sigma^{2}_{k}}(x^{\prime},\mathrm{d}x),\quad\forall x^{\prime}\in\mathbb{R}.$ (12)
In addition, the sequence $(c_{k})_{k\geq 0}$ is such that $c\rho^{2}/(1+c\sigma^{2})\leq c_{k}\leq\rho^{2}/\sigma^{2}$ for all $k\geq 1$, and there exist constants $(\sigma_{\infty},C_{\star})\in(0,\infty)^{2}$ and $\epsilon_{\star}\in(0,1)$ such that $|\rho_{k}|\leq C_{\star}\epsilon_{\star}^{k}$ and such that $|\sigma_{k}^{2}-\sigma_{\infty}^{2}|\leq C^{2}_{\star}\epsilon_{\star}^{2k}$ for all $k\geq 0$.
Proof.
The result in the first part of the lemma can be deduced from the calculations in Del Moral and Jacod, (2001) while the convergence result for the sequence $(\sigma_{k})_{k\geq 1}$ can be directly obtained from Del Moral and Horton, (2023, Theorem 3.5). However, for sake of completeness the whole lemma is proved in what follows.
To prove the first part of the lemma remark first that, for all $a\in(0,\infty)$, we have
$M_{\rho,\sigma^{2}}(x^{\prime},\mathrm{d}x)G_{a}(x)\propto G_{\tilde{a}}(x^{\prime})M_{\tilde{\rho},\tilde{\sigma}^{2}}(x^{\prime},\mathrm{d}x),\quad\forall x^{\prime}\in\mathbb{R}$ (13)
where
$\tilde{a}=\frac{a\rho^{2}}{1+a\sigma^{2}},\quad\tilde{\rho}=\frac{\rho}{1+a\sigma^{2}},\quad\tilde{\sigma}^{2}=\frac{\sigma^{2}}{1+a\sigma^{2}}.$ (14)
We now prove (12) by induction on $k$. For $k=1$ the result given in (12) directly follows by applying (13)-(14) with $a=c$. Assume now that (12) holds for some $k\geq 1$. Then, by using first the inductive hypothesis and then by applying (13)-(14) with $a=c+c_{k}$, for all $x^{\prime\prime}\in\mathbb{R}$ we have
$M_{\rho,\sigma^{2}}(x^{\prime\prime},\mathrm{d}x^{\prime})G_{c}(x^{\prime})Q^{k}(x^{\prime},\mathrm{d}x)$ $\propto M_{\rho,\sigma^{2}}(x^{\prime\prime},\mathrm{d}x^{\prime})G_{c}(x^{\prime})G_{c_{k}}(x^{\prime})M_{\rho_{k},\sigma^{2}_{k}}(x^{\prime},\mathrm{d}x)$
$=M_{\rho,\sigma^{2}}(x^{\prime\prime},\mathrm{d}x^{\prime})G_{c+c_{k}}(x^{\prime})M_{\rho_{k},\sigma^{2}_{k}}(x^{\prime},\mathrm{d}x)$
$\propto G_{c_{k+1}}(x^{\prime\prime})M_{\tilde{\rho}_{k},\tilde{\sigma}_{k}^{2}}(x^{\prime\prime},\mathrm{d}x^{\prime})M_{\rho_{k},\sigma^{2}_{k}}(x^{\prime},\mathrm{d}x)$
with
$\tilde{\rho}_{k}=\frac{\rho}{1+(c+c_{k})\sigma^{2}},\quad\tilde{\sigma}_{k}^{2}=\frac{\sigma^{2}}{1+(c+c_{k})\sigma^{2}}.$
Therefore, for all $x^{\prime\prime}\in\mathbb{R}$ we have
$Q^{k+1}(x^{\prime\prime},\mathrm{d}x)$ $\propto G_{c_{k+1}}(x^{\prime\prime})\int_{\mathbb{R}}M_{\tilde{\rho}_{k},\tilde{\sigma}_{k}^{2}}(x^{\prime\prime},\mathrm{d}x^{\prime})M_{\rho_{k},\sigma^{2}_{k}}(x^{\prime},\mathrm{d}x)$
$=G_{c_{k+1}}(x^{\prime\prime})M_{\tilde{\rho}_{k}\rho_{k},\rho_{k}^{2}\tilde{\sigma}^{2}_{k}+\sigma^{2}_{k}}(x^{\prime\prime},\mathrm{d}x)$
$=G_{c_{k+1}}(x^{\prime\prime})M_{\rho_{k+1},\sigma^{2}_{k+1}}(x^{\prime\prime},\mathrm{d}x)$
and the proof of (12) is complete.
To prove the second part of the lemma let $f:[0,\infty)\rightarrow[0,\infty)$ be defined by
$f(x)=\frac{(c+x)\rho^{2}}{1+(c+x)\sigma^{2}},\quad x\in[0,\infty)$
so that $c_{k}=f(c_{k-1})$ for all $k\geq 1$. Easy computations show that the function $f$ is non-decreasing on $[0,\infty)$, and thus the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and such that $c\rho^{2}/(1+c\sigma^{2})\leq c_{k}\leq\rho^{2}/\sigma^{2}$ for all $k\geq 1$.
To study the sequence $(\rho_{k})_{k\geq 0}$ assume first that $|\rho|<1+c\sigma^{2}$. In this case, we trivially have that $|\rho_{k}|\leq\epsilon_{\star}^{k}$ for all $k\geq 0$ with $\epsilon_{\star}=|\rho|/(1+c\sigma^{2})$. Assume now that $|\rho|\geq 1+c\sigma^{2}$ and note that, for all $k\geq 1$, we have
$\frac{|\rho|}{1+(c+c_{k})\sigma^{2}}<1$ $\Leftrightarrow c_{k}>\frac{|\rho|-(1+c\sigma^{2})}{\sigma^{2}}.$ (15)
To proceed further we remark that, since the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and bounded, the sequence converges in $[0,\infty)$ and we let $c_{\infty}=\lim_{k\rightarrow\infty}c_{k}$. In addition, simple calculations show that
$c_{\star}:=\frac{\rho^{2}-(1+c\sigma^{2})+\sqrt{(1+c\sigma^{2}-\rho^{2})^{2}+4c\sigma^{2}\rho^{2}}}{2\sigma^{2}}$
is the unique value of $x\in[0,\infty)$ such that $x=f(x)$, and since we are assuming that $|\rho|\geq 1+c\sigma^{2}>1$ the constant $c_{\star}$ is such that
$c_{\star}\geq\frac{\rho^{2}-(1+c\sigma^{2})+\sqrt{(1+c\sigma^{2}-\rho^{2})^{2}}}{2\sigma^{2}}=\frac{\rho^{2}-(1+c\sigma^{2})}{\sigma^{2}}>\frac{\rho-(1+c\sigma^{2})}{\sigma^{2}}.$ (16)
On the other hand, since $f$ is continuous on $[0,\infty)$ we have
$c_{\infty}=\lim_{k\rightarrow\infty}c_{k+1}=\lim_{k\rightarrow\infty}f(c_{k})=f(\lim_{k\rightarrow\infty}c_{k})=f(c_{\infty})$
showing that $c_{\infty}=c_{\star}$, and thus $\lim_{k\rightarrow\infty}c_{k}=c_{\star}$. Together with (16), this implies that there exists a $k_{\star}\in\mathbb{N}$ such that
$c_{k_{\star}}>\frac{|\rho|-(1+c\sigma^{2})}{\sigma^{2}}$
and thus, using (15) and the fact that the sequence $(c_{k})_{k\geq 0}$ is non-decreasing, we have
$\frac{|\rho|}{1+(c+c_{k})\sigma^{2}}\leq\epsilon_{\star}:=\frac{|\rho|}{1+(c+c_{k_{\star}})\sigma^{2}}<1,\quad\forall k\geq k_{\star}$
implying that $|\rho_{k}|\leq C\epsilon_{\star}^{k}$ for all $k\geq 0$ and with $C=\big{(}|\rho|/(1+c\sigma^{2})\big{)}^{k_{\star}+1}$.
Finally, to study the sequence $(\sigma_{k})_{k\geq 0}$ note first that
$\sigma_{k}^{2}$ $=\sum_{s=0}^{k-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}},\quad\forall k\geq 1$
and thus, for all integers $1\leq k<m$ we have, with $C$ and $\epsilon_{\star}$ as above,
$|\sigma_{m}^{2}-\sigma_{k}^{2}|=\Big{|}\sum_{s=k}^{m-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}\Big{|}\leq\frac{\sigma^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\rho_{s}^{2}\leq\frac{\sigma^{2}C^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\epsilon_{\star}^{2s}\leq\epsilon_{\star}^{2k}\frac{\sigma^{2}C^{2}}{(1-\epsilon_{\star})(1+c\sigma^{2})}.$ (17)
Using (17) it is easily verified that the sequence $(\sigma^{2}_{k})_{k\geq 1}$ is Cauchy and therefore converges in $[0,\infty)$, and we let $\sigma^{2}_{\infty}=\lim_{k\rightarrow\infty}\sum_{s=0}^{k-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}$. Remark that since the sequence $(\sigma^{2}_{k})_{k\geq 1}$ is non-decreasing and such that $\sigma^{2}_{k}\geq\sigma^{2}/(1+c\sigma^{2})$ it follows that $\sigma^{2}_{\infty}>0$, and remark that, by (17), we have
$|\sigma^{2}_{\infty}-\sigma_{k}^{2}|\leq\epsilon_{\star}^{2k}\frac{\sigma^{2}C^{2}}{(1-\epsilon_{\star})(1+c\sigma^{2})},\quad\forall k\geq 1.$
The proof of the lemma is complete. ∎
A.2.3 A second useful property of the model of Section 2.1
The following lemma will be used to prove Lemma 2:
Lemma 6.
There exist constants $(\varrho_{1},\varrho_{2})\in(0,\infty)^{2}$ such that, for any function $f:\mathbb{R}\rightarrow\mathbb{R}$ verifying $\sup_{x\in\mathbb{R}}|f(x)|e^{-\varrho_{1}x^{2}}<\infty$, there is a constant $C_{f}\in(0,\infty)$ for which we have
$\sup_{t\geq 1}\big{|}\hat{\eta}_{t}(f\mathbbm{1}_{[-b,b]^{c}})\big{|}\leq C_{f}e^{-\varrho_{2}b^{2}},\quad\sup_{t\geq 1}\big{|}\eta_{t}(f\mathbbm{1}_{[-b,b]^{c}})\leq C_{f}e^{-\varrho_{2}b^{2}},\quad\forall b\in(0,\infty).$
Proof.
We start with some preliminary computations. Let $b\in[2,\infty)$, $(a,r)\in\mathbb{R}^{2}$, $s\in(0,\infty)$,
$\bar{\delta}_{a,r,s^{2}}=\min\bigg{\{}\frac{1}{s^{2}},\frac{a}{as^{2}+r^{2}}\bigg{\}}$ (18)
and $\Psi(\cdot;a,s^{2})$ denote the cumulative density function of the $\mathcal{N}_{1}(a,s^{2})$ distribution. In addition, for all $\delta\in(0,\infty)$ and function $f:\mathbb{R}\rightarrow\mathbb{R}$ we let $C_{f,\delta}=\sup_{x\in\mathbb{R}}|f(x)|e^{-(\delta/2)x^{2}}\in(0,\infty]$.
We now let $\delta\in(0,\bar{\delta}_{a,r,s^{2}})$ be fixed and $f:\mathbb{R}\rightarrow\mathbb{R}$ be such that $C_{f,\delta}<\infty$. Then, for all $x\in\mathbb{R}$ we have
$\begin{split}\big{|}G_{a}(x)M_{r,s^{2}}\big{(}x,f\mathbbm{1}_{[-b,b]^{c}}\big{)}\big{|}&\leq\frac{C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{x^{2}}{2}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}M_{\frac{r}{1-\delta s^{2}},\frac{s^{2}}{1-\delta s^{2}}}\big{(}x,[-b,b]^{c}\big{)}\\ &\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{x^{2}}{2}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}\Phi\Big{(}(b,\infty);\frac{|rx|}{1-\delta s^{2}},\frac{s^{2}}{1-\delta s^{2}}\Big{)}.\end{split}$ (19)
To proceed further let $x\in\mathbb{R}$ be such that $b\geq 1+3|rx|/(1-\delta s^{2})$. Then, recalling that for all $z>0$ we have $\Phi\big{(}(z,\infty);0,1\big{)}\leq e^{-z^{2}/2}/z$, it follows that
$\Phi\Big{(}(b,\infty);\frac{|rx|}{1-\delta s^{2}},\frac{s^{2}}{1-\delta s^{2}}\Big{)}\leq\exp\bigg{(}-\frac{1-\delta s^{2}}{2s^{2}}\Big{(}b-\frac{|rx|}{1-\delta s^{2}}\Big{)}^{2}\bigg{)}\frac{s}{(1-\delta s^{2})^{1/2}}$
which, together with (19), implies that
$|G_{a}(x)M_{r,s^{2}}(x,f\mathbbm{1}_{[-b,b]^{c}})|$ $\leq\frac{2sC_{f,\delta}}{1-\delta s^{2}}\exp\Big{(}-\alpha b^{2}-\beta x^{2}+\gamma b|x|\Big{)}$ (20)
with
$\alpha=\frac{1-\delta s^{2}}{2s^{2}},\quad\beta=\frac{1}{2}\Big{(}\frac{r^{2}}{s^{2}}+a\Big{)},\quad\gamma=\frac{|r|}{s^{2}}.$
If $r\neq 0$ we have, recalling that we are assuming that $x$ is such that $b\geq 1+3|rx|/(1-\delta s^{2})$,
$-\alpha b^{2}+\gamma b|x|\leq-\alpha b^{2}+\gamma\frac{1-\delta s^{2}}{3|r|}b^{2}$ $=-\frac{1-\delta s^{2}}{2s^{2}}b^{2}+\frac{1-\delta s^{2}}{3s^{2}}b^{2}=-\frac{1-\delta s^{2}}{6s^{2}}b^{2}$
while if $r=0$ we have $-\alpha b^{2}+\gamma b|x|=-\alpha b^{2}\leq-(1-\delta s^{2})/(6s^{2})$. Together with (20), this shows that
$b\geq 1+3|rx|/(1-\delta s^{2})\implies\big{|}G_{a}(x)M_{r,s^{2}}\big{(}x,f\mathbbm{1}_{[-b,b]^{c}}\big{)}\big{|}$ $\leq\frac{2sC_{f,\delta}}{1-\delta s^{2}}\exp\Big{(}-\frac{1-\delta s^{2}}{6s^{2}}b^{2}\Big{)}.$ (21)
Assume now that $x$ is such that $b<1+3|rx|/(1-\delta s^{2})$, so that $x^{2}>(1-\delta s^{2})^{2}(b-1)^{2}/(9r^{2})$. (Remark that since $b\geq 2$ we can have $b<1+3|rx|/(1-\delta s^{2})$ only if $r\neq 0$). Then, noting that since $\delta\in(0,\bar{\delta}_{a,r,s^{2}})$ we have $a-\delta r^{2}/(1-\delta s^{2})>0$, it directly follows from (19) that
$\begin{split}b<1+3|rx|/(1-\delta s^{2})\implies\big{|}G_{a}(x)&M_{r,s^{2}}\big{(}x,f\mathbbm{1}_{[-b,b]^{c}}\big{)}\big{|}\\ &\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{x^{2}}{2}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}\\ &\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{(b-1)^{2}(1-\delta s^{2})^{2}}{18r^{2}}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}\\ &\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{b^{2}(1-\delta s^{2})^{2}}{36r^{2}}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}\end{split}$ (22)
where the last inequality uses the fact that $(b-1)\geq b/2$ since $b\geq 2$.
Then, by using (21) and (22), it follows that
$\big{|}G_{a}(x)M_{r,s^{2}}\big{(}x,f\mathbbm{1}_{[-b,b]^{c}}\big{)}\big{|}\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\bigg{(}1+\frac{s}{(1-\delta s^{2})^{1/2}}\bigg{)}e^{-b^{2}K(a,r,s^{2},\delta)},\quad\forall x\in\mathbb{R}$ (23)
with
$K(a,r,s^{2},\delta)=\min\bigg{\{}\frac{1-\delta s^{2}}{6s^{2}},\frac{(1-\delta s^{2})^{2}}{36r^{2}}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{\}}.$
We are now in position to prove the lemma. To show the first part of the lemma let $t\geq 2$, $b\in[2,\infty)$, and let $(c_{k})_{k\geq 0}$, $(\rho_{k})_{k\geq 0}$ and $(\sigma_{k})_{k\geq 0}$ be as defined in Lemma 5. Then, by using (4) and Lemma 5, for any function $f:\mathbb{R}\rightarrow\mathbb{R}$ we have
$\hat{\eta}_{t}\big{(}f\mathbbm{1}_{[-b,b]^{c}}\big{)}=\frac{\int_{\mathbb{R}}G_{c_{t-1}}(x)M_{\rho_{t-1},\sigma^{2}_{t-1}}(x,f\mathbbm{1}_{[-b,b]^{c}})\hat{\eta}_{1}(\mathrm{d}x)}{\hat{\eta}_{1}(G_{c_{t-1}})}.$ (24)
By Lemma 5, we have $c_{t-1}\geq c_{1}$, $|\rho_{t-1}|\leq|\rho|$ and $\sigma_{t-1}^{2}\leq\sigma^{2}_{\infty}$, with $\sigma^{2}_{\infty}\in(0,\infty)$ as in Lemma 5. Therefore, letting
$\bar{\delta}^{\prime}=\min\bigg{\{}\frac{1}{\sigma^{2}_{\infty}},\frac{c_{1}}{2\rho^{2}+c_{1}\sigma^{2}_{\infty}}\bigg{\}}$
it follows that
$c_{t-1}-\frac{\delta^{\prime}\rho_{t-1}^{2}}{1-\delta^{\prime}\sigma_{t-1}^{2}}\geq\frac{c_{t-1}}{2}\geq\frac{c_{1}}{2},\quad\forall\delta^{\prime}\in(0,\bar{\delta}^{\prime}).$ (25)
We now let $\delta^{\prime}\in(0,\bar{\delta^{\prime}})$ be fixed and $f:\mathbb{R}\rightarrow\mathbb{R}$ be such that $C_{f,\delta^{\prime}}<\infty$. Then, noting that $\delta^{\prime}<\bar{\delta}^{\prime}\leq\bar{\delta}_{c_{t-1},\rho_{t-1},\sigma^{2}_{t-1}}$ (with $\bar{\delta}_{a,r,s^{2}}$ defined in (18) for any $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$), by applying (23) with $(a,r,s^{2},\delta)=(c_{t-1},\rho_{t-1},\sigma^{2}_{t-1},\delta^{\prime})$ we have
$\begin{split}|G_{c_{t-1}}(x)M_{\rho_{t-1},\sigma^{2}_{t-1}}(x,f\mathbbm{1}_{[-b,b]^{c}})|&\leq\frac{2C_{f,\delta^{\prime}}}{(1-\delta^{\prime}\sigma_{t-1}^{2})^{1/2}}\bigg{(}1+\frac{\sigma_{t-1}}{(1-\delta^{\prime}\sigma_{t-1}^{2})^{1/2}}\bigg{)}e^{-K_{t-1}b^{2}}\\ &\leq\frac{2C_{f,\delta^{\prime}}}{(1-\delta^{\prime}\sigma_{\infty}^{2})^{1/2}}\bigg{(}1+\frac{\sigma_{\infty}}{(1-\delta^{\prime}\sigma_{\infty}^{2})^{1/2}}\bigg{)}e^{-K_{t-1}b^{2}}\quad\forall x\in\mathbb{R}\end{split}$ (26)
with
$\begin{split}K_{t-1}&=\min\bigg{\{}\frac{1-\delta^{\prime}\sigma_{t-1}^{2}}{6\sigma_{t-1}^{2}},\frac{(1-\delta^{\prime}\sigma_{t-1}^{2})^{2}}{36\rho_{t-1}^{2}}\Big{(}c_{t-1}-\frac{\delta^{\prime}\rho_{t-1}^{2}}{1-\delta^{\prime}\sigma_{t-1}^{2}}\Big{)}\bigg{\}}\\ &\geq\varrho:=\min\bigg{\{}\frac{1-\delta^{\prime}\sigma_{\infty}^{2}}{6\sigma_{\infty}^{2}},\frac{(1-\delta^{\prime}\sigma_{\infty}^{2})^{2}}{36\rho^{2}}\frac{c_{1}}{2}\bigg{\}}>0\end{split}$ (27)
and where the first inequality holds by (25).
On the other hand, by Lemma 5 we have $c_{t-1}\leq\rho^{2}/\sigma^{2}$ and thus
$\hat{\eta}_{1}(G_{c_{t-1}})\geq\hat{\eta}_{1}(G_{\rho^{2}/\sigma^{2}})=\frac{\eta_{1}(G_{c+\rho^{2}/\sigma^{2}})}{\eta_{1}(G_{c})}.$
By combining this latter result with (24), (26) and (27), we obtain that
$|\hat{\eta}_{t}(f\mathbbm{1}_{[-b,b]^{c}})|$ $\leq e^{-\varrho b^{2}}\frac{2C_{f,\delta^{\prime}}}{\eta_{1}(G_{c+\rho^{2}/\sigma^{2}})(1-\delta^{\prime}\sigma_{\infty}^{2})^{1/2}}\bigg{(}1+\frac{\sigma_{\infty}}{(1-\delta^{\prime}\sigma_{\infty}^{2})^{1/2}}\bigg{)},\quad\forall t\geq 2,\quad\forall b\in[2,\infty)$
showing the first part of the lemma.
The prove the second part of the lemma let $f:\mathbb{R}\rightarrow\mathbb{R}$ and note that, for any $b\in(0,\infty)$ we have, using (4) and Lemma 5,
$\eta_{t+1}(f\mathbbm{1}_{[-b,b]^{c}})=\frac{\int_{\mathbb{R}}G_{c_{t-1}}(x)M_{\rho\rho_{t-1},\sigma^{2}+\rho\sigma^{2}_{t-1}}(x,f\mathbbm{1}_{[-b,b]^{c}})\hat{\eta}_{1}(\mathrm{d}x)}{\hat{\eta}_{1}(G_{c_{t-1}})},\quad\forall t\geq 1.$
Then, the second part of the lemma readily follows from the computations used to show the first part of the lemma, and the proof of the lemma is complete. ∎
A.2.4 A general 1-dimensional Koksma-Hlawka inequality
The next result is a general $s=1$ dimensional Koksma-Hlawka inequality, where the integration is w.r.t. an arbitrary distribution $\mu\in\mathcal{P}(\mathbb{R})$. By contrast, in the original Koksma-Hlawka inequality $s\in\mathbb{N}$ is arbitrary but the integration is w.r.t. the $\mathcal{U}(0,1)^{s}$ distribution (see e.g. Dick and Pillichshammer,, 2010, Proposition 2.18, page 33). It is also worth mentioning that Aistleitner and Dick, (2015) derived a version of the Koksma-Hlawka inequality where the integration is w.r.t. an arbitrary probability distribution on $(0,1)^{s}$ for some $s\in\mathbb{N}$.
Lemma 7.
Let $-\infty\leq a<b\leq\infty$ and let $f:\mathbb{R}\rightarrow\mathbb{R}$ be continuously differentiable on the interval $(a,b)$ and such that $V(f\mathbbm{1}_{(a,b)})<\infty$, and let $\pi,\mu\in\mathcal{P}(\mathbb{R})$. Then, with the convention that $f(a)=0$ when $a=-\infty$ and with the convention that $f(b)=0$ when $b=\infty$, we have
$\begin{split}\Big{|}\int_{a}^{b}f(x)\big{(}\pi(\mathrm{d}x)-\mu(\mathrm{d}x)\big{)}\Big{|}\leq\Big{(}|f(a)|+|f(b)|+V(f\mathbbm{1}_{(a,b)})\Big{)}\|\pi-\mu\|.\end{split}$
Proof.
Assume first $(a,b)\in\mathbb{R}^{2}$. Then, using the integration by part formula for Riemann-Stieltjes integrals (see Apostol,, 1973, Theorem 7.6, page 144), we have
$\begin{split}\int_{a}^{b}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}f^{\prime}(x)\mathrm{d}x&=\int_{a}^{b}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}\mathrm{d}f(x)\\ &=f(b)\big{(}F_{\pi}(b)-F_{\mu}(b)\big{)}-f(a)\big{(}F_{\pi}(a)-F_{\mu}(a)\big{)}\\ &-\int_{a}^{b}f(x)\mathrm{d}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}\end{split}$ (28)
and the result of the lemma follows upon noting that $\int_{a}^{b}f(x)\mathrm{d}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}=\int_{a}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))$.
Assume now that $a=-\infty$ and that $b\in\mathbb{R}$, and note that under the assumptions of the lemma we have $\sup_{x\in\mathbb{R}}|f(x)|<\infty$. Then, by (28),
$\bigg{|}\int_{-\infty}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))\bigg{|}$ $=\bigg{|}\lim_{a\rightarrow-\infty}\int_{a}^{b}f(x)\mathrm{d}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}\bigg{|}$
$=\bigg{|}f(b)\big{(}F_{\pi}(b)-F_{\mu}(b)\big{)}-\lim_{a\rightarrow-\infty}\int_{a}^{b}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}f^{\prime}(x)\mathrm{d}x\bigg{|}$
$\leq\|\pi-\mu\|\big{(}|f(b)|+V(\mathbbm{1}_{(-\infty,b)}f)\big{)}$
proving the result of the lemma when $a=-\infty$ and $b\in\mathbb{R}$. The proof of the result of the lemma for the two remaining case (i.e. for $b=\infty$ and $a\in\mathbb{R}\cup\{-\infty\}$) is similar and is therefore omitted to save space. ∎
A.2.5 A discrepancy bound for 1-dimensional importance sampling
From Lemma 7 we readily obtain the following corollary:
Corollary 1.
Let $H\in\mathcal{C}^{1}(\mathbb{R})$ be such that $V(H)<\infty$ and such that $H(x)\geq 0$ for all $x\in\mathbb{R}$, and let $\mu,\pi\in\mathcal{P}(\mathbb{R})$. Then,
$\|\Psi_{H}(\mu)-\Psi_{H}(\pi)\|\leq\frac{\|H\|_{\infty}+2V(H)}{\pi(H)}\|\mu-\pi\|.$
Proof.
The result follows from Lemma 7 and the fact that, for all $B\in\mathcal{B}_{1}$, we have
$|\Psi_{H}(\mu)(B)-\Psi_{H}(\pi)(B)|\leq\frac{|(\mu-\pi)(H\mathbbm{1}_{B})|+|(\mu-\pi)(H)|}{\pi(H)}.$
A.2.6 A discrepancy bound for sampling from univariate mixtures
The following lemma will be used to study jointly the resampling and mutation steps of Algorithm 1. Its proof is inspired by that of Hlawka and Mück, (1972, “Statz 2”).
Lemma 8.
Let $M\in\mathbb{N}$, $\pi_{M}(\mathrm{d}x)=\sum_{m=1}^{M}W_{m}\delta_{\{x_{m}\}}\in\mathcal{P}(\mathbb{R})$ and, for all $x\in\{x_{1},\dots,x_{M}\}$, let $\mu_{x}\in\mathcal{P}(\mathbb{R})$. In addition, let $\mu=\sum_{m=1}^{M}W_{m}\delta_{\{x_{m}\}}\otimes\mu_{x_{m}}\in\mathcal{P}(\mathbb{R}^{2})$ and $\varphi_{\mu}:(0,1)^{2}\rightarrow\mathbb{R}^{2}$ be defined by
$\varphi_{\mu}(u)=\Big{(}x_{1},F^{-1}_{\mu_{x_{1}}}(u_{2})\Big{)},\quad x_{1}=F^{-1}_{\pi_{M}}(u_{1}),\quad u=(u_{1},u_{2})\in(0,1)^{2}.$
Finally, let $N\geq 1$, $\{u^{n}\}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ and assume that for all $a\in\mathbb{R}$ the mapping $\{x_{1},\dots,x_{M}\}\ni x\mapsto F_{\mu_{x}}(a)$ is monotone. Then,
$\sup_{B\in\mathcal{B}_{2}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big{(}\varphi_{\mu}(u^{n})\big{)}-\mu(B)\Big{|}\leq\sqrt{96}\,D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}^{1/2}+12D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}.$ (29)
Proof.
Let $B=(-\infty,a_{1}]\times(-\infty,a_{2}]$ for some $(a_{1},a_{2})\in\mathbb{R}^{2}$. Without loss of generality we assume below that the set $\{x_{m}\}_{m=1}^{M}$ is labelled in such a way that $x_{1}\leq x_{2}\leq\dots\leq x_{M}$. In addition, we assume that $a_{1}\geq x_{1}$ since otherwise the set $B$ is such that
$\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big{(}\varphi_{\mu}(u^{n})\big{)}-\mu(B)\Big{|}=0.$
To prove the lemma we let $G(B)=\big{\{}u\in(0,1)^{2}:\varphi_{\mu}(u)\in B\big{\}}$ so that
$\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big{(}\varphi_{\mu}(u^{n})\big{)}-\mu(B)=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{G(B)}(u^{n})-\lambda_{2}(G(B)).$ (30)
Letting $j_{B}=\max\{i\in\{1,\dots,M\}:x_{i}\leq a_{1}\}$, and using the convention that $W_{i-1}=0$ for $i=1$, we note that
$G(B)=\bigcup_{i=1}^{j_{B}}\Big{(}W_{i-1},W_{i}\Big{]}\times\big{(}0,F_{\mu_{x_{i}}}(b_{2})\big{]}\subseteq(0,1]^{2}.$ (31)
To proceed further we let $L\in\mathbb{N}$ and $\mathsf{P}$ be the partition of $(0,1]^{2}$ in congruent squares of side $\frac{1}{L}$ such that $W\in\mathsf{P}$ if and only if $W=(\xi_{1}/L,(\xi_{1}+1)/L]\times(\xi_{2}/L,(\xi_{2}+1)/L]$ for some $(\xi_{1},\xi_{2})\in\{0,\dots,L-1\}^{2}$. Then, we let
$\mathsf{U}_{1}=\Big{\{}W\in\mathsf{P}:\,\,W\subset\mathring{G(B)}\Big{\}},\quad\mathsf{U}_{2}=\Big{\{}W\in\mathsf{P}:\,\,W\cap\partial(G(B))\neq\emptyset\Big{\}}$
and we let $U_{1}=\cup_{W\in\mathsf{U}_{1}}W$ and $U_{2}=\cup_{W\in\mathsf{U}_{2}}W$ with the convention that $\cup_{W\in\mathsf{S}}=\emptyset$ if $\mathsf{S}=\emptyset$. With this notation in place, and letting $U_{1}^{\prime}=G(B)\setminus U_{1}$, we have
$\begin{split}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{G(B)}(u^{n})-\lambda_{2}(G(B))&=\Big{(}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})\Big{)}+\Big{(}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\Big{)}\end{split}$ (32)
and we start by studying the first term on the r.h.s. of (32).
To do so assume first that $\mathsf{U}_{1}\neq\emptyset$ and, for all $\xi_{1}\in\{0,\dots,L-1\}$, let
$S_{\xi_{1}}=\Big{\{}k\in\{0,\dots,L-1\}:\,\,(\xi_{1}/L,(\xi_{1}+1)/L]\times(k/L,(k+1)/L]\in\mathsf{U}_{1}\Big{\}}$
and let $I=\{\xi_{1}\in\{0,\dots,L-1\}:\,S_{\xi_{1}}\neq\emptyset\}$. Then, using the expression (31) for $G(B)$, it is direct to see that for all $\xi_{1}\in I$ the set
$Q(\xi_{1}):=\bigcup_{k\in S_{\xi_{1}}}(\xi_{1}/L,(\xi_{1}+1)/L]\times(k/L,(k+1)/L]$
belongs to $\mathcal{B}_{2}$ and that we have $\cup_{\xi_{1}\in I}Q(\xi_{1})=U_{1}$. Therefore,
[…]
where the second inequality uses the fact that, since $\{u^{n}\}_{n=1}^{N}$ is assumed to be a point set in $(0,1)^{2}$, we have (see Niederreiter,, 1992, Remarks 2-3 and Proposition 2.4, pages 14-15)
$\sup_{(a,b]\subset(0,1)^{2}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{(a,b]}(u^{n})-\lambda_{2}\big{(}(a,b]\big{)}\Big{|}\leq 4D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}.$ (34)
If $\mathsf{U}_{1}=\emptyset$ we have $\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})=0$ and thus (33) also holds in this case.
We now study the second term on the r.h.s. of (32). To this end note that we can cover $U_{1}^{\prime}$ with sets in $\mathsf{U}_{2}$, and thus
$-\lambda_{2}(U_{2})\leq-\lambda_{2}(U_{1}^{\prime})\leq\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\leq\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{2}}(u^{n})-\lambda_{2}(U_{2})+\lambda_{2}(U_{2}).$
By using this latter result, we obtain that
$\begin{split}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\Big{|}&\leq\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{2}}(u^{n})-\lambda_{2}(U_{2})\Big{|}+\lambda_{2}(U_{2})\leq|\mathsf{U}_{2}|\Big{(}4D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}+L^{-2}\Big{)}\end{split}$ (35)
where the second inequality holds by (34).
Using (31) and the fact that, by assumption, the function $\{x_{1},\dots,x_{M}\}\ni x\mapsto F_{\mu_{x}}(b_{2})$ is either non-increasing or non-decreasing, it is readily checked that $|\mathsf{U}_{2}|\leq 2L$. Together with (30) and (32)-(35), this implies that
$\sup_{B\in\mathcal{B}_{2}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big{(}\varphi_{\mu}(u^{n})\big{)}-\mu(B)\Big{|}\leq 12LD^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}+2L^{-1}$
and the result of the lemma follows by applying this latter inequality with
$L=\Big{\lceil}\big{\{}6D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}\big{\}}^{-1/2}\Big{\rceil}.$
A.2.7 A stability result for the particle system generated by Algorithm 1 based on low discrepancy point sets
Lemma 9.
Let $N\geq 1$ and, for all $t\geq 1$, let $\{u_{t}^{n}=(u_{t,1}^{n},u_{t,2}^{n})\}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ such that, for $i=1,2$, we have $u_{t,i}\neq a2^{-k}$ for all $k\in\{1,\dots,7\}$ and all $a\in\{0,\dots,2^{7}\}$, and such that $D^{*}_{2}(\{u_{t}^{n}\}_{n=1}^{N})<2^{-10}$. Consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}=\{u_{t}^{n}\}_{n=1}^{N}$ for all $t\geq 1$ and let $\xi\in(0,\infty)$ be such that
$\begin{split}\xi>\max\bigg{\{}&|F_{\eta_{1}}^{-1}(1/4)|,|F_{\eta_{1}}^{-1}(3/4)|,\frac{c\sigma^{2}\Phi^{-1}(3/4)^{2}+2\log(256/3)}{2c\sigma\Phi^{-1}(3/4)},\sigma\Phi^{-1}(3/4)+\sigma\Phi^{-1}(7/8)\bigg{\}}.\end{split}$
Then, $\inf_{t\geq 1}\eta_{t}^{N}\big{(}[-\xi,\xi]\big{)}\geq 2^{-8}$ and $\inf_{t\geq 1}\hat{\eta}_{t}^{N}\big{(}[-\xi,\xi]\big{)}\geq 1/4$.
Remark 5.
It can be shown that if, in the statement of the lemma, $\{U_{t}^{n}\}_{n=1}^{N}$ is a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$ then the conclusion of the lemma holds $\mathbb{P}$-a.s. for $N=64$ and all $N\geq 128$.
Proof.
To prove the lemma let $t\geq 1$ and remark first that since $\{u_{t}^{n}\}_{n=1}^{N}$ is a point set in $(0,1)^{2}$ we have (see Niederreiter,, 1992, Remarks 2-3 and Proposition 2.4, pages 14-15)
$\begin{split}\sup_{[a,b]\subset[0,1)^{2}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{[a,b]}(u_{t}^{n})-\lambda_{2}\big{(}[a,b]\big{)}\Big{|}\leq 4D^{*}_{2}(\{u_{t}^{n}\}_{n=1}^{N}).\end{split}$ (36)
Let $[a,b]\subset(0,1)^{2}$ be such that $\lambda_{2}([a,b])\geq 2^{-7}$. Then, using (36), it follows that under the assumptions on $\{u_{t}^{n}\}_{n=1}^{N}$ we have
$\sum_{n=1}^{N}\mathbbm{1}_{[a,b]}(u_{t}^{n})\geq N\Big{(}\lambda_{2}([a,b])-4D^{*}_{2}(\{u_{t}^{n}\}_{n=1}^{N})\Big{)}\geq N\big{(}2^{-7}-2^{-8}\big{)}=N2^{-8}$
showing that, for all $t\geq 1$, every closed interval $[a,b]\subset(0,1)^{2}$ of volume at least equal to $2^{-7}$ contains at least $N2^{-8}$ elements of $P_{t,N}:=\{u_{t}^{n}\}_{n=1}^{N}$.
We now let
[…]
so that proving the lemma amounts to showing that
$\frac{1}{N}\sum_{n=1}^{N}\delta_{[-\xi,\xi]}(X_{t}^{n})\geq 2^{-8},\quad\sum_{n\in I_{t}\cup J_{t}}W_{t}^{n}\geq\frac{1}{4},\quad\forall t\geq 1.$ (37)
To show that (37) holds for $t=1$ remark that $X_{1}^{n}=F_{\eta_{1}}^{-1}(u_{1,1}^{n})$ for all $n\in\{1,\dots,N\}$ and remark that the set $Q:=\cup_{i=0}^{2^{6}-1}\big{[}(2^{5}+i)2^{-7},(2^{5}+i+1)2^{-7}\big{]}$ is such that $\sum_{i=1}^{N}\mathbbm{1}_{Q}(u_{1,1}^{n})=\sum_{i=1}^{N}\mathbbm{1}_{[1/4,3/4]}(u_{1,1}^{n})$ under the assumptions on $\{u_{1}^{n}\}_{n=1}^{N}$. Therefore, it follows from the above computations that the interval $[1/4,3/4]$ contains at least $N/4$ elements of $\{u_{1,1}^{n}\}_{n=1}^{N}$, and thus the set $\{X_{1}^{n}\}_{n=1}^{N}$ has at least $N/4$ elements in the interval $[F_{\eta_{1}}^{-1}(1/4),F_{\eta_{1}}^{-1}(3/4)]\subset[-\xi,\xi]$. Then, by using the definition of $G(\cdot)$, it follows that $\sum_{n\in I_{1}\cup J_{1}}W_{1}^{n}\geq 1/4$, and consequently (37) holds for $t=1$.
Assume now that (37) holds for some $t\geq 1$. Then, we have $\sum_{n\in I_{t}\cup J_{t}}W_{t}^{n}\geq 1/4$ and thus we have $\sum_{n\in I_{t}}W_{t}^{n}\geq 1/8$ and/or $\sum_{n\in J_{t}}W_{t}^{n}\geq 1/8$.
Assume first that $\sum_{n\in I_{t}}W_{t}^{n}\geq 1/8$. Then, using the definition of $G(\cdot)$, it is easily checked that there exists an $i\in\{0,\dots,15\}$ such that $0\leq F^{-1}_{\hat{\eta}_{t}^{N}}(u)\leq\xi$ for all $u\in[i/16,(1+i)/16]$. Therefore, since as proved above the set $[i/16,(1+i)/16]\times[1/8,2/8]$ contains at least $N2^{-8}$ elements of $P_{t+1,N}$, it follows that there exists a set $\hat{I}_{t}\subseteq\{1,\dots,N\}$ such that $|\hat{I}_{t}|\geq N2^{-8}$ and such that, for all $n\in\hat{I}_{n}$,
$\hat{X}_{t}^{n}+\sigma\Phi^{-1}(1/8)\leq X_{t+1}^{n}=\hat{X}_{t}^{n}+\sigma\Phi^{-1}(u_{t+1,2}^{n})\leq\hat{X}_{t}^{n}+\sigma\Phi^{-1}(2/8),\quad 0\leq\hat{X}_{t}^{n}\leq\xi.$ (38)
To proceed further remark that if $(\hat{X}_{t}^{n},X_{t+1}^{n})$ is as in (38) then
$|X_{t+1}^{n}|$ $\leq\max\big{\{}|\xi+\sigma\Phi^{-1}(2/8)|,\sigma|\Phi^{-1}(1/8)|\big{\}}=\max\big{\{}|\xi-\sigma\Phi^{-1}(6/8)|,\sigma\Phi^{-1}(7/8)\big{\}}\leq\xi-\delta$
where the inequality holds with $\delta=\sigma\Phi^{-1}(6/8)$ and uses the condition imposed on $\xi$ in the statement of the lemma. By combining the above calculations, it follows that if $\sum_{n\in I_{t}}W_{t}^{n}\geq 1/8$ then there exists a set $\hat{I}_{t}\subseteq\{1,\dots,N\}$ such that $|\hat{I}_{t}|\geq N2^{-8}$ and such that $|X_{t+1}^{n}|\leq\xi-\delta<\xi$ for all $n\in\hat{I}_{t}$.
Using a similar argument as above, one can easily show that if $\sum_{n\in J_{t}}W_{t}^{n}\geq 1/8$ then there exists a set $\hat{J}_{t}\subseteq\{1,\dots,N\}$ such that $|\hat{J}_{t}|\geq N2^{-8}$ and such that, for all $n\in\hat{J}_{t}$,
$|X_{t+1}^{n}|$ $\leq\max\big{\{}|-\xi+\sigma\Phi^{-1}(6/8)|,\sigma\Phi^{-1}(7/8)\big{\}}=\max\big{\{}|\xi-\sigma\Phi^{-1}(6/8)|,\sigma\Phi^{-1}(7/8)\big{\}}\leq\xi-\delta.$
Therefore, under the inductive hypothesis, we have
$\sum_{n=1}^{N}\delta_{[-\xi+\delta,\xi-\delta]}(X_{t+1}^{n})\geq N2^{-8}$ (39)
and thus
$\begin{split}\sum_{n\not\in(I_{t+1}\cup J_{t+1})}W_{t+1}^{n}&\leq\frac{Ne^{-\frac{c}{2}\xi^{2}}}{Ne^{-\frac{c}{2}\xi^{2}}+N2^{-8}e^{-\frac{c}{2}(\xi-\delta)^{2}}}=\frac{2^{8}}{2^{8}+e^{-\frac{c}{2}\delta^{2}+c\delta\xi}}<\frac{3}{4}\end{split}$ (40)
where the last inequality holds under the condition on $\xi$ imposed in the statement of the lemma. Combining (39) and (40) show that (37) holds at time $t+1$ and the proof of (37), and thus of the lemma, is complete. ∎
A.2.8 A discrepancy bound for Algorithm 1 based on low discrepancy point sets
Lemma 10.
Let $(T,N)\in\mathbb{N}^{2}$ and $\{u_{t}^{n}\}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ for all $t\in\{1,\dots,T\}$, and consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}=\{u_{t}^{n}\}_{n=1}^{N}$ for all $t\in\{1,\dots,T\}$. Assume that there exist a non-empty and bounded interval $I\subset\mathbb{R}$ and a constant $\iota\in(0,1]$ such that
$\hat{\eta}_{t}^{N}(I)\geq\iota,\quad\eta_{t}^{N}(I)\geq\iota,\quad\forall t\in\{1,\dots,T\}.$
Then, there exists a constant $C\in(0,\infty)$, depending only on $c$, $\sigma$, $\rho$, $\iota$, $I$ and $\eta_{1}$, such that, with $\epsilon_{\star}\in(0,1)$ as in Lemma 5,
$\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\inf_{\gamma\in(0,1)}\Big{(}\sqrt{2400}\delta^{1/2}+60\delta\Big{)}^{1-\gamma}\frac{C^{\gamma}}{1-\epsilon_{\star}^{\gamma}},\quad\delta=\sup_{t\in\{1,\dots,T\}}D_{2}^{*}(\{u_{t}^{n}\}_{n=1}^{N}).$
In addition, $\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\sqrt{96}\delta^{1/2}+12\delta+\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.
Proof.
In what follows we let $(c_{t})_{t\geq 0}$, $(\rho_{t})_{t\geq 0}$ and $(\sigma_{t})_{t\geq 0}$ be as defined in Lemma 5, and for all $a\in\mathbb{R}$ we let $f_{a}=\mathbbm{1}_{(-\infty,a]}$ and
$g_{a,k}(x)=\Phi\Big{(}\frac{a-\rho_{k}x}{\sigma_{k}}\Big{)},\quad h_{a,k}(x)=\Phi\Big{(}\frac{a-\rho_{k}x}{\sigma_{k}}\Big{)}-\Phi\Big{(}\frac{a}{\sigma_{k}}\Big{)},\quad\forall x\in\mathbb{R},\quad\forall k\geq 1.$
Next, we let $\phi:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be such that $\phi(\eta)=\Psi_{G}(\eta M)$ for all $\eta\in\mathcal{P}(\mathbb{R})$, and for all $t\geq 1$ we let $\phi^{t}=\phi^{t-1}\phi$ with the convention that $\phi^{t-1}(\eta)=\eta$ for all $\eta\in\mathcal{P}(\mathbb{R})$ when $t=1$.
Using this notation and following the approach introduced by Del Moral, (2004, Chapter 7) for studying PFs we have, using the convention that $\hat{\eta}^{N}_{0}M=\eta_{1}$ (so that $\phi(\hat{\eta}^{N}_{0})=\Psi_{G}(\eta_{1})$),
$\hat{\eta}_{t}^{N}-\hat{\eta}_{t}$ $=\sum_{p=1}^{t}\Big{(}\phi^{t-p}(\hat{\eta}^{N}_{p})-\phi^{t-p}\big{(}\phi(\hat{\eta}^{N}_{p-1})\big{)}\Big{)},\quad\forall t\geq 1$ (41)
and we now study the terms inside the sums.
To this aim assume for now that $T\geq 3$, and let $t\in\{2,\dots,T\}$ and $p\in\{1,\dots,t-1\}$ be fixed. Then, for all $a\in\mathbb{R}$ we have, by Lemma 5,
$\begin{split}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})&=\frac{\hat{\eta}^{N}_{p}Q^{t-p}(f_{a})}{\hat{\eta}^{N}_{p}Q^{t-p}(\mathbb{R})}=\frac{\sum_{n=1}^{N}W_{p}^{n}G_{c_{t-p}}(X_{p}^{n})g_{a,t-p}(X_{p}^{n})}{\sum_{n=1}^{N}W_{p}^{n}G_{c_{t-p}}(X_{p}^{n})}=\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(g_{a,t-p})\end{split}$ (42)
and, similarly,
$\phi^{t-p}\big{(}\phi(\hat{\eta}^{N}_{p-1}))(f_{a})=\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(g_{a,t-p}).$ (43)
Then, by combining (42) and (43), we obtain that
$\begin{split}\sup_{a\in\mathbb{R}}\big{|}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})-&\phi^{t-p}\big{(}\phi_{p}(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\big{|}\\ &=\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(g_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(g_{a,t-p})\Big{|}\\ &=\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(h_{a,t-p})\Big{|}.\end{split}$ (44)
To proceed further remark that, by Lemma 4, we have $\sup_{(a,k)\in\mathbb{R}\times\mathbb{N}}V(h_{a,k})\leq 1$ and thus, by Lemma 7,
$\begin{split}\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}&\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(h_{a,t-p})\Big{|}\leq\Big{\|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}\Big{\|}.\end{split}$ (45)
Noting that
$\|G_{a}\|_{\infty}=1,\quad V(G_{a})=\int_{\mathbb{R}}|G^{\prime}_{a}(x)|\mathrm{d}x=a\int_{\mathbb{R}}|x|G_{a}(x)\mathrm{d}x=2,\quad\forall a\in(0,\infty)$ (46)
it follows from Corollary 1 that
$\big{\|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}\Big{\|}$ $\leq\frac{5\|\hat{\eta}^{N}_{p}-\hat{\eta}^{N}_{p-1}M\|}{\hat{\eta}^{N}_{p-1}M(G_{c+c_{t-p}})}\leq\frac{5\|\hat{\eta}^{N}_{p}-\hat{\eta}^{N}_{p-1}M\|}{\hat{\eta}^{N}_{p-1}M(G_{c+c_{\star}})}$ (47)
where the second inequality uses the fact that, by Lemma 5, we have $\sup_{k\geq 1}c_{k}\leq c_{\star}:=\rho^{2}/\sigma^{2}$. Then, using (47) and noting that under the assumptions of the lemma there exists a constant $\iota^{\prime}\in(0,1]$ such that
$\inf_{k\in\mathbb{N}_{0}}\inf_{x\in I}\hat{\eta}_{k}^{N}M(G_{c+c_{\star}})\geq\iota^{\prime},$ (48)
it follows that
$\begin{split}\big{\|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}\Big{\|}\leq\frac{5\|\eta^{N}_{p-1}-\hat{\eta}^{N}_{p-1}M\|}{\iota^{\prime}}.\end{split}$ (49)
By Lemma 8 and with $\delta$ as in the statement of the lemma, we have
$\|\eta^{N}_{p}-\hat{\eta}^{N}_{p-1}M\|\leq\sqrt{96}\delta^{1/2}+12\delta$ (50)
and thus, by combining (45), (49) and (50), we obtain that
$\begin{split}\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}&(h_{a,t-p})\Big{|}\leq\frac{\sqrt{2400}\delta^{1/2}+60\delta}{\iota^{\prime}}.\end{split}$ (51)
On the other hand, by Lemma 4 we have
$|h_{a,k}(x)|\leq\frac{|\rho_{k}x|}{\sigma_{k}\sqrt{2\pi}},\quad\forall(a,x)\in\mathbb{R}^{2},\quad\forall k\geq 1$
and therefore, noting that $|x|G_{a}(x)\leq a^{-1/2}e^{-1/2}$ for all $x\in\mathbb{R}$ and all $a\in(0,\infty)$, it follows that
$G_{c+c_{t-p}}(x)|h_{a,t-p}(x)|\leq\frac{|\rho_{t-p}|e^{-1/2}}{\sigma_{t-p}\sqrt{(c+c_{t-p})2\pi}}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}},\quad\forall a\in\mathbb{R}$
where the second inequality uses the fact that $\inf_{k\geq 1}\sigma^{2}_{k}=\sigma^{2}_{1}=\sigma^{2}/(1+c\sigma^{2})>0$ while $\inf_{k\geq 1}c_{k}\geq 0$.
Using this result with and recalling that $c_{\star}=\rho^{2}/\sigma^{2}\geq\sup_{k\geq 1}c_{k}$, it follows that we have both
$\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})\Big{|}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\eta_{p}^{N}(G_{c+c_{t-p}})}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\eta_{p}^{N}(G_{c+c_{\star}})}$
and
$\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(h_{a,t-p})\Big{|}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\hat{\eta}^{N}_{p-1}M(G_{c+c_{t-p}})}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\hat{\eta}^{N}_{p-1}M(G_{c+c_{\star}})}$
and thus
$\begin{split}\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(h_{a,t-p})\Big{|}\leq\frac{\bar{C}}{\iota^{\prime}}|\rho_{t-p}|\end{split}$ (52)
with $\bar{C}=\max\{1,2/(c\sigma_{1}^{2})^{1/2}\}$ and assuming without generality that the constant $\iota^{\prime}>0$ such that (48) holds is sufficiently small to ensure that $\big{(}\iota\inf_{x\in I}G_{c+c_{\star}}(x)\big{)}\geq\iota^{\prime}$, with the constant $\iota$ and the interval $I$ as in the statement of the lemma.
Let $\gamma\in(0,1)$ be fixed. Then, by combining (44), (51) and (52) it follows that
$\begin{split}\sup_{a\in\mathbb{R}}\big{|}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})-\phi^{t-p}\big{(}\phi_{p}(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\big{|}&\leq\frac{|\rho_{t-p}|^{\gamma}\bar{C}^{\gamma}\big{(}\sqrt{2400}\delta^{1/2}+60\delta\big{)}^{1-\gamma}}{\iota^{\prime}}.\end{split}$ (53)
The inequality in (53) holds for all $t\in\{2,\dots,T\}$ and all $p\in\{1,\dots,t-1\}$ (assuming that $T\geq 3$), and we now consider the remaining cases where $p=t$ for some $t\in\{1,\dots,T\}$. To do so let $t\in\{1,\dots,T\}$ and $p=t$ and note that
$\sup_{a\in\mathbb{R}}\big{|}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})-\phi^{t-p}\big{(}\phi(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\big{|}$ $=\sup_{a\in\mathbb{R}}\big{|}\hat{\eta}^{N}_{t}(f_{a})-\phi(\hat{\eta}^{N}_{t-1})(f_{a})\big{|}$
$=\sup_{a\in\mathbb{R}}\big{|}\Psi_{G}(\eta^{N}_{t})(f_{a})-\Psi_{G}(\hat{\eta}^{N}_{t-1}M)(f_{a})\big{|}$
$\leq\frac{5\|\eta^{N}_{t}-\hat{\eta}^{N}_{t-1}M\|}{\hat{\eta}^{N}_{t-1}M(G_{c})}$
$\leq\frac{5\|\eta^{N}_{t}-\hat{\eta}^{N}_{t-1}M\|}{\hat{\eta}^{N}_{t-1}M(G_{c+c_{\star}})}$
where the first inequality holds by Corollary 1 and uses (46), and where the last inequality uses the fact that $c_{\star}=\rho^{2}/\sigma^{2}\geq 0$.
Then, using (48) and the fact that, by Lemma 8, we have $\|\eta^{N}_{t}-\hat{\eta}^{N}_{t-1}M\|\leq\sqrt{96}\delta^{1/2}+12\delta$ for all $t\in\{1,\dots,T\}$, it follows that for all $t\in\{1,\dots,T\}$ we have, for $p=t$ and recalling that $\rho_{0}=1$,
$\sup_{a\in\mathbb{R}}\big{|}\phi^{p-t}(\hat{\eta}^{N}_{p})(f_{a})-\phi^{p-t}\big{(}\phi(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\big{|}$ $\leq(\iota^{\prime})^{-1}\min\big{\{}1,\sqrt{2400}\delta^{1/2}+60\delta\big{\}}$
$\leq(\iota^{\prime})^{-1}\min\big{\{}1,\big{(}\sqrt{2400}\delta^{1/2}+60\delta\big{)}^{1-\gamma}\big{\}}$
$\leq(\iota^{\prime})^{-1}|\rho_{t-p}|^{\gamma}\bar{C}^{\gamma}\big{(}\sqrt{2400}\delta^{1/2}+60\delta\big{)}^{1-\gamma}$
showing that (53) holds for any integers $1\leq p\leq t\leq T$.
We now let $C_{\star}\in(0,\infty)$ and $\epsilon_{\star}\in(0,1)$ be as in Lemma 5, so that $|\rho_{k}|\leq C_{\star}\epsilon_{\star}^{k}$ for all $k\geq 0$. Then, by using (41) and (53), it follows that
$\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|=\sup_{t\in\{1,\dots,T\}}\sup_{a\in\mathbb{R}}\big{|}\hat{\eta}_{t}^{N}(f_{a})-\hat{\eta}_{t}(f_{a})\big{|}$ $\leq\sum_{p=1}^{T}\sup_{a\in\mathbb{R}}\Big{|}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})-\phi^{t-p}\big{(}\phi(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\Big{|}$
$\leq(\iota^{\prime})^{-1}\big{(}\sqrt{2400}\delta^{1/2}+60\delta\big{)}^{1-\gamma}\frac{(C_{\star}\bar{C})^{\gamma}}{1-\epsilon_{\star}^{\gamma}}$
and since $\gamma\in(0,1)$ is arbitrary the first part of the lemma follows.
To show the second part of the lemma let $t\in\{2,\dots,T\}$ (assuming that $T\geq 2$) and note that
$\|\eta_{t}^{N}-\eta_{t}\|$ $\leq\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|+\|\hat{\eta}^{N}_{t-1}M-\hat{\eta}_{t-1}M\|\leq\sqrt{96}\delta^{1/2}+12\delta+\|\hat{\eta}^{N}_{t-1}M-\hat{\eta}_{t-1}M\|$ (54)
where the second inequality holds by Lemma 8. Noting that for all $a\in\mathbb{R}$ we have
$M(f_{a})(x)=\Phi\bigg{(}\frac{a-\rho x}{\sigma}\bigg{)}=:\tilde{g}_{a}(x),\quad\forall x\in\mathbb{R}$
where, using Lemma 4, $V(\tilde{g}_{a})=V(h_{a,\rho,\sigma})\leq 1$, it follows from Lemma 7 that (assuming that $T\geq 2$)
$\|\hat{\eta}^{N}_{t-1}M-\hat{\eta}_{t-1}M\|=\sup_{a\in\mathbb{R}}|\hat{\eta}^{N}_{t-1}(\tilde{g}_{a})-\hat{\eta}_{t-1}(\tilde{g}_{a})|\leq\|\hat{\eta}^{N}_{t-1}-\hat{\eta}_{t-1}\|,\quad\forall t\in\{2,\dots,T\}.$
By combining this latter result with (54) we obtain that
$\sup_{t\in\{2,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\sqrt{96}\delta^{1/2}+12\delta+\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}^{N}_{t}-\hat{\eta}_{t}\|.$
Then, the second part of the lemma follows upon noting that $\|\eta_{1}^{N}-\eta_{1}\|\leq\sqrt{96}\delta^{1/2}+12\delta$ by Lemma 8. The proof of the lemma is complete. ∎
A.3 Proof of Lemmas 1-3, Theorems 1-2 and Propositions 1-3
Lemma 1 is a direct consequence of Lemmas 9-10. Lemma 2 is a direct consequence of Lemmas 1, 6 and 7. Theorem 2 directly follows from (9) and Lemma 1. Proposition 2 is a direct consequence of Lemma 1 and Theorem 1. Consequently, below we only prove Lemma 3, Theorem 1 and Propositions 1- 3.
A.3.1 Proof of Proposition 1
Proof.
The result of the proposition trivially holds if $\rho=0$ and therefore below we assume that $\rho\neq 0$.
Let $(c_{t})_{t\geq 0}$, $(\rho_{t})_{t\geq 0}$, $(\sigma_{t})_{t\geq 0}$, $C_{\star}\in(0,\infty)$ and $\epsilon_{\star}\in(0,1)$ be as in Lemma 5, and for all $a\in\mathbb{R}$ let $f_{a}=\mathbbm{1}_{(-\infty,a]}$ and, for all $r\in\mathbb{R}$ and $s\in(0,\infty)$, let $h_{a,r,s}$ be as defined in (11) and let $p_{a,s}=\Phi(a/s)$.
We now let $a\in\mathbb{R}$ be fixed. Then, using (4) and Lemma 5, we have
$\hat{\eta}_{t+1}(f_{a})=\frac{\hat{\eta}_{1}Q^{t}(f_{a})}{\hat{\eta}_{1}Q^{t}(\mathbb{R})}=\Psi_{G_{c_{t}}}(\hat{\eta}_{1})\big{(}M_{\rho_{t},\sigma_{t}^{2}}(f_{a})\big{)}=\Psi_{G_{c_{t}}}(\hat{\eta}_{1})(h_{a,\rho_{t},\sigma_{t}})+p_{a,\sigma_{t}},\quad\forall t\geq 1.$ (55)
Using Lemma 4 and the fact that $\inf_{t\geq 1}\sigma_{t}\geq\sigma_{1}>0$, it follows that
$|h_{a,\rho_{t},\sigma_{t}}(x)|$ $\leq\frac{|x\rho_{t}|}{\sigma_{1}\sqrt{2\pi}},\quad\forall x\in\mathbb{R},\quad\forall t\geq 1.$
By combining this latter result and (55), and by using the fact that $\sup_{x\in\mathbb{R}}|x|G_{q}(x)\leq q^{-1/2}e^{-1/2}$ for all $q\in(0,\infty)$, it follows that
$\big{|}\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{t}}\big{|}\leq\big{|}\Psi_{G_{c_{t}}}(\hat{\eta}_{1})(h_{a,\rho_{t},\sigma_{t}})\big{|}\leq\frac{|\rho_{t}|e^{-1/2}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{2\pi c_{t}}}\leq\frac{|\rho_{t}|\sqrt{1+c\sigma^{2}}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{c\rho^{2}}},\quad\forall t\geq 1$ (56)
where the last inequality uses Lemma 5. By Lemma 5 we have $G_{c_{t}}(x)\geq G_{\rho^{2}/\sigma^{2}}(x)$ for all $x\in\mathbb{R}$ and $|\rho^{t}|\leq C_{\star}\epsilon_{\star}^{t}$ for all $t\geq 1$, and thus, by (56),
$\big{|}\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{t}}\big{|}\leq C_{1}|\rho_{t}|\leq C_{1}\epsilon_{\star}^{t},\quad\forall t\geq 1$ (57)
with
$C_{1}=\frac{C_{\star}\sqrt{1+c\sigma^{2}}}{\hat{\eta}_{1}(G_{\rho^{2}/\sigma^{2}})\sigma_{1}\sqrt{c\rho^{2}}}.$
On the other hand, using the mean value theorem, for all $t\geq 1$ there exists a $v_{t,a}\in([-1,1])$ such that
$|p_{a,\sigma_{t}}-p_{a,\sigma_{\infty}}|$ $\leq\frac{|\sigma_{t}-\sigma_{\infty}|}{\sigma_{t}+v_{t,a}(\sigma_{t}-\sigma_{\infty})}\,\,\frac{|a|}{\sigma_{t}+v_{t,a}(\sigma_{t}-\sigma_{\infty})}\varphi\Big{(}\frac{a}{\sigma_{t}+v_{t,a}(\sigma_{t}-\sigma_{\infty})}\Big{)}.$
Together with the fact that $|x|\varphi(x)\leq 1$ for all $x\in\mathbb{R}$, and recalling $\inf_{t\geq 1}\sigma_{t}\geq\sigma_{1}>0$, this shows that
$|p_{a,\sigma_{t}}-p_{a,\sigma_{\infty}}|\leq\frac{|\sigma_{t}-\sigma_{\infty}|}{\sigma_{1}},\quad\forall t\geq 1.$ (58)
By Lemma 5, we have $|\sigma^{2}_{t}-\sigma^{2}_{\infty}|\leq C^{2}_{\star}\epsilon_{\star}^{2t}$ for all $t\geq 1$ and thus, using the fact that $xy\leq(x^{2}+y^{2})/2$ for all $(x,y)\in\mathbb{R}^{2}$, we have
$|\sigma_{t}-\sigma_{\infty}|^{2}\leq|\sigma^{2}_{t}-\sigma^{2}_{\infty}|\leq C^{2}_{\star}\epsilon_{\star}^{2t},\quad\forall t\geq 1.$ (59)
Noting that in the above calculations $a\in\mathbb{R}$ is arbitrary, by combining (57), (58) and (59) we obtain that
$\|\hat{\eta}_{t+1}-\mathcal{N}_{1}(0,\sigma^{2}_{\infty})\|=\sup_{a\in\mathbb{R}}\big{|}\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{\infty}}\big{|}\leq\big{(}C_{1}+C_{\star}/\sigma_{1}\big{)}\epsilon_{\star}^{t},\quad\forall t\geq 1.$
To complete the proof of the proposition it remains to show that
$\|\eta_{t+1}-\mathcal{N}_{1}(0,\rho\sigma^{2}_{\infty}+\sigma^{2})\|\leq\|\eta_{t}-\mathcal{N}_{1}(0,\sigma^{2}_{\infty})\|,\quad\forall t\geq 1.$ (60)
To this aim we let $\hat{\eta}_{\infty}=\mathcal{N}_{1}(0,\sigma^{2}_{\infty})$ and note that, for all $t\geq 2$, we have
$\|\eta_{t}-\mathcal{N}_{1}(0,\rho\sigma^{2}_{\infty}+\sigma^{2})\|$ $=\|\hat{\eta}_{t-1}M-\hat{\eta}_{\infty}M\|=\sup_{a\in\mathbb{R}}|\hat{\eta}_{t-1}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|.$
Since $\sup_{a\in\mathbb{R}}V(h_{a,\rho,\sigma})\leq 1$ by Lemma 4, it follows from Lemma 7 that
$\sup_{a\in\mathbb{R}}|\hat{\eta}_{t-1}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|\leq\|\hat{\eta}_{t-1}-\hat{\eta}_{\infty}\|$
and (60) follows. The proof of the proposition is complete. ∎
A.3.2 Proof of Lemma 3
Proof.
Let $\delta\in(0,1)$ and remark that, since $\{U_{1}^{n}\}_{n=1}^{N}$ is a set if $N$ independent $\mathcal{U}(0,1)^{s}$ random numbers, it follows from Aistleitner and Hofer, (2014, Theorem 1) that for all $\gamma\in(0,1)$ we have
$\mathbb{P}\Big{(}D^{*}_{s}\big{(}\{U_{1}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}\geq 1-\gamma,\quad\forall N\geq N_{s,\delta,\gamma}:=\frac{160s+33\log(1/\gamma)}{\delta^{2}}.$
Therefore, since ($(\{U_{t}^{n}\}_{n=1}^{N})_{t\geq 1}$ is assumed to be a sequence of independent sets, it follows that for all $\gamma\in(0,1)$ we have
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}D^{*}_{s}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}$ $=\mathbb{P}\Big{(}D^{*}_{s}\big{(}\{U_{1}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}^{T}\geq(1-\gamma)^{T},\quad\forall N\geq N_{s,\delta,\gamma}.$
We now let $q\in(0,1)$ so that, by applying the latter result with $\gamma=1-(1-q)^{1/T}$, we have
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}D^{*}_{s}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}\geq 1-q,\quad\forall N\geq N_{s,\delta,1-(1-q)^{1/T}}.$
Using Bernoulli’s inequality (see e.g. Li and Yeh,, 2013), we have $1-(1-q)^{1/T}\geq T^{-1}q$ and thus
$N_{s,\delta,1-(1-q)^{1/T}}\leq\frac{160s-33\log(T^{-1}q)}{\delta^{2}}=\frac{160s+33\log(T)+33\log(1/q)}{\delta^{2}}$
and the proof of the lemma is complete. ∎
A.3.3 Proof of Theorem 1
Proof.
We start with two preliminary results. To state the first one let $\pi_{M}(\mathrm{d}x)=\sum_{m=1}^{M}W_{m}\delta_{\{x_{m}\}}\in\mathcal{P}(\mathbb{R})$ for some $M\in\mathbb{N}$, for all $x\in\{x_{1},\dots,x_{M}\}$ let $\mu_{x}\in\mathcal{P}(\mathbb{R})$, let $\mu=\sum_{m=1}^{M}W_{m}\mu_{x_{m}}\in\mathcal{P}(\mathbb{R})$, $\{u^{n}\}_{n=1}^{N}$ be a point set in $(0,1)$ and $\mu^{N}=N^{-1}\sum_{n=1}^{N}\delta_{F^{-1}_{\mu}(u^{n})}$. Then, it is direct to see that $\|\mu^{N}-\mu\|=D^{*}_{1}(\{u^{n}\}_{n=1}^{N})$. To state the second preliminary result let ($(\{U_{t}^{n}\}_{n=1}^{N})_{t\geq 1}$ be as in the statement of the theorem, and let $\mu_{1}^{N}=\eta_{1}^{N}$ and
$\mu_{t+1}^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{F^{-1}_{\hat{\mu}_{t}^{N}M}(U_{t+1,1}^{n})},\quad\hat{\mu}^{N}_{t}=\Psi_{G}(\mu_{t}^{N}),\quad\forall t\geq 1.$
Then, it is direct to see that we have $(\hat{\eta}_{t}^{N})_{t\geq 1}\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{dist}}}}}{{=}}(\hat{\mu}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{dist}}}}}{{=}}(\mu_{t}^{N})_{t\geq 1}$.
Using these two results, it follows that to prove the theorem we can use the version of Lemma 1 we obtain if, in Lemma 8, we replace the right-hand size of (29) by $2D_{1}^{*}(\{u_{1}^{n}\}_{n=1}^{N})$, that is, to prove the lemma we can do as if the conclusion of Lemma 1 holds with $\delta_{N,T}=\sup_{t\in\{1,\dots,T\}}D_{1}^{*}(\{U_{t,1}^{n}\}_{n=1}^{N})^{2}$. Using this observation the result of the theorem then follows from Lemma 3. ∎
A.3.4 Proof of Proposition 3
Proof.
Let $\kappa\in(0,1)$ and let $T_{\kappa}\in\mathbb{N}$ and $a_{\kappa}\in(0,\infty)$ be such that we have both $\eta_{t}([-a_{\kappa},a_{\kappa}])\leq\kappa$ and $\hat{\eta}_{t}([-a_{\kappa},a_{\kappa}])\leq\kappa$ for all $T\geq T_{\kappa}$. Remark that such constants $T_{\kappa}$ and $a_{\kappa}$ exist by Proposition 1.
Next, let $\varrho_{\kappa}=\min\big{\{}1-\Phi(4a_{\kappa}/\sigma),1/2\big{\}}$ so that, using (5) and with $(\mathcal{F}^{N}_{t})_{t\geq 1}$ as defined in Section 3.1, we have
$\mathbb{P}\Big{(}\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}^{N}_{t-1}\Big{)}\geq\varrho^{N}_{\kappa},\quad\forall t\geq 1.$ (61)
Noting that for all $t\geq T_{\kappa}$ we have
$1-\mathbb{P}\Big{(}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa|\mathcal{F}_{t-1}^{N}\Big{)}\geq\mathbb{P}\Big{(}\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}_{t-1}^{N}\Big{)}$
it follows from (61) that
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa\Big{)}\leq(1-\varrho^{N}_{\kappa})^{T-T_{\kappa}+1}.$
To complete the proof of the proposition let $q\in(0,1)$ and note that
$(1-\varrho^{N}_{\kappa})^{T-T_{\kappa}+1}<1-q,\quad\forall N<\frac{\log\big{(}(1-(1-q)^{\frac{1}{T-T_{\kappa}+1}})^{-1}\big{)}}{\log(1/\varrho_{\kappa})}\leq\frac{\log(T-T_{\kappa}+1)-\log(q)}{\log(1/\varrho_{\kappa})}$
where the inequality uses the fact that, by Bernoulli’s inequality (see e.g. Li and Yeh,, 2013), we have $(1-q)^{1/(T-T_{\kappa}+1)}\leq 1-q/(T-T_{\kappa}+1)$. The result of the proposition follows. ∎

### v4

@@
 This is particularly important for applications which require to accurately estimate a very long sequence of filtering distributions $\{\hat{\eta}_{t}\}_{t=1}^{T}$; for instance, according to Kalra and Paddock, (2016), in order to demonstrate the reliability of self-driving cars their onboard algorithms must be tested for hundreds of billions of miles.
-The second contribution of this paper is to provide a first theoretical analysis of the probability $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$.
-Considering the same filtering problem as the one described above, we prove that there exists a finite constant $C_{\kappa}$ such that for any $q\in(0,1)$ and time horizon $T\geq 1$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq C_{\kappa}(\log(T)+\log(1/q))$.
-Consequently, to guarantee that we still have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ as $T$ increases, the number of particles only needs to grow at speed $\log(T)$ while, to decrease $q$, the number of particles should only grow at speed $\log(1/q)$.
-To the best of our knowledge Del Moral, (2013) provides the only previous attempt to study the ability of PFs to jointly estimate multiple filtering distributions.
-In particular, in Del Moral, (2013, Corollary 14.3.3, page 435) it is shown that, for some finite constant $C_{\kappa}$, for all $q\in(0,1)$ and $T\in\mathbb{N}$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\geq\kappa)\leq q$ for all $N\geq C_{\kappa}\log(T)\log(1/q)$ and for any function $f$ with oscillations bounded by one.
-This result is however obtained under strong assumptions on the SSM (which are not verified for the filtering problem that we consider in this work) and it only provides a theoretical guarantee for estimating the filtering expectations $\{\hat{\eta}_{t}(f)\}_{t=1}^{T}$ of a single function.
-Despite these reassuring results concerning the time evolution of PF estimates, it is not entirely satisfactory to delegate (life-critical) filtering tasks to algorithms that we know will fail infinitely often.
+The second contribution of this paper is to provide a first sharp result on the dependence of the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ to $N$ and $T$.
+Considering the same filtering problem as the one described above, we prove that there exists a finite constant $C_{\kappa}$ such that for any $q\in(0,1)$ and time horizon $T\geq 1$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq N_{\kappa,T,q}:=C_{\kappa}(\log(T)+\log(1/q))$.
+For a given $\kappa$, the dependence of $N_{\kappa,T,q}$ to $T$ and $q$ is shown to be sharp, and thus to guarantee that we still have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ as $T$ increases the number of particles only needs to grow at speed $\log(T)$ while, to decrease $q$, the number of particles should only grow at speed $\log(1/q)$.
+It is worth noting at this stage that Corollary 14.5.7 in Del Moral, (2013) establishes that $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq N^{\prime}_{\kappa,T,q}:=C^{\prime}_{\kappa}(\log(T+e)(1+\log(1/q))$ and for some finite constant $C^{\prime}_{\kappa}$.
+We also note that it has not been realized (and therefore proved) in this reference that $N$ must increase with $T$ to ensure that $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\geq\kappa)$ remains bounded by $q$ as $T$ increases (and thus that for any fixed $N\geq 1$ we have $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$).
+We finally remark that this result is obtained under (strong) assumptions which are not verified for the filtering problem that we consider in this work.
+Despite the reassuring results we proved regarding the time evolution of PF estimates, it is not entirely satisfactory to delegate (life-critical) filtering tasks to algorithms that we know will fail infinitely often.
 Following a previous discussion, to resolve this limitation of PFs some de-randomization is needed, and a possible approach to do so is to replace its underpinning independent $\mathcal{U}(0,1)$ random variables by randomized quasi-Monte (RQMC) point sets.
@@
 Given the popularity of PF algorithms in many applications it is important to understand the behaviour of the resulting joint estimation error, $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.
-Surprisingly, with the exception of Del Moral, (2013, Corollary 14.3.3, page 435), this latter quantity (or a variant of it) does not seem to have been previously studied in the literature.
+Surprisingly, with the exception of Del Moral, (2013, Corollary 14.5.7, page 449), this latter quantity (or a variant of it) does not seem to have been previously studied in the literature.
 The results we obtained in this paper are reassuring, but an important question that has yet to be studied is how the dimension of the filtering problem impacts the behaviour of $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.

### v5

@@
-supReferences
+Safety of particle filters: Some results on the time evolution of particle filter estimates
 Abstract
-[…]
+Particle filters (PFs) form a class of Monte Carlo algorithms that propagate over time a set of $N\geq 1$ particles which can be used to estimate, in an online fashion, the sequence of filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$ defined by a state-space model.
+Despite the popularity of PFs, the study of the time evolution of their estimates has received barely any attention in the literature.
+Denoting by $(\hat{\eta}_{t}^{N})_{t\geq 1}$ the PF estimate of $(\hat{\eta}_{t})_{t\geq 1}$ and letting $\kappa\in(0,1/2)$, in this work we first show that for any number of particles $N$ it holds that, with probability one, we have $\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa$ for infinitely many time instants $t\geq 1$, with $\|\cdot\|$ the Kolmogorov distance between probability distributions.
+Considering a simple filtering problem we then provide reassuring results concerning the ability of PFs to estimate jointly a finite set $\{\hat{\eta}_{t}\}_{t=1}^{T}$ of filtering distributions by studying the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$.
+Finally, on the same toy filtering problem, we prove that sequential quasi-Monte Carlo, a randomized quasi-Monte Carlo version of PF algorithms, offers greater safety guarantees than PFs in the sense that, for this algorithm, it holds that $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|=0$ with probability one.
 Mathieu Gerber
@@
 Particle filters (PFs) are used for real-time inference in state-space models (SSMs) in various applications, including robotics (Thrun,, 2002; Hailu et al.,, 2024), self-driving cars (Berntorp and Di Cairano,, 2019; Hafez et al.,, 2024), assisted surgery (Kummert et al.,, 2021) and ballistics objects tracking (Kim et al.,, 2023).
-Denoting by $t\in\mathbb{N}$ the time index and by $\hat{\eta}_{t}$ the conditional distribution of the state variable at time $t$ given the available observations $\{y_{s}\}_{s=1}^{t}$, a PF is a Monte Carlo algorithm that propagates over time a set of $N\in\mathbb{N}$ particles that can be used to estimate, in an online fashion, the sequence of the so-called filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$.
+Denoting by $t\in\mathbb{N}:=\{1,2,\dots\}$ the time index and by $\hat{\eta}_{t}$ the conditional distribution of the state variable at time $t$ given the available observations $\{y_{s}\}_{s=1}^{t}$, a PF is a Monte Carlo algorithm that propagates over time a set of $N\in\mathbb{N}$ particles that can be used to estimate, in an online fashion, the sequence of the so-called filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$.
 For instance, in self-driving cars applications $\hat{\eta}_{t}$ is the conditional distribution of the position of the car at time $t$ given its GPS localisation (known to have a precision of a few meters) and its distance with respect to some nearby landmarks, under the assumed SSM.
 The popularity of PFs stems from the fact that this class of algorithms can be deployed on a very large class of SSMs.
-In addition, it is widely believed that the quality of the PF estimate $\hat{\eta}_{t}^{N}$ of $\hat{\eta}_{t}$ does not deteriorate over time, a belief supported by several theoretical analyses showing that, for some class of functions $\mathfrak{F}$ and $p\in\mathbb{N}$, we have (see Caffarel et al.,, 2025, and references therein)
+In addition, it is widely believed that the quality of the PF estimate $\hat{\eta}_{t}^{N}$ of $\hat{\eta}_{t}$ does not deteriorate over time, a belief supported by several theoretical analyses showing that, for some class of functions $\mathfrak{F}$ and constant $p\in\mathbb{N}$, we have (see Caffarel et al.,, 2025, and references therein)
 $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\mathbb{E}\big[|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|^{p}\big]=0,\quad\forall f\in\mathfrak{F}.$ (1)
@@
 1.2 Contributions of the paper
-[…]
-For a given $\kappa\in(0,1)$ the PF estimate $(\hat{\eta}_{t}^{N})_{t\geq 1}$ of $(\hat{\eta}_{t})_{n\geq 1}$ is therefore necessarily such that we have $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$.
-Given the ever increasing use of PFs in life critical applications, notably due to the increasing prevalence of autonomous systems (Ogunsina et al.,, 2024) and the role that PFs play in these systems, it is then crucial to understand the behaviour of $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ as a function of the number of particle $N$ and of the time horizon $T$.
-This is particularly important for applications which require to accurately estimate a very long sequence of filtering distributions $\{\hat{\eta}_{t}\}_{t=1}^{T}$; for instance, according to Kalra and Paddock, (2016), in order to demonstrate the reliability of self-driving cars their onboard algorithms must be tested for hundreds of billions of miles.
+In this paper we first show that such a theoretical guarantee for PFs cannot exist.
+Formally, denoting by $\|\cdot\|$ the Kolmogorov distance between probability distributions, we prove that for any $\kappa\in(0,1/2)$, with probability one and for any number of particles $N\geq 1$, we have $\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa$ for infinitely many time instants $t\geq 1$.
+The fact that in some challenging situations a PF may fail to accurately estimate some filtering distributions is not surprising: it is well-known that this problem will usually happen when the distance between two successive filtering distributions $\hat{\eta}_{t}$ and $\hat{\eta}_{t+1}$ is large, e.g.
+because the SSM is not time homogenous or because the observations $y_{t}$ and $y_{t+1}$ are very different from each other.
+However, and crucially, we prove the above property of PFs on what is arguably the simplest filtering scenario, namely for a one-dimensional and time-homogenous linear Gaussian SSM with observations $y_{t}=0$ for all $t\geq 1$.
+As we will see in what follows, this negative result for PF estimates is trivial to prove, and arises from the very simple and intuitive reason that too much randomness is used to propagate the particle system over time.
+We stress that this problem is not limited to PFs but is common and intrinsic to any sequential (importance) sampling algorithms, including the ensemble Kalman filter (Roth et al.,, 2017).
+For any given $\kappa\in(0,1)$, the PF estimate $(\hat{\eta}_{t}^{N})_{t\geq 1}$ of the sequence $(\hat{\eta}_{t})_{n\geq 1}$ is therefore necessarily such that $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$.
+Given the ever-increasing use of PFs in life critical applications, notably due to the increasing prevalence of autonomous systems (Ogunsina et al.,, 2024) and the role that PFs play in these systems, it is then crucial to understand the behaviour of the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ as a function of the number of particles $N$ and of the time horizon $T$.
 The second contribution of this paper is to provide a first sharp result on the dependence of the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ to $N$ and $T$.
-Considering the same filtering problem as the one described above, we prove that there exists a finite constant $C_{\kappa}$ such that for any $q\in(0,1)$ and time horizon $T\geq 1$ we have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq N_{\kappa,T,q}:=C_{\kappa}(\log(T)+\log(1/q))$.
-For a given $\kappa$, the dependence of $N_{\kappa,T,q}$ to $T$ and $q$ is shown to be sharp, and thus to guarantee that we still have $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ as $T$ increases the number of particles only needs to grow at speed $\log(T)$ while, to decrease $q$, the number of particles should only grow at speed $\log(1/q)$.
-It is worth noting at this stage that Corollary 14.5.7 in Del Moral, (2013) establishes that $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ for all $N\geq N^{\prime}_{\kappa,T,q}:=C^{\prime}_{\kappa}(\log(T+e)(1+\log(1/q))$ and for some finite constant $C^{\prime}_{\kappa}$.
-We also note that it has not been realized (and therefore proved) in this reference that $N$ must increase with $T$ to ensure that $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\geq\kappa)$ remains bounded by $q$ as $T$ increases (and thus that for any fixed $N\geq 1$ we have $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$).
-We finally remark that this result is obtained under (strong) assumptions which are not verified for the filtering problem that we consider in this work.
+Considering the same filtering problem as the one described above, we prove that there exists a constant $C\in(0,\infty)$ such that, for any $(\kappa,q)\in(0,1)^{2}$ and any $T\geq 1$, the inequality $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ holds for all $N\geq N_{T,\kappa,q}:=Cv_{\kappa}\log(T/q)$, with $v_{\kappa}=\{1+\log(1+\kappa^{-1})\}^{2}\kappa^{-2}$.
+We also establish that, for a given $\kappa$, the dependence of $N_{T,\kappa,q}$ on the two parameters $T$ and $q$ is sharp.
+To the best of our knowledge, prior to this work the only study of the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)$ was given in Del Moral, (2013, Corollary 14.5.7, page 449).
+In this reference, it is shown that the inequality $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)\leq q$ holds for all $N\geq N^{\prime}_{T,\kappa,q}:=C^{\prime}\kappa^{-2}\log(1+T)(1+\log(1/q))$, with $C^{\prime}$ a finite constant.
+This result is however obtained under strong assumptions, which are for instance not satisfied for the filtering problem that we consider in this work.
+We also note that $N^{\prime}_{T,\kappa,q}$ has worse dependence on the parameters $(T,q)$ than $N_{T,\kappa,q}$, although its dependence on $\kappa$ is slightly better.
+Lastly, and importantly, it has not been realized (and therefore proved) in Del Moral, (2013) that $N$ must increase with $T$ to ensure that the probability $\mathbb{P}(\sup_{t\in\{1,\dots,T\}}|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|\geq\kappa)$ remains bounded above by $q$ as $T$ increases, and thus that for any fixed $N\geq 1$ we have $\mathbb{P}(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa)=1$.
 Despite the reassuring results we proved regarding the time evolution of PF estimates, it is not entirely satisfactory to delegate (life-critical) filtering tasks to algorithms that we know will fail infinitely often.
-Following a previous discussion, to resolve this limitation of PFs some de-randomization is needed, and a possible approach to do so is to replace its underpinning independent $\mathcal{U}(0,1)$ random variables by randomized quasi-Monte (RQMC) point sets.
+Following a previous discussion, to resolve this limitation of PFs some de-randomization is needed, and a possible approach to do so is to replace its underpinning independent $\mathcal{U}(0,1)$ random variables by randomized quasi-Monte point sets.
 This idea has been proposed and studied in Gerber and Chopin, (2015) with the aim of obtaining a filtering algorithm, called sequential quasi-Monte Carlo (SQMC), which has a faster convergence rate (as $N\rightarrow\infty$) than PFs.
-Letting $(\tilde{\eta}^{N}_{t})_{t\geq 1}$ denote the estimate of $(\hat{\eta}_{t})_{t\geq 1}$ obtained with SQMC, the third contribution of this paper is to show, for the one-dimensional filtering problem outlined above, $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\tilde{\eta}^{N}_{t}-\hat{\eta}_{t}\|=0$ with probability one.
-To analyse the time evolution of PF and SQMC estimates we introduce a new approach for studying PF algorithms which builds on tools developed in the PF and in the quasi-Monte Carlo literature.
-Informally speaking, in a first step we analyse the behaviour of the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ for a given realization of the $\mathcal{U}(0,1)$ random numbers underpinning the algorithm.
-Then, in a second step, we leverage the known properties of sets of independent $\mathcal{U}(0,1)$ random numbers and of RQMC point sets to apply the result obtained in the first step to PFs and SQMC.
+Letting $(\tilde{\eta}^{N}_{t})_{t\geq 1}$ denote the estimate of $(\hat{\eta}_{t})_{t\geq 1}$ obtained with SQMC, the third contribution of this paper is to show, for the one-dimensional filtering problem outlined above, that we have $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\tilde{\eta}^{N}_{t}-\hat{\eta}_{t}\|=0$ with probability one.
 1.3 Notation and organization of the paper
-In what follows all the random variables are defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$.
-To introduce some notation let $s\in\mathbb{N}$.
-Then, we let $\mathcal{B}(\mathbb{R}^{s})$ denote the Borel $\sigma$-algebra on $\mathbb{R}^{s}$, $\mathcal{P}(\mathbb{R}^{s})$ denote the set of probability measures on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$, $\lambda_{s}$ denote the Lebesgue measure on $\mathbb{R}^{s}$ and
-$\mathcal{B}_{s}=\Big{\{}B\subset\mathbb{R}^{s}:B=\prod_{i=1}^{s}(-\infty,a_{i}],\,(a_{1},\dots a_{s})\in\mathbb{R}^{s}\Big{\}}.$
-In addition, for any signed-measure $\mu$ on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$ we let $\|\mu\|=\sup_{B\in\mathcal{B}_{s}}|\mu(B)|$ and, for a point set $\{u^{n}\}_{n=1}^{N}$ in $[0,1)^{s}$, we denote by $D^{*}_{s}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}$ its star discrepancy, that is
-$D^{*}_{s}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}:=\sup_{b\in[0,1)^{s}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{[0,b)}(u^{n})-\lambda_{s}\big{(}[0,b)\big{)}\Big{|}.$
-Next, for a probability distribution $\pi\in\mathcal{P}(\mathbb{R}^{s})$, a measurable function $f:\mathbb{R}^{s}\rightarrow\mathbb{R}$ and a (potentially un-normalized) kernel $K$ acting on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$, we let $\pi(f)=\int_{\mathbb{R}^{s}}f(x)\pi(\mathrm{d}x)$, $\pi K$ denote the measure such that $\pi K(A)=\int_{\mathbb{R}^{s}}K(x,A)\pi(\mathrm{d}x)$ for all $A\in\mathcal{B}(\mathbb{R}^{s})$, and $(K^{k})_{k\geq 1}$ denote the sequence of (potentially un-normalized) kernels acting on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$ defined by
-$K^{k}(x,A)=\int_{\mathbb{R}^{s}}K(x,\mathrm{d}x^{\prime})K^{k-1}(x^{\prime},A),\quad A\in\mathcal{B}(\mathbb{R}^{s}),\quad k\geq 1$
-with the convention that $K^{k-1}(x^{\prime},\mathrm{d}x)=\delta_{\{x^{\prime}\}}(\mathrm{d}x)$ for all $x^{\prime}\in\mathbb{R}^{s}$ when $k=1$.
-Finally, for any probability distribution $\pi\in\mathcal{P}(\mathbb{R})$ we let $F_{\pi}$ denote the cumulative density function of $\pi$ and $F^{-1}_{\pi}$ denote its (generalized-)inverse.
+In what follows all the random variables are defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$, and for any $s\in\mathbb{N}$ we let $\mathcal{B}(\mathbb{R}^{s})$ denote the Borel $\sigma$-algebra on $\mathbb{R}^{s}$, $\mathcal{P}(\mathbb{R}^{s})$ denote the set of probability measures on $(\mathbb{R}^{s},\mathcal{B}(\mathbb{R}^{s}))$, $\lambda_{s}$ denote the Lebesgue measure on $\mathbb{R}^{s}$, and for any point set $\{u^{n}\}_{n=1}^{N}$ in $0,1)^{s}$ we denote by $D^{*}_{s}\big(\{u^{n}\}_{n=1}^{N}\big)$ its star discrepancy, that is
+$D^{*}_{s}\big(\{u^{n}\}_{n=1}^{N}\big)=\sup_{b\in[0,1)^{s}}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{[0,b)}(u^{n})-\lambda_{s}\big([0,b)\big)\Big|.$
+Next, for any probability measure $\pi\in\mathcal{P}(\mathbb{R})$ we let $F_{\pi}$ denote its cumulative density function and $F^{-1}_{\pi}$ denote the (generalized-)inverse of $F_{\pi}$.
+In this notion, the Kolmogorov distance between the probability measures $\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R})$ is given by $\|\pi_{1}-\pi_{2}\|=\sup_{a\in\mathbb{R}}|F_{\pi_{1}}(a)-F_{\pi_{2}}(a)|$, and we denote by $\|\pi_{1}-\pi_{2}\|_{\mathrm{D}}$ the discrepancy metric between these two probability measures, that is we let
+$\|\pi_{1}-\pi_{2}\|_{\mathrm{D}}=\sup_{-\infty<a<b<\infty}\big|\pi_{1}([a,b])-\pi_{2}([a,b])\big|.$
+We recall the reader that these two norms on $\mathcal{P}(\mathbb{R})$ are related together through the inequalities
+$\|\pi_{1}-\pi_{2}\|\leq\|\pi_{1}-\pi_{2}\|_{\mathrm{D}}\leq 2\|\pi_{1}-\pi_{2}\|,\quad\forall\pi_{1},\pi_{2}\in\mathcal{P}(\mathbb{R}).$ (2)
+With the exception of Propositions 2 and 3, the results presented below are expressed in term of the Kolmogorov metric, which is arguably more popular than the discrepancy metric.
+These results can however be easily expressed in term of the latter metric by using (2).
+Propositions 2 and 3 are expressed using the discrepancy metric for reasons that will be provided in due course (see Remarks 2 and 3).
+To introduce some further notation we let $\pi\in\mathcal{P}(\mathbb{R})$, $f:\mathbb{R}\rightarrow\mathbb{R}$ be a measurable function and $K$ be a (potentially un-normalized) kernel acting on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$.
+Then, we let $\pi(f)=\int_{\mathbb{R}}f(x)\pi(\mathrm{d}x)$, $\pi K$ denote the measure on $\mathcal{B}(\mathbb{R})$ such that $\pi K(A)=\int_{\mathbb{R}}K(x,A)\pi(\mathrm{d}x)$ for all $A\in\mathcal{B}(\mathbb{R})$, and $(K^{k})_{k\geq 1}$ denote the sequence of (potentially un-normalized) kernels acting on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$ defined by
+$K^{k}(x,A)=\int_{\mathbb{R}}K(x,\mathrm{d}x^{\prime})K^{k-1}(x^{\prime},A),\quad(x,A)\in\mathbb{R}\times\mathcal{B}(\mathbb{R}),\quad k\geq 1.$
+Lastly, below we let $\Phi(\cdot)$ denote the cumulative density function of the $\mathcal{N}(0,1)$ distribution.
 The rest of the paper is organized as follows.
-Section 2 introduces formally the filtering problem and algorithms discussed in this work, and the time evolution of the resulting PF estimates is studied in Section 3.
-The result outlined in the introductory section for SQMC is given in Section 4 and Section 5 concludes.
-All the proof are gathered in the appendix.
+In the next subsection we provide a brief reminder on scrambled $(t,s)$-sequences.
+Section 2 introduces formally the filtering problem and algorithms studied in this work, and our main results are presented in Section 3.
+Main results not proved in Section 3 are proved in Section 4, while Section 5 contains the proof of Proposition 1 presented Section 2 as well as the proof of Lemma 1 stated Section 4, which is the main technical contribution of the paper.
+1.4 Reminder on scrambled $(t,s)$-sequences
+Let $t\geq 0$, $s\geq 1$ and $b\geq 2$ be three integers, and let
+$\left\{\prod_{j=1}^{s}\left[a_{j}b^{-d_{j}},(a_{j}+1)b^{-d_{j}}\right)\subseteq[0,1)^{s},\,a_{j},\,d_{j}\in\mathbb{N},\,a_{j}<b^{d_{j}},\,j=1,...,s\right\}$
+be the set of all $b$-ary boxes.
+Then, for an integer $m\geq t$ the point set $\{u^{n}\}_{n=1}^{b^{m}}$ is called a $(t,m,s)$-net in base $b$ if every $b$-ary box of volume $b^{t-m}$ contains exactly $b^{t}$ points, while a sequence $(u^{n})_{n\geq 1}$ of points in $0,1)^{s}$ is called a $(t,s)$-sequence in base $b$ if, for any integers $a\geq 0$ and $m\geq t$, the point set $\{u^{n}\}_{n=ab^{m}+1}^{(a+1)b^{m}}$ is a $(t,m,s)$-net in base $b$.
+From the above definitions it should be clear that, from a practical perspective, $(t,s)$-sequences in base $b$ with $t=0$ and $b=2$ are preferable, but such sequences exist only for $s\in\{1,2\}$ (Niederreiter,, 1992, Corollary 4.24, page 62).
+In the context of this work, the crucial property of a $(t,s)$-sequence $(u^{n})_{n\geq 1}$ in base $b$ is that, for some constant $C_{s,t,b}\in(0,\infty)$ depending only on $s$, $t$ and $b$, we have (see Niederreiter,, 1992, Section 4.1)
+$D^{*}_{s}\big(\{u^{n}\}_{n=1}^{N}\big)\leq C_{s,t,b}\frac{\log(N)^{s}}{N},\quad\forall N\geq b^{t}.$ (3)
+Finally, a sequence $(U^{n})_{n\geq 1}$ of $(0,1)^{s}$-valued random variables is a scrambled $(t,s)$-sequence in base $b$ if, $\mathbb{P}$-a.s., $(U^{n})_{n\geq 1}$ is a $(t,s)$-sequence in base $b$ and if $U^{n}\sim\mathcal{U}(0,1)^{s}$ for all $n\geq 1$.
+Consequently, a scrambled $(t,s)$-sequence in base $b$ is a sequence $(U^{n})_{n\geq 1}$ of dependent $\mathcal{U}(0,1)^{s}$ random variables such that $D^{*}_{s}\big(\{U^{n}\}_{n=1}^{N}\big)\leq C_{s,t,b}N^{-1}\log(N)^{s}$ with $\mathbb{P}$-probability one and for all $N\geq 1$, and with the constant $C_{s,t,b}$ as in (3).
+A scrambled $(t,s)$-sequence is obtained by randomizing, or scrambling, a $(t,s)$-sequence, and various scrambling methods have been proposed in the literature.
+It is worth mentioning that the first $N$ points of a scrambled sequence can be generated in $\mathcal{O}(N\log(N))$ operations, but the memory requirements of the algorithm needed to perform this task depend on the scrambling procedure that is used (Hong and Hickernell,, 2003).
+We refer the reader to Chapter 3 in Dick and Pillichshammer, (2010) and to Owen, (2023) for a discussion on scrambling algorithms.
 2 Model and filtering algorithms
@@
 $Y_{t}=X_{t}+c^{-1/2}Z_{t},\quad X_{t+1}=\rho X_{t}+\sigma W_{t+1},\quad\forall t\geq 1$ (4)
-with $X_{1}\sim\eta_{1}$ and where the $Z_{t}$’s and $W_{t+1}$’s are independent $\mathcal{N}_{1}(0,1)$ random variables.
+with $X_{1}\sim\eta_{1}$ and where the $Z_{t}$’s and the $W_{t+1}$’s are independent $\mathcal{N}(0,1)$ random variables.
 We now let $(y_{t})_{t\geq 1}$ be a sequence in $\mathbb{R}$ and, for all $t\geq 1$, we denote by $\hat{\eta}_{t}$ the filtering distribution of $X_{t}$, that is the conditional distribution of $X_{t}$ given that $Y_{s}=y_{s}$ for all $s\in\{1,\dots,t\}$ under the model (4).
 In addition, for all $t\geq 2$ we denote by $\eta_{t}$ the predictive distribution of $X_{t}$, that is the conditional distribution of $X_{t}$ given that $Y_{s}=y_{s}$ for all $s\in\{1,\dots,t-1\}$ under the model (4).
-We assume henceforth that $y_{t}=0$ for all $t\geq 1$ and, for all $z\in\mathbb{R}$, we let $M(z,\mathrm{d}x)=\mathcal{N}_{1}(\rho z,\sigma^{2})$, $G(z)=e^{-cz^{2}/2}$ and $Q(z,\mathrm{d}x)=M(z,\mathrm{d}x)G(x)$.
+We assume henceforth that $y_{t}=0$ for all $t\geq 1$ and, for all $z\in\mathbb{R}$, we let $M(z,\mathrm{d}x)=\mathcal{N}(\rho z,\sigma^{2})$, $G(z)=e^{-cz^{2}/2}$ and $Q(z,\mathrm{d}x)=M(z,\mathrm{d}x)G(x)$.
 With this notation and assumption on $(y_{t})_{t\geq 1}$, the two sequences of distributions $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are defined by
@@
 $\eta_{t}(\mathrm{d}x_{t})=\hat{\eta}_{t-1}M(\mathrm{d}x_{t}),\quad\hat{\eta}_{t}(\mathrm{d}x_{t})=\frac{\hat{\eta}_{1}Q^{t-1}(\mathrm{d}x_{t})}{\hat{\eta}_{1}Q^{t-1}(\mathbb{R})},\quad t\geq 2.$ (6)
-The following proposition shows that the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ admit a well-defined limit in $\mathcal{P}(\mathbb{R})$ and that, as $t\rightarrow\infty$, the impact of the initial distribution $\eta_{1}$ on $\hat{\eta}_{t}$ and on $\eta_{t}$ vanishes exponentially fast:
+The following proposition shows that the sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ admit a well-defined limit in $\mathcal{P}(\mathbb{R})$ and that, as $t\rightarrow\infty$, the impact of the initial distribution $\eta_{1}$ on $\hat{\eta}_{t}$ and on $\eta_{t}$ vanishes exponentially fast.
 Proposition 1.
-There exist constants $(C_{\star},\sigma^{2}_{\infty})\in(0,\infty)^{2}$ and $\epsilon_{\star}\in(0,1)$ such that
-$\big{\|}\eta_{t+1}-\mathcal{N}_{1}(0,\rho\sigma^{2}_{\infty}+\sigma^{2})\big{\|}\leq\|\hat{\eta}_{t}-\mathcal{N}_{1}(0,\sigma^{2}_{\infty})\|\leq C_{\star}\epsilon_{\star}^{t},\quad\forall t\geq 1.$
+There exist constants $(C_{\star},\sigma^{2}_{\infty})\in(0,\infty)^{2}$ and $\epsilon_{\star}\in(0,1)$, where only $C_{\star}$ dependents on $\eta_{1}$, such that
+$\|\hat{\eta}_{t}-\mathcal{N}(0,\sigma^{2}_{\infty})\|\leq C_{\star}\epsilon_{\star}^{t},\quad\big\|\eta_{t}-\mathcal{N}(0,\rho^{2}\sigma^{2}_{\infty}+\sigma^{2})\big\|\leq C_{\star}\epsilon_{\star}^{t},\quad\forall t\geq 1.$
 Remark 1.
 When $\eta_{1}$ is a Gaussian distribution the conclusion of the lemma can be obtained from the results in Del Moral and Horton, (2023).
-We stress that, despite the simplicity of the filtering problem that we consider in this work, little is known about the finite $N$ behaviour of the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}$ of $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$.
+We stress that, despite the simplicity of the filtering problem that we consider in this work, little is known about the finite $N$ behaviour of the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}$ of the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$.
 Notably, using Corollary 3 in Caffarel et al., (2025) only allows to establish that, for all functions $f$ in some class of functions (containing unbounded functions) and under the assumption that $|\rho|<1$, we have $\sup_{t\geq 1}\mathbb{E}\big[|\eta_{t}^{N}(f)-\eta_{t}(f)|]\leq CN^{-\beta/2}$ for some unknown constants $C$ and $\beta\in(0,1]$.
-2.2 Particle filters and SQMC
-In the specific case where $\eta_{1}=\mathcal{N}_{1}(\mu_{1},\sigma_{1}^{2})$ for some constants $\mu_{1}\in\mathbb{R}$ and $\sigma_{1}^{2}\in(0,\infty)$ the SSM defined in (2) is a linear Gaussian SSM and the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are sequences of Gaussian distributions whose means and variances can be computed using the Kalman filter.
+2.2 Particle filters and sequential-quasi Monte Carlo
+In the specific case where $\eta_{1}=\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ for some constants $\mu_{1}\in\mathbb{R}$ and $\sigma_{1}^{2}\in(0,\infty)$, the SSM defined in (4) is a linear Gaussian SSM and the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are sequences of Gaussian distributions whose means and variances can be computed using the Kalman filter.
 For other choices of initial distribution $\eta_{1}$ these two sequences are intractable and therefore need to be approximated.
@@
 When in Algorithm 1 the $U_{t}^{n}$’s are independent $\mathcal{U}(0,1)^{2}$ random variables we recover the bootstrap PF with multinomial resampling, which is arguably the simplest PF algorithm.
-It is well-known that the variance of PF estimates can be reduced by replacing multinomial resampling by more advanced resampling schemes and, for instance, Algorithm 1 reduces to the (bootstrap) PF with ordered stratified resampling when the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables while the sets ($(\{U_{t,1}^{n}\}_{n=1}^{N})_{t\geq 1}$ used in the resampling steps are independent sets of $N$ dependent $\mathcal{U}(0,1)$ random variables (see Gerber et al.,, 2019, for more details on resampling algorithms for PFs).
-Sequential quasi-Monte Carlo goes one step further than PF with stratified resampling in the de-randomization process of the vanilla PF (i.e.
-of the PF with multinomial resampling) by also introducing, at each time $t\geq 1$, some dependence in the $\mathcal{U}(0,1)$ random numbers $\{U_{t,2}^{n}\}_{n=1}^{N}$ used to perform the mutation step.
-In addition, for all $t\geq 1$ the $N$ dependent $\mathcal{U}(0,1)^{2}$ random variables $\{U_{t}^{n}\}^{N}_{n=1}$ used in SQMC spread evenly over $(0,1)^{2}$ with $\mathbb{P}$-probability one, that is, with $\mathbb{P}$-probability one these sets of random variables are such that $\sup_{t\geq 1}D^{*}_{2}(\{U_{t}^{n}\}^{N}_{n=1})\leq r_{N}$ for some sequence $r_{N}\rightarrow\infty$.
-Formally, we recover SQMC when, in Algorithm 1, for all $t\geq 1$ the set $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(t,2)$-sequence in based $b$ (see Section 4.1 for a definition), and these random sets are independent.
-In this case, with $\mathbb{P}$-probability one we have $\sup_{t\geq 1}D^{*}_{2}(\{U_{t}^{n}\}_{n=1}^{N})=\mathcal{O}(N^{-1}\log(N)^{2})$.
-Moreover, there exists a finite universal constant $C$ such that, for all $t\geq 1$ and any square integrable function $f:[0,1]^{2}\rightarrow\mathbb{R}$, the variance of the estimate $\frac{1}{N}\sum_{n=1}^{N}f(U_{t}^{n})$ of $\int_{[0,1]^{2}}f(u)\mathrm{d}u$ is smaller than $C$ times the variance we would have obtained if $\{U_{t}^{n}\}_{n=1}^{N}$ was a collection of independent $\mathcal{U}(0,1)^{2}$ random numbers (see Gerber,, 2015, and references therein).
-This latter property is important as it ensures that if for a $p=2$ the result in (1) holds for the vanilla PF then it also holds for SQMC.
-On the other hand, for cost per iteration of SQMC is $\mathcal{O}(N\log(N))$ against $\mathcal{O}(N)$ for PFs.
+By contrast, we recover SQMC when, in Algorithm 1, for all $t\geq 1$ the set $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(t,2)$-sequence in based $b$ (see Section 1.4 for a definition), and these random sets are independent.
 1:let $\{U_{1}^{n}=(U_{1,1}^{n},U_{1,2}^{n})\}_{n=1}^{N}$ be a random point set in $(0,1)^{2}$
@@
 5:  let $\{U_{t}^{n}=(U_{t,1}^{n},U_{t,2}^{n})\}_{n=1}^{N}$ be a random point set in $(0,1)^{2}$
-6:     let $\hat{X}_{t-1}^{n}=F^{-1}_{\hat{\eta}_{t-1}^{N}}(U_{t,1}^{n})$ , $X_{t}^{n}=F^{-1}_{M(\hat{X}_{t-1}^{n},\mathrm{d}x)}(U_{t,2}^{n})$ and $W_{t}^{n}=G(X_{t}^{n})/\sum_{m=1}^{N}G(X_{t}^{m})$
-7:     let $\eta_{t}^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}$ and $\hat{\eta}_{t}^{N}=\sum_{n=1}^{N}W_{t}^{n}\delta_{\{X_{t}^{n}\}}$
-8:end for
+6:  let $\hat{X}_{t-1}^{n}=F^{-1}_{\hat{\eta}_{t-1}^{N}}(U_{t,1}^{n})$ [resampling]
+7:  let $X_{t}^{n}=F^{-1}_{M(\hat{X}_{t-1}^{n},\mathrm{d}x)}(U_{t,2}^{n})$ [mutation]
+8:  let $W_{t}^{n}=G(X_{t}^{n})/\sum_{m=1}^{N}G(X_{t}^{m})$
+9:  let $\eta_{t}^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}$ and $\hat{\eta}_{t}^{N}=\sum_{n=1}^{N}W_{t}^{n}\delta_{\{X_{t}^{n}\}}$
+10:end for
 Algorithm 1 Generic filtering algorithm
 (Operations with index $n$ must be performed for all $n\in\{1,\dots,N\}$)
-2.3 Time uniform bounds for deterministic filtering algorithms
-In the previous subsection we focussed on stochastic versions of Algorithm 1, that is on the case where $U_{t}^{n}\sim\mathcal{U}(0,1)^{2}$ for all $n\in\{1,\dots,N\}$ and all $t\geq 1$.
-Alternatively, one may consider determinist versions of the algorithm in which the $U_{t}^{n}$’s are non-random.
-We do not necessarily advocate for the use of such filtering algorithms but, as it will be clear in what follows, studying Algorithm 1 based on deterministic elements of $(0,1)^{2}$ is a key step for analysing the time evolution of $\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ and of $\|\eta_{t}^{N}-\eta_{t}\|$ for both PF and SQMC.
-In particular, the following lemma will play a key role to study these two algorithms:
-Lemma 1.
-Let $N\geq 1$ and for all $t\geq 1$ let $\big{\{}u_{t}^{n}=(u_{t,1},u_{t,2}^{n})\big{\}}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ such that, for $i=1,2$, we have $u_{t,i}\neq a2^{-k}$ for all $k\in\{1,\dots,7\}$ and all $a\in\{0,\dots,2^{7}\}$, and such that $\sup_{t\in\{1,\dots,T\}}D^{*}(\{u_{t}^{n}\}_{n=1}^{N})<2^{-10}$ for some $T\in\mathbb{N}\cup\{\infty\}$.
-Consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}=\{u_{t}^{n}\}_{n=1}^{N}$ for all $t\geq 1$.
-Then, there exists a constant $\bar{C}\in(0,\infty)$, depending only on $(c,\sigma,\rho)$ and $\eta_{1}$, such that
-$\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\bar{C}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\quad\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\bar{C}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}}$
-with $\epsilon_{\star}\in(0,1)$ as in Proposition 1 and with $\delta_{N,T}=\sup_{t\in\{1,\dots,T\}}D^{*}(\{u_{t}^{n}\}_{n=1}^{N})$.
-We stress that, in the lemma, we did not aim at optimizing the assumptions on ($(\{u_{t}^{n}\}_{n=1}^{N})_{\geq 1}$ but rather at obtaining a result that holds under weak assumptions on this sequence of sets.
-In particular, when $\{u_{t}^{n}\}_{n=1}^{N}$ is the realization of an RQMC point set the upper bound on $D^{*}(\{u_{t}^{n}\}_{n=1}^{N})$ imposed in the lemma can be drastically reduced (see Remark 5 in the Appendix).
-In practice we are usually interested in estimating the expectation of some function $f$ under the filtering distributions $(\hat{\eta}_{t})_{t\geq 1}$ and under the predictive distributions $(\eta_{t})_{t\geq 1}$.
-Building on the result of Lemma 1, the following lemma will be used below to address this problem using PFs and SQMC:
-Lemma 2.
-Consider the set up of Lemma 1 with $T\geq 2$.
-Then, there exist constants $(\varrho_{1},\varrho_{2})\in(0,\infty)^{2}$, depending only on $c$, $\sigma$, $\rho$ and $\eta_{1}$, such that for any function $f\in\mathcal{C}^{1}(\mathbb{R})$ verifying $\sup_{x\in\mathbb{R}}|f(x)|e^{-\varrho_{1}x^{2}}<\infty$ there is a constant $C_{f}\in(0,\infty)$, depending only on $f$, $c$, $\sigma$, $\rho$ and $\eta_{1}$, for which we have, for all $a\in(0,\infty]$ and all $\gamma\in(0,1)$, and with $\epsilon_{\star}$ and $\bar{C}$ as in Lemma 1,
-$\sup_{t\in\{2,\dots,T\}}\bigg{(}\big{|}\hat{\eta}_{t}^{N}\big{(}f\mathbbm{1}_{[-a,a]}\big{)}-\hat{\eta}_{t}(f)\big{|}\vee\big{|}\eta_{t}^{N}\big{(}f\mathbbm{1}_{[-a,a]}\big{)}-\eta_{t}(f)\big{|}\Big{)}\leq\delta_{N,T}^{\gamma/2}\frac{\bar{C}F(a)}{1-\epsilon_{\star}^{1-\gamma}}+C_{f}e^{-\varrho_{2}a^{2}}$
-where $F(a)=|f(a)|+|f(-a)|+\int_{-a}^{a}|f^{\prime}(x)|\mathrm{d}x$ with the convention that $F(a)=\lim_{z\rightarrow\infty}F(z)$ when $a=\infty$.
-3 Time evolution of particle filter estimates
-3.1 Almost sure behaviour of particle filter estimates
-We consider the use of Algorithm 1 to approximate the sequences of distributions $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ defined in Section 2.1, where the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables.
-As explained in Section 2.2, this framework includes the bootstrap PF with both multinomial resampling and some more advanced resampling techniques.
-In fact, as it will be clear from the calculations that follow, the main result of this subsection (Proposition 2) holds for any resampling mechanism.
-Let $N\geq 1$, $\mathcal{F}^{N}_{0}=\{0,\Omega\}$ be the trivial $\sigma$-algebra and for all $t\geq 1$ let
-$\mathcal{F}^{N}_{t}=\sigma\Big{(}\big{\{}(X_{s}^{n},\hat{X}_{s}^{n}),\,n=1,\dots,N,\,\,s=1,\dots,t\Big{)}.$
-Next, let $[a,b]\subset\mathbb{R}$ be an arbitrary non-empty interval and note that
-$\inf_{x^{\prime}:\,\rho x^{\prime}\in[a,b]}M\big{(}x^{\prime},[a,b]^{c}\big{)}\geq 1-\Phi\big{(}2(b-a)/\sigma\big{)},\quad\inf_{x^{\prime}:\,\rho x^{\prime}\not\in[a,b]}M\big{(}x^{\prime},[a,b]^{c}\big{)}\geq\frac{1}{2}.$
-Therefore, letting $\varrho=\min\big{\{}1-\Phi(2(b-a)/\sigma),1/2\big{\}}>0$, it follows that
+3 Main results
+3.1 Almost sure behaviour of particle filters
+The following proposition is the first result announced in the introductory section (see also Remark 2):
+Proposition 2.
+Let $N\geq 1$ and consider Algorithm 1 where $\big(\{U_{t,2}^{n}\}_{n=1}^{N}\big)_{t\geq 1}$ is a sequence of independent sets of $N$ independent $\mathcal{U}(0,1)$ random variables. Then,
+$\mathbb{P}\big(\hat{\eta}_{t}^{N}\big([a,b]\big)=0,\,\,i.o.\big)=\mathbb{P}\big(\eta_{t}^{N}\big([a,b]\big)=0,\,\,i.o.\big)=1,\quad\forall[a,b]\subset\mathbb{R}$ (7)
+and we have
+$\mathbb{P}\big(\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\geq\kappa,\,\,i.o.\big)=\mathbb{P}\big(\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\geq\kappa,\,\,i.o.\big)=1,\quad\forall\kappa\in(0,1).$
+Remark 2.
+Using (2), it follows that if we replace the discrepancy metric $\|\cdot\|_{\mathrm{D}}$ by the Kolmogorov metric $\|\cdot\|$ then the conclusion of Proposition 2 holds only for all $\kappa\in(0,1/2)$.
+3.1.1 Proof of Proposition 2
+Proof of Proposition 2.
+To prove the first statement of the proposition we let $\mathcal{F}^{N}_{0}=\{0,\Omega\}$ be the trivial $\sigma$-algebra,
+$\mathcal{F}^{N}_{t}=\sigma\Big(\big\{(X_{s}^{n},\hat{X}_{s}^{n}),\,n=1,\dots,N,\,\,s=1,\dots,t\Big),\quad\forall t\geq 1$
+and $a,b]\subset\mathbb{R}$ be arbitrary. Then, letting $\varrho=\min\big\{1-\Phi(2(b-a)/\sigma),1/2\big\}>0$, we easily check that
+$\inf_{x^{\prime}:\,\rho x^{\prime}\in[a,b]}M\big(x^{\prime},[a,b]^{c}\big)\geq 1-\Phi\big(2(b-a)/\sigma\big)\geq\varrho,\quad\inf_{x^{\prime}:\,\rho x^{\prime}\not\in[a,b]}M\big(x^{\prime},[a,b]^{c}\big)\geq\frac{1}{2}\geq\varrho$
+from which we deduce that
 $\mathbb{P}\big(\eta_{t}^{N}([a,b])=0\big|\mathcal{F}^{N}_{t-1}\big)\geq\varrho^{N}>0,\quad\forall t\geq 1,\quad\mathbb{P}-a.s.$ (8)
-Then, using the second Borel-Cantelli lemma for dependent events given in Durrett, (2019, Theorem 4.3.4, page 225), the result in (5) implies that with $\mathbb{P}$-probability one we have $\eta_{t}^{N}([a,b])=0$ infinitely often.
-Since $N\geq 1$ and the interval $[a,b]$ are arbitrary, we obtain in particular that
-$\mathbb{P}\big{(}\hat{\eta}_{t}^{N}\big{(}[-a,a]^{c}\big{)}=1,\,\,i.o.\big{)}=\mathbb{P}\big{(}\eta_{t}^{N}\big{(}[-a,a]^{c}\big{)}=1,\,\,i.o.\big{)}=1,\quad\forall N\geq 1,\quad\forall a\in(0,\infty).$ (6)
-In words, for any number of particles $N\geq 1$ and interval $[-a,a]\subset\mathbb{R}$, with $\mathbb{P}$-probability one the set $\{X_{t}^{n}\}_{n=1}^{N}$ contains no elements in $[-a,a]$ for infinitely many time instants $t\geq 1$.
-By combining (6) with Proposition 1 we readily obtain the first result announced in the introductory section:
+By using the second Borel-Cantelli lemma for dependent events given in Durrett, (2019, Theorem 4.3.4, page 225), it follows from (8) that with $\mathbb{P}$-probability one we have $\eta_{t}^{N}([a,b])=0$ infinitely often.
+Since we have the implication $\eta_{t}^{N}([a,b])=0\implies\hat{\eta}_{t}^{N}([a,b])=0$, the proof of (7) is complete.
+To prove the second statement of the proposition we let $\kappa\in(0,1)$ be arbitrary, and we let $a_{\kappa}\in(0,\infty)$ and $T_{\kappa}\in\mathbb{N}$ be such that $\hat{\eta}_{t}([-a_{\kappa},a_{\kappa}])\geq\kappa$ for all $t\geq T_{\kappa}$.
+Remark that such constants $a_{\kappa}$ and $T_{\kappa}$ exist by Proposition 1.
+Then, for any $t\geq T_{\kappa}$ we have the implication
+$\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0\implies\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\geq|\hat{\eta}_{t}^{N}([a_{\kappa},a_{\kappa}])-\hat{\eta}_{t}([-a_{\kappa},a_{\kappa}])|\geq\kappa$
+and thus it follows from (7) that $\mathbb{P}\big(\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\geq\kappa,\,\,i.o.\big)=1$.
+Following similar calculations we obtain that $\mathbb{P}\big(\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\geq\kappa,\,\,i.o.\big)=1$, completing the prove of the proposition.
+∎
+3.1.2 Some remarks
+The result stated in (7) is the key result of Proposition 2, the second part of the proposition following directly from (7) and Proposition 1.
+Inspecting the proof of Proposition 2 reveals that the negative result (7) concerning the behaviour of PFs arises for a simple and intuitive reason.
+On the one hand, since for all $t\geq 1$ the uniform random variables $\{U_{t,2}^{n}\}_{n=1}^{N}$ used by a PF are independent, for any interval $a,b]\subset\mathbb{R}$ there exists a constant $\varrho>0$ such that, after each mutation step of the algorithm (Line 7 in Algorithm 1), all the particles are outside the interval $[a,b]$ with probability at least $\varrho^{N}$.
+On the other hand, because the different mutation steps of a PF rely on independent sets of uniform random variables, with probability one this event occurs infinity often, by a suitable version of the second Borel-Cantelli lemma.
+Following the discussion in the previous paragraph, and more formally following the calculations in Section 3.1.1, it should be clear that the result in (7) is not specific to the considered filtering problem but will hold in general.
+Notably, the problem identified in (7) is entirely caused by the mutation steps of PF algorithms.
+This observation implies, for instance, that (7) still holds if for all $t\geq 1$ we replace the function $G(x)=e^{-cx^{2}/2}$ by any measurable function $G_{t}:\mathbb{R}\rightarrow(0,\infty)$.
+We also stress that (7) is not due to the fact that we consider an unbounded state-space.
+To illustrate this point let $\mathsf{X}=[-l,l]$ for some constant $l\in(0,\infty)$, and let $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ be as defined in (5)-(6) where, for all $x^{\prime}\in\mathbb{R}$, $M(x^{\prime},\mathrm{d}x)$ denote the $\mathcal{N}(\rho x^{\prime},\sigma^{2})$ distribution truncated on $\mathsf{X}$ and where $\eta_{1}$ is such that $\eta_{1}(\mathsf{X})=1$.
+In this context, by repeating the calculations in Section 3.1.1, it is trivial to verify that if in Algorithm 1 the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables then
+$\mathbb{P}\big(\hat{\eta}_{t}^{N}\big([-a,a]\big)=0,\,\,i.o.\big)=\mathbb{P}\big(\eta_{t}^{N}\big[-a,a]\big)=0,\,\,i.o.\big)=1,\quad\forall N\geq 1,\quad\forall a\in(0,l).$
+Interestingly, in the literature it has been realized that if, in Algorithm 1, the function $G$ is defined by $G(x)=\mathbbm{1}_{\mathbb{R}\setminus I}(x)$ for some compact interval $I\subset\mathbb{R}$ then it may happen that, at some time $t\geq 1$, all the particles $\{X_{t}^{n}\}_{n=1}^{N}$ fall into $I$.
+In this case, the particle system is said to be extinct, and some estimates for the probability that this event happens before a given time instant $T\geq 2$ are provided in Del Moral, (2004, Section 7.4.1, page 231).
+Surprisingly, it has not been noticed that the problem whereby all the particles $\{X_{t}^{n}\}_{n=1}^{N}$ are located in an arbitrary given compact interval $I\subset\mathbb{R}$ (i) arises independently of the definition of the function $G$, and (ii) implies that PFs cannot be used to reliably estimate filtering distributions over an infinite time horizon.
+3.2 Finite time horizon and probabilistic behaviour of particle filters
+By Proposition 2, as time progresses a PF will inevitably reach a time instant $\tau$ at which it will fail at estimating well the filtering and/or the predictive distribution of the SSM of interest.
+From a practical point of view it is of key interest to study the probability that such an undesirable event happens in a finite number $T$ of time steps.
+The following theorem addresses this question:
+Theorem 1.
+Let $N\geq 1$ and consider Algorithm 1 where $\big(\{U_{t}^{n}\}_{n=1}^{N}\big)_{t\geq 1}$ is a sequence of independent sets of $N$ independent $\mathcal{U}(0,1)^{2}$ random variables.
+Then, there exists a constant $\bar{C}_{1}\in(0,\infty)$ (independent of $N$) such that, for all $q\in(0,1)$ and all $T\geq 1$, with probability at least 1q1-q we have
+$\max\bigg\{\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\bigg\}\leq\bar{C}_{1}\delta_{N,T,q}^{1/2}\log\big(1+\delta_{N,T,q}^{-1/2}\big)$
+with $\delta_{N,T,q}=N^{-1}\big(1+\log(T/q)\big)$.
+The next result is a direct consequence of Theorem 1.
+Corollary 1.
+Consider the set-up of Theorem 1. Then, there exists a constant $\bar{C}^{\prime}_{1}\in(0,\infty)$ such that, for any $(q,\kappa)\in(0,1)^{2}$ and $T\geq 1$, we have
+$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa\Big)\geq 1-q,\quad\forall N\geq N_{T,\kappa,q}$
+with $N_{T,\kappa,q}=\bar{C}^{\prime}_{1}\big\{1+\log(1+\kappa^{-1})\big\}^{2}\kappa^{-2}\big(1+\log(T/q)\big)$.
+The lower bound on $N$ provided in Corollary 1 to ensure a given (probabilistic) estimation error depends on $(T,q)$ only through the quantity $\log(T/q)$.
+As shown in the next proposition, this dependence in the parameters $T$ and $q$ is sharp:
 Proposition 3.
-Let $N\geq 1$ and consider Algorithm 1 where ($\big{(}\{U_{t,2}^{n}\}_{n=1}^{N}\big{)}_{t\geq 1}$ is a sequence of independent sets of independent $\mathcal{U}(0,1)$ random variables. Then,
-$\mathbb{P}\big{(}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa,\,\,i.o.\big{)}=\mathbb{P}\big{(}\|\eta_{t}^{N}-\eta_{t}\|\geq\kappa,\,\,i.o.\big{)}=1,\quad\forall N\geq 1,\quad\forall\kappa\in(0,1).$
-3.2 Discussion
-As shown in the above calculations, the negative result given in (6) (and thus in Proposition 2) for PFs (i.e.
-for Algorithm 1 where the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables) arises for the following simple and intuitive reason.
-On the one hand, since for all $t\geq 1$ the uniform random variables $\{U_{t,2}^{n}\}_{n=1}^{N}$ used by a PF are independent, each mutation step of the algorithm moves all the particle outside any given interval $[a,b]\subset\mathbb{R}$ with a constant positive probability.
-On the other hand, because the different mutation steps of a PF rely on independent sets of uniform random variables this event will necessarily happen infinity often.
-Interestingly, in the literature it has been realized that if, in Algorithm 1, the function $G$ is defined by $G(x)=\mathbbm{1}_{\mathbb{R}\setminus I}(x)$ for some bounded interval $I\subset\mathbb{R}$ then, for a PF, it may happen that at some time $t\geq 1$ all the particles $\{X_{t}^{n}\}_{n=1}^{N}$ fall into $I$, in which case the particle system is said to be extinct (Del Moral,, 2004, see, Section 7.2.2, page 219).
-Surprisingly, it has not been realized that this positive probability to have, at any time $t\geq 1$, all the particles located in a given interval implies that PFs are not suitable for estimating an infinite sequence of filtering distributions.
-Above we establish (6) using only the properties of the mutation steps of Algorithm 1.
-This observation has two key implications.
-Firstly, considering more sophisticated resampling schemes than those allowed by Algorithm 1 will not prevent (6) to hold.
-Secondly, this latter result remains true if for all $t\geq 1$ we replace the function $G(x)=e^{-cx^{2}/2}$ by some arbitrary function $G_{t}:\mathbb{R}\rightarrow(0,\infty)$.
-Notably, for the SSM considered in this work (6) holds for any sequence of observations $(y_{t})_{t\geq 1}$.
-More generally, by looking at the calculations used to prove (6) it seems hard to think of a meaningful SSM for which (6) does not hold.
-In particular, we stress that (6) is not due to the fact that we consider an unbounded state-space.
-To illustrate this point let $\mathsf{X}=[-l,l]$ for some constant $l\in(0,\infty)$, and let $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ be as defined in (3)-(4) where, for all $x^{\prime}\in\mathbb{R}$, $M(x^{\prime},\mathrm{d}x)$ denote the $\mathcal{N}_{1}(\rho x^{\prime},\sigma^{2})$ distribution truncated on $\mathsf{X}$ and where $\eta_{1}$ is such that $\eta_{1}(\mathsf{X})=1$.
-Then, by repeating the above calculations, it is trivial to verify that if in Algorithm 1 the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables then
-$\mathbb{P}\Big{(}\hat{\eta}_{t}^{N}\big{(}[-l,l]\setminus[-a,a]\big{)}=1,\,\,i.o.\Big{)}=\mathbb{P}\Big{(}\eta_{t}^{N}\big{(}[-l,l]\setminus[-a,a]\big{)}=1,\,\,i.o.\Big{)}=1,\quad\forall N\geq 1,\quad\forall a\in(0,l).$
-3.3 Finite time horizon and probabilistic behaviour of particle filter estimates
-As proved above, as time progress a PF will inevitably reach a time instant $\tau$ at which it will fail at estimating well the filtering and/or the predictive distribution of the SSM of interest.
-From a practical point of view it is of key interest to study the probability that such an undesirable event happens in a finite number $T$ of time steps.
-Interestingly, Lemma 1, obtained for deterministic versions of Algorithm 1, can be used to address this problem.
-This is the case because if, in the algorithm, the $U_{t}^{n}$’s are independent $\mathcal{U}(0,1)^{2}$ random variables then $\sup_{t\in\{1,\dots,T\}}D^{*}_{2}(\{U_{t}^{n}\}_{n=1}^{N})\leq\sqrt{2/N}$ with high probability, provided that $N$ is sufficiently large.
-More precisely, using Aistleitner and Hofer, (2014, Theorem 1) we have the following key result:
-Lemma 3.
-Let $(s,T)\in\mathbb{N}^{2}$, $(\delta,q)\in(0,1)^{2}$ and, for all $N\geq 1$, let $\{\{U_{t}^{n}\}_{n=1}^{N}\}_{t=1}^{T}$ be $T$ independent sets of $N$ independent $\mathcal{U}(0,1)^{s}$ random variables.
-Then,
-$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}D^{*}_{s}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}\geq 1-q,\quad\forall N\geq N_{s,T,q}:=\frac{160s+33\log(T)-33\log(q)}{\delta^{2}}.$
-By combining Lemma 1 and Lemma 3 we easily obtain the following result for PF with multinomial resampling:
-Theorem 1.
-Let $N\geq 1$ and consider Algorithm 1 where ($\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}_{t\geq 1}$ is a sequence of independent sets of $N$ independent $\mathcal{U}(0,1)^{2}$ random variables.
-Then, there exists a constant $\bar{C}_{1}\in[1,\infty)$ (independent of $N$) such that, for all $q\in(0,1)$ and all $T\in\mathbb{N}$, we have
-$\mathbb{P}\Bigg{(}\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\bar{C}_{1}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T,q}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\,\,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\bar{C}_{1}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T,q}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}}\Bigg{)}\geq 1-q$
-with $\epsilon_{\star}\in(0,1)$ as in Proposition 1 and with $\delta_{N,T,q}=N^{-1}\big{(}160+33\log(T)-33\log(q)\big{)}$.
+Consider the set-up of Theorem 1 and let $\bar{q}\in(0,1/2)$.
+Then, for all $\kappa\in(0,1]$, there exist constants $C_{\kappa}\in(0,\infty)$ and $T_{\kappa}\in\mathbb{N}$ such that, for all $q\in(0,\bar{q}]$ and all $T\geq T_{\kappa}$, we have
+$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\leq\kappa\Big)<1-q,\quad\forall N<C_{\kappa}\log(T/q).$
 Remark 3.
-A direct application of Lemma 1 and of Lemma 3 would lead the conclusion of Theorem 1 to hold with $\delta_{N,T,q}$ replaced by $\delta_{N,T,q}^{1/2}$.
-In Theorem 1 we exploit a particular property of independent uniform random variables to improve this result (see the proof of the theorem for more details).
-Remark 3.
-In Theorem 1 the condition $C_{1}\geq 1$ comes from Lemma 3, and implies that the probabilistic bounds given in the theorem are meaningful (i.e.
-smaller than one) only when $N>N_{1,T,q}$, with $N_{1,T,q}$ as defined in Lemma 3.
-A direct implication of Theorem 1 is that if, for some $\kappa\in(0,1)$, we have
-$N\geq\Big{(}160+33\log(T)-33\log(q)\Big{)}\inf_{\gamma\in(0,1)}\kappa^{-\frac{2}{\gamma}}\bigg{(}\frac{\bar{C}_{1}}{1-\epsilon_{\star}^{1-\gamma}}\bigg{)}^{\frac{2}{\gamma}}$ (7)
-then with probability at least $1-q$ we have $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ and $\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa$.
-This constitutes a reassuring result concerning the time evolution of PF estimates for two reasons.
-Firstly, it shows that, as the time horizon $T$ that we consider increases, the number of particles $N$ should grow at most at speed $\log(T)$ to ensure that the PF algorithm estimates the whole sets of distributions $\{\hat{\eta}_{t}\}_{t=1}^{T}$ and $\{\eta_{t}\}_{t=1}^{T}$ with a constant (probabilistic) error.
-Secondly, Theorem 1 establishes that $N$ should grow at most at speed $\log(1/q)$ if we want to increase the confidence level $1-q$.
-The following proposition shows that these results concerning the dependence in $T$ and in $q$ of the number of particles needed to have both $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ and $\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa$ with probability at least $1-q$ are sharp:
-Proposition 3.
-Consider the set-up of Theorem 1.
-Then, for all $\kappa\in(0,1)$ there exist constants $\varrho_{\kappa}\in(0,1)$ and $T_{\kappa}\in\mathbb{N}$ such that, for all $q\in(0,1)$ and all $T\geq T_{\kappa}$, we have
-$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa\Big{)}<1-q,\quad\forall N<\frac{\log(T)-\log(q)}{\log(1/\varrho_{\kappa})}.$
-By using Lemma 2 and Theorem 1 we can easily obtain probabilistic bounds for the joint estimation of all the filtering and predictive expectations $\{\hat{\eta}_{t}(f)\}_{t=1}^{T}$ and $\{\eta_{t}(f)\}_{t=1}^{T}$ for a large class of functions, as illustrated with the following proposition:
-Proposition 4.
-Consider the set-up of Theorem 1 and let $\gamma\in(0,1)$ and $\mathfrak{F}\subset\mathcal{C}^{1}(\mathbb{R})$.
-Then, there exists a constant $C_{\mathfrak{F}}\in(0,\infty)$ such that, for all $T\geq 2$, all $q\in(0,1)$ and all $N\geq 1$, we have
-$\mathbb{P}\bigg{(}\sup_{t\in\{2,\dots,T\}}\sup_{f\in\mathfrak{F}}\Big{(}\big{|}\hat{\eta}_{t}^{N}\big{(}f\mathbbm{1}_{[-a_{N},a_{N},]}\big{)}-\hat{\eta}_{t}(f)\big{|}\vee\big{|}\eta_{t}^{N}\big{(}f\mathbbm{1}_{[-a_{N},a_{N}]}\big{)}-\eta_{t}(f)\big{|}\Big{)}\leq C_{\mathfrak{F}}\,\delta_{N,T,q}^{\gamma\beta/2}\bigg{)}\geq 1-q$
-for
-  • $\mathfrak{F}=\big{\{}f\in\mathcal{C}^{1}(\mathbb{R}):\,\int_{\mathbb{R}}|f^{\prime}(x)|\mathrm{d}x=1\big{\}}$, if $a_{N}=\infty$ and if $\beta=1$.
-  • $\mathfrak{F}=\big{\{}f\in\mathcal{C}^{1}(\mathbb{R}):\,|f(x)|\leq 1+|x|\text{ and }|f^{\prime}(x)|\leq 1,\,\forall x\in\mathbb{R}\big{\}}$, if $a_{N}=(1\wedge\delta_{N,T,q})^{-\gamma\alpha/2}$ for some constant $\alpha\in(0,1)$ and if $\beta=1-\alpha$.
-  • $\mathfrak{F}=\big{\{}f\in\mathcal{C}^{1}(\mathbb{R}):\,|f(x)|\leq 1+e^{\zeta|x|}\text{and }|f^{\prime}(x)|\leq e^{\zeta|x|},\,\,\forall x\in\mathbb{R}\big{\}}$ for some constant $\zeta\in(0,\infty)$, if $a_{N}=-\gamma\alpha/(2\zeta)\log(1\wedge\delta_{N,T,q})$ for some constant $\alpha\in(0,1)$ and if $\beta=1-\alpha$.
-In the last two examples given in Proposition 4, estimating e.g.
-$\hat{\eta}_{t}(f)$ with the truncated estimator $\hat{\eta}_{t}^{N}(f\mathbbm{1}_{[-a_{N},a_{N}]})$ is unusual (the typical estimate being the empirical counterpart $\hat{\eta}_{t}^{N}(f)$) but has an intuitive explanation.
-In these two examples the function $f$ is unbounded and thus, informally speaking, even a small estimation error of $\hat{\eta}_{t}$ in a region $R\subset\mathbb{R}$ where $\inf_{x\in R}|f(x)|$ is large will have a significant impact on the overall estimation error $|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|$.
-4 Almost sure behaviour of sequential quasi-Monte Carlo
-4.1 Remainder on scrambled $(t,s)$-sequences
-Let $s\geq 1$ and $b\geq 2$ be two integers,
-$\left\{\prod_{j=1}^{s}\left[a_{j}b^{-d_{j}},(a_{j}+1)b^{-d_{j}}\right)\subseteq[0,1)^{s},\,a_{j},\,d_{j}\in\mathbb{N},\,a_{j}<b^{d_{j}},\,j=1,...,s\right\}$
-be the set of all $b$-ary boxes and $t\in\mathbb{N}_{0}$.
-Then, for an integer $m\geq t$ the point set $\{u^{n}\}_{n=1}^{b^{m}}$ is called a $(t,m,s)$-net in base $b$ if every $b$-ary box of volume $b^{t-m}$ contains exactly $b^{t}$ points, while a sequence $(u^{n})_{n\geq 1}$ of points in $[0,1)^{s}$ is called a $(t,s)$-sequence in base $b$ if, for any integers $a\geq 0$ and $m\geq t$, the point set $\{u^{n}\}_{n=ab^{m}+1}^{(a+1)b^{m}}$ is a $(t,m,s)$-net in base $b$.
-From these definitions it should be clear that, from a practical perspective, $(t,s)$-sequences in base $b$ with $t=0$ and $b=2$ are preferable but such sequences exist only for $s\in\{1,2\}$, respectively (Niederreiter,, 1992, Corollary 4.24, page 62).
-In the context of this work, the crucial property of a $(t,s)$-sequence $(u^{n})_{n\geq 1}$ in base $b$ is to be such that, for some constant $C_{s,t,b}\in(0,\infty)$ depending only on $s$,$t$ and $b$, we have (see Niederreiter,, 1992, Section 4.1)
-$D^{*}_{s}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}\leq C_{s,t,b}\frac{\log(N)^{s}}{N},\quad\forall N\geq b^{t}.$ (8)
-Finally, a sequence $(U^{n})_{n\geq 1}$ of $(0,1)^{s}$-valued random variables is a scrambled $(t,s)$-sequence in base $b$ if, $\mathbb{P}$-a.s., $(U^{n})_{n\geq 1}$ is a $(t,s)$-sequence in base $b$ and if $U^{n}\sim\mathcal{U}(0,1)^{s}$ for all $n\geq 1$.
-A scrambled $(t,s)$-sequence is obtained by randomizing, or scrambling, a $(t,s)$-sequence, and various scrambling methods have been proposed in the literature.
-The first $N$ points of a scrambled sequence can be generated in $\mathcal{O}(N\log(N))$ operations but the memory requirements of the algorithm needed to perform this task depend on the scrambling procedure that is used (Hong and Hickernell,, 2003).
-We refer to Chapter 3 in Dick and Pillichshammer, (2010) and to Owen, (2023) for a discussion on scrambling algorithms.
-4.2 Main result
-In the context of this work we only need point sets in $(0,1)^{2}$ and thus, following the discussion of the previous subsection, we limit our attention to scrambled $(0,2)$-sequences in base $b=2$.
-More precisely, in what follows we study Algorithm 1 in the case where $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$.
-In this situation we have, using Niederreiter, (1992, Theorem 4.7, page 54 and Theorem 4.14, page 59),
-$\sup_{t\geq 1}D^{*}_{2}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta_{N},\quad\forall N\geq 1\quad\mathbb{P}-a.s.$
+Using (2), it follows that if we replace the discrepancy metric $\|\cdot\|_{\mathrm{D}}$ by the Kolmogorov metric $\|\cdot\|$ then the conclusion of Proposition 2 holds only for all $\kappa\in(0,1/2)$.
+3.2.1 Proof of Proposition 3
+Proof of Proposition 3.
+Let $\kappa\in(0,1)$ be arbitrary and remark that, by Proposition 1 and (2), there exist constants $a_{\kappa}\in(0,\infty)$ and $T_{\kappa}\in\mathbb{N}$ such that, for all $t\geq T_{\kappa}$, we have both $\eta_{t}([-a_{\kappa},a_{\kappa}])>\kappa$ and $\hat{\eta}_{t}([-a_{\kappa},a_{\kappa}])>\kappa$.
+In addition, let $\varrho_{\kappa}=\min\big\{1-\Phi(4a_{\kappa}/\sigma),1/2\big\}$ and, for all $N\geq 1$, let $(\mathcal{F}^{N}_{t})_{t\geq 0}$ be as defined in the proof of Proposition 2.
+Then, by using (8), we have
+$\mathbb{P}\Big(\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}^{N}_{t-1}\Big)\geq\varrho^{N}_{\kappa},\quad\forall t\geq 1,\quad\forall N\geq 1.$ (9)
+Therefore, noting that for all $t\geq T_{\kappa}$ and $N\geq 1$ we have
+$1-\mathbb{P}\Big(\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\leq\kappa,$ $\,\,\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\leq\kappa|\mathcal{F}_{t-1}^{N}\Big)\geq\mathbb{P}\Big(\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}_{t-1}^{N}\Big)$
+it follows from (9) that, for all $T\geq T_{\kappa}$ and $N\geq 1$,
+$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\leq\kappa\Big)\leq(1-\varrho^{N}_{\kappa})^{T-T_{\kappa}+1}.$ (10)
+To conclude the proof of the proposition we let $T\geq T_{\kappa}$, $\bar{q}\in(0,1/2)$ and $q\in(0,\bar{q}]$ be fixed, and to simplify the notation we let $v_{T}=T-T_{\kappa}+1$.
+Then, we first note that
+$(1-\varrho^{N}_{\kappa})^{v_{T}}<1-q,\quad\forall N<N_{T,\kappa,q}:=\frac{-\log\big(1-(1-q)^{1/v_{T}}\big)}{\log(1/\varrho_{\kappa})}$ (11)
+where, using Bernoulli’s inequality (see e.g. Li and Yeh,, 2013, Theorem 3.1),
+$(1-q)^{1/v_{T}}=\frac{(1-q)^{1+1/v_{T}}}{1-q}\geq\frac{1-q-q/v_{T}}{1-q}.$
+The latter inequality implies that $1-(1-q)^{1/v_{T}}\leq(q/v_{T})/(1-q)\leq(q/v_{T})/(1-\bar{q})$, and thus
+$N_{T,\kappa,q}\geq\frac{\log(v_{T}/q)+\log(1-\bar{q})}{\log(1/\varrho_{\kappa})}\geq\frac{\log(v_{T}/q)}{\log(1/\varrho_{\kappa})}\delta_{\bar{q}},\quad\delta_{\bar{q}}:=1+\frac{\log(1-\bar{q})}{\log(1/\bar{q})}$ (12)
+where $\delta_{\bar{q}}>0$ since $\bar{q}\in(0,1/2)$. The result of the proposition follows from (10)-(12), upon noting that if $T_{\kappa}>1$ then
+$\frac{\log(v_{T}/q)}{\log(T/q)}\geq\frac{\log(v_{T})}{\log(T)}\geq\inf_{x\in[T_{\kappa},\infty)}\frac{\log(x-T_{\kappa}+1)}{\log(x)}>0.$
+where the first inequality uses the fact that $T\geq v_{T}$.
+3.2.2 Some remarks
+Corollary 1 constitutes a reassuring result concerning the time evolution of PF estimates.
+Indeed, it notably shows that, as the time horizon $T$ that we consider increases, the number of particles $N$ only needs to grow at speed $\log(T)$ to ensure that the PF algorithm estimates the whole sets of distributions $\{\hat{\eta}_{t}\}_{t=1}^{T}$ and $\{\eta_{t}\}_{t=1}^{T}$ with a constant (probabilistic) error.
+It is important to stress that the result of Proposition 3 is not specific to the considered filtering problem but will hold in general.
+Indeed, by inspecting the calculations of Section 3.2.1, it should be clear that the conclusion of Proposition 3 holds as soon as the result of Proposition 1 and (7) hold.
+The former result is standard while, as we argued in Section 3.1.2, the latter result is expected to hold for a large class of filtering problems.
+3.3 Almost sure behaviour of sequential quasi-Monte Carlo
+We now turn to the study of SQMC.
+In the context of this work we only need point sets in $(0,1)^{2}$ and thus, following the discussion in Section 1.4, we limit our attention to scrambled $(0,2)$-sequences in base $b=2$.
+More precisely, in this subsection we study Algorithm 1 in the case where $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$, in which case we have, by Niederreiter, (1992, Theorem 4.7, page 54 and Theorem 4.14, page 59),
+$\sup_{t\geq 1}D^{*}_{2}\big(\{U_{t}^{n}\}_{n=1}^{N}\big)\leq\delta_{N},\quad\forall N\geq 1\quad\mathbb{P}-a.s.$ (13)
 with
 $\delta_{N}=\begin{cases}\frac{\log(N)+3}{2N},&\text{ if }N\in\{2^{k},\,k\in\mathbb{N}\}\\ \frac{\log(N)^{2}+11\log(2)\log(N)+18\log(2)^{2}}{N8\log(2)^{2}},&\text{otherwise}\end{cases},\quad\forall N\geq 1.$ (14)
-A direct application of Lemma 1 then gives the following almost sure and time uniform guarantee for the resulting algorithm:
+For the resulting filtering algorithm we have the following almost sure and time uniform guarantee:
 Theorem 2.
 Let $N\geq 1$ and consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$.
-Then, there exists a constant $\bar{C}_{2}\in(0,\infty)$ (independent of $N$) such that, with $\delta_{N}$ as defined in (9) and with $\epsilon_{\star}\in(0,1)$ as in Proposition 1,
-$\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\bar{C}_{2}\inf_{\gamma\in(0,1)}\frac{\delta_{N}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\quad\sup_{t\geq 1}\|\eta_{t}^{N}-\eta_{t}\|\leq\bar{C}_{2}\inf_{\gamma\in(0,1)}\frac{\delta_{N}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\quad\mathbb{P}-a.s.$
+Then, there exists a constant $\bar{C}_{2}\in(0,\infty)$ (independent of $N$) such that, with $\delta_{N}$ as defined in (14),
+$\max\bigg\{\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|,\,\,\sup_{t\geq 1}\|\eta_{t}^{N}-\eta_{t}\|\bigg\}\leq\bar{C}_{2}\,\delta_{N}^{1/2}\log\big(1+\delta_{N}^{-1/2}\big),\quad\mathbb{P}-a.s.$
 Remark 4.
 Theorem 2 does not assume that the random point sets $\{U_{t}^{n}\}_{n=1}^{N}$’s are independent.
-By using Lemma 2 and Theorem 2 we can easily obtain almost sure time uniform bounds for the SQMC estimates of the filtering and predictive expectations $(\hat{\eta}_{t}(f))_{t\geq 1}$ and $(\eta_{t}(f))_{t\geq 1}$ for a large class of functions.
-For instance, for SQMC the conclusion of Proposition 4 holds with $T=\infty$, $q=1$ and with $\delta_{N,T,q}$ replaced by $\delta_{N}$.
-4.3 Discussion
+Theorem 2 implies that the SQMC estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}$ of the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are such that $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|=\sup_{t\geq 1}\|\eta_{t}^{N}-\eta_{t}\|=0$ with $\mathbb{P}$-probability one.
+Moreover, it allows to readily obtain the following result:
+Corollary 2.
+Consider the set-up of Theorem 2. Then, there exists a constant $\bar{C}^{\prime}_{2}\in(0,\infty)$ such that, for any $(q,\kappa)\in(0,1)^{2}$ and $T\geq 1$, the following implication holds:
+$\frac{N}{\log(1+N)^{2}}\geq\bar{C}_{2}^{\prime}\big\{1+$ $\log(1+\kappa^{-1})\big\}^{2}\kappa^{-2}\implies\mathbb{P}\Big(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa\Big)=1.$
+3.3.1 Some remarks
 To simplify the discussion we focus below on the estimation of the filtering distributions but the same comments hold for the estimation of the predictive distributions $(\eta_{t})_{t\geq 1}$.
-Due to the presence of the $\log(N)$ term in the definition of $\delta_{N}$, the dependence in $N$ of bound given in Theorem 2 for $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ is slightly worse than that of the probabilistic bound obtained in Theorem 1 for PFs.
-On the other hand, in the context of this work, using RQMC within PF algorithms has the advantage to produce a non-linear filleting algorithm having time uniform and almost sure guarantees.
-In particular, it follows from Theorem 2 that if $N\in\{2^{k},\,k\in\mathbb{N}\}$ and, for some $\kappa\in(0,1)$, we have
-$N\geq\frac{\log(N)+3}{2}\inf_{\gamma\in(0,1)}\kappa^{-\frac{2}{\gamma}}\bigg{(}\frac{\bar{C}_{2}}{1-\epsilon_{\star}^{1-\gamma}}\bigg{)}^{\frac{2}{\gamma}}$ (10)
-then $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ with $\mathbb{P}$-probability one.
-Unfortunately, we cannot compare the lower bound on $N$ given in (10) with the one obtained in (7), used to ensure that with probability at least $1-q$ the PF estimates $\{\hat{\eta}_{t}^{N}\}_{t=1}^{T}$ are such that $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$, the reason being that there is no reason for the constants $\bar{C}_{1}$ and $\bar{C}_{2}$ appearing in these bounds to be identical.
-Given the properties of scrambled $(t,s)$-sequences, reminded in Section 4.1, it is however reasonable to expect that, in the worst case, the constant $\bar{C}_{2}$ in (10) can only be slightly larger than the constant $\bar{C}_{1}$ appearing in (7).
-In any cases, if we want to compute estimates $\{\hat{\eta}_{t}^{N}\}_{t=1}^{T}$ of $\{\hat{\eta}_{t}\}_{t=1}^{T}$ such that $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ with probability at least $1-q$, then it follows from Theorem 1 and Theorem 2 that there exists a value $T_{\kappa,q}\in\mathbb{N}$ such that SQMC is computationally cheaper than PFs for performing this task for any $T\geq T_{\kappa,q}$.
-Based on (7) and (10), it is reasonable to expect that $T_{\kappa,q}$ is an increasing function of both $\kappa$ and $q$.
-The conclusions drawn in this subsection are based on a toy filtering problem, and future research should aim at establishing if they remain valid when considering more realistic filtering tasks.
-5 Concluding remarks
-Given the popularity of PF algorithms in many applications it is important to understand the behaviour of the resulting joint estimation error, $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.
-Surprisingly, with the exception of Del Moral, (2013, Corollary 14.5.7, page 449), this latter quantity (or a variant of it) does not seem to have been previously studied in the literature.
-The results we obtained in this paper are reassuring, but an important question that has yet to be studied is how the dimension of the filtering problem impacts the behaviour of $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.
-Lemma 1, the key result that allows us to study $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ for PFs as well as the quantity $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ for SQMC, is obtained for a toy filtering problem but we expect that it can be generalized by following a similar proof technique.
-The proof of this lemma, which studies the PF estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ for a given realization $(u_{i})_{i\geq 1}$ of the $\mathcal{U}(0,1)$ random numbers underpinning the algorithm, relies notably on three key building blocks, namely a result showing that the SSM forgets exponentially quickly its initial distribution $\eta_{1}$ as $t\rightarrow\infty$ (illustrated by Proposition 1), a bound for $\big{\|}\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}-\hat{\eta}_{t-1}^{N}M\big{\|}$ (Lemma 8 in the appendix) and a stability result for the particles generated by the algorithm (Lemma 9 in the appendix).
-The first of these three building blocks is standard, in the sense that in the literature on PFs all the existing bounds for $\sup_{t\geq 1}\mathbb{E}[|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|]$ have been established under the assumption that the model forgets exponentially quickly its initial distribution (see Caffarel et al.,, 2025, and references therein), an assumption that has been shown to hold for a large class of SSMs (see e.g.
-Del Moral et al.,, 2023).
-By contrast, to prove Lemma 1 we derive a bound for $\big{\|}\frac{1}{N}\sum_{n=1}^{N}\delta_{\{X_{t}^{n}\}}-\hat{\eta}_{t-1}^{N}M\big{\|}$ which applies only to some one-dimensional SSMs and a stability result for the particle system which is tailored to the specific filtering problem addressed in this work.
-We stress that this latter result is quite strong, as it shows (under a reasonable assumption on $(u_{i})_{i\geq 1}$) that there exists a bounded interval $I\subset\mathbb{R}$ that contains, at any time instant $t\geq 1$, at least a proportion $\varsigma>0$ of the particles.
-Our conjecture is that a result similar to that of Lemma 1 can be established for more realistic filtering problems where this strong stability property of the particle system cannot hold.
-For instance, we expect that this is the case for the SSM considered in this paper if, instead of assuming that $y_{t}=0$ for all $t\geq 1$, we assume that the sequence of observations $(y_{t})_{t\geq 1}$ is unbounded but such that $\sup_{t\geq 1}|y_{t+1}-y_{t}|<\infty$.
-In this work, we focussed on the use of scrambled $(t,s)$-sequences within PF algorithms to improve the safety of the filtering procedure, that is to obtain filtering algorithms for which we have $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|=0$ with $\mathbb{P}$-probability one.
-Future research should aim at determining if other approaches can be used to derive safe filtering algorithms.
-Appendix A Proofs
-A.1 Additional notation
-For a function $f\in\mathcal{C}^{1}(\mathbb{R})$ we let $V(f)=\int_{\mathbb{R}}|f^{\prime}(x)|\mathrm{d}x$ with the convention that $V(f)=0$ if $f^{\prime}(x)=0$ for all $x\in\mathbb{R}$, and, abusing notation, for a function $f:\mathbb{R}\rightarrow\mathbb{R}$ continuously differentiable on some non-empty interval $(a,b)$ we let $V(\mathbbm{1}_{(a,b]}f)=V(\mathbbm{1}_{(a,b)}f)=\int_{a}^{b}|f^{\prime}(x)|\mathrm{d}x$.
-For all $(z,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $M_{r,s^{2}}(z,\mathrm{d}x)=\mathcal{N}_{1}(rz,s^{2})$ and $G_{s}(z)=e^{-sz^{2}/2}$, and for any bounded function $H:\mathbb{R}\rightarrow(0,\infty)$ we let $\Psi_{H}:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be defined by
+Due to the presence of the $\log(N)$ term in the definition of $\delta_{N}$, the dependence in $N$ of the bound given in Theorem 2 for $\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$ is slightly worse than that of the probabilistic bound obtained in Theorem 1 for PFs.
+On the other hand, in the context of this work, SQMC has the advantage to be a filtering algorithm having time uniform and almost sure guarantees.
+Unfortunately, we cannot compare the lower bound on $N$ given in Corollary 2 with the one obtained in Corollary 1 to ensure that, with probability at least 1q1-q, the PF estimates $\{\hat{\eta}_{t}^{N}\}_{t=1}^{T}$ are such that $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$: there is indeed no reason for the constants $\bar{C}_{1}$ and $\bar{C}_{2}$ appearing in these bounds to be identical.
+Given the properties of scrambled $(t,s)$-sequences reminded in Section 1.4, it is however reasonable to expect that, in the worst scenario, the constant $\bar{C}_{2}$ appearing in Corollary 2 can only be slightly larger than the constant $\bar{C}_{1}$ appearing in Corollary 1.
+4 Proof of Theorems 1 and 2 and of Corollaries 1 and 2
+4.1 Preliminary result: A general discrepancy bound for filtering algorithms
+Let $\Phi,\Psi:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be defined
+$\Psi(\mu)(\mathrm{d}x)=\frac{G(x)\mu(\mathrm{d}x)}{\mu(G)},\quad\Phi(\mu)=\Psi(\mu)M,\quad\mu\in\mathcal{P}(\mathbb{R})$
+and note that $\eta_{t+1}=\Phi(\eta_{t})$ for all $t\geq 1$.
+Using this notation, the following result plays a central role in the proof of Theorems 1 and 2:
+Lemma 1.
+There exists a constant $C\in(0,\infty)$ such that, for any sequence $(\mu_{t})_{t\geq 1}$ in $\mathcal{P}(\mathbb{R})$ and letting $\hat{\mu}_{t}=\Psi(\mu_{t})$ for all $t\geq 1$, we have, for any $T\geq 1$,
+$\sup_{t\in\{1,\dots,T\}}\|\hat{\mu}_{t}-\hat{\eta}_{t}\|\leq C\,\Delta_{T}\log\big(1+\Delta_{T}^{-1}\big),\quad\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\eta_{t}\|\leq C\,\Delta_{T}\log\big(1+\Delta_{T}^{-1}\big)$
+where $\Delta_{T}=\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\Phi(\mu_{t-1})\|$ with the convention that $\Phi(\mu_{0})=\eta_{1}$.
+Proof.
+See Section 5.4. ∎
+Remark 5.
+Let $(\mu_{t})_{t\geq 1}$ be a sequence in $\mathcal{P}(\mathbb{R})$ such that, for all $t\geq 1$, we have $\mu_{t+1}=\Phi_{t}(\mu_{t})$ for some mapping $\Phi_{t}:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$.
+Then, informally speaking, Lemma 1 implies that if $\Phi_{t}$ is close to $\Phi$ for all $t\in\{1,\dots,T\}$ then the estimation error $\|\mu_{t}-\eta_{t}\|$ will be small for all $t\in\{1,\dots,T\}$.
+This message conveyed by Lemma 1 is very intuitive, but, to the best of our knowledge, has not been proved to be true before.
+4.2 Proof of Theorem 1
+Proof.
+In what follows we let $T\geq 1$ be fixed and we use the convention that $\hat{\eta}^{N}_{0}M=\eta_{1}$.
+Then, by Lemma 1, to obtain the result of Theorem 1 it suffices to control the behaviour of $\sup_{t\in\{1\,\dots,T\}}\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|$ as a function of $N$.
+To this aim, remark first that, with $\mathbb{P}$-probability one and for all $t\geq 1$ and all $N\geq 1$, the function $F_{\hat{\eta}^{N}_{t-1}M}:\mathbb{R}\rightarrow(0,1)$ is continuous.
+Therefore, by the Dvoretzky-Kiefer-Wolfowitz inequality (see Massart,, 1990, Corollary 1), for all $(\gamma,\kappa)\in(0,1)$ we have
+$\sup_{t\in\{1,\dots,T\}}\mathbb{P}\Big(\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|\leq\kappa\big|\mathcal{F}_{t-1}^{N}\Big)\geq 1-\gamma,\quad\forall N\geq N_{\kappa,\gamma}:=\frac{\log(2/\gamma)}{2\kappa^{2}}$ (15)
+with the $\sigma$-algebra $\mathcal{F}_{t}^{N}$ as defined in the proof of Proposition 2 for all $t\geq 0$ and all $N\geq 1$. Using (15), we can conclude that
+$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|\leq\kappa\Big)\geq(1-\gamma)^{T},\quad\forall N\geq N_{\kappa,\gamma},\quad\forall(\gamma,\kappa)\in(0,1).$
+We now let $q\in(0,1)$. Then, by applying the latter result with $\gamma=1-(1-q)^{1/T}$, we obtain that
+$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|\leq\kappa\Big)\geq 1-q,\quad\forall N\geq N_{\kappa,1-(1-q)^{1/T}}.$ (16)
+By Bernoulli’s inequality (see e.g. Li and Yeh,, 2013), we have $1-(1-q)^{1/T}\geq T^{-1}q$ and thus
+$N_{\kappa,1-(1-q)^{1/T}}\leq\frac{\log(2)+\log(T/q)}{2\kappa^{2}}\leq\frac{\log(2)+\log(T/q)}{\kappa^{2}}.$
+Together with (16), this shows that
+$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\hat{\eta}_{t-1}M\|\leq N^{-1/2}\sqrt{\log(2)+\log(T/q)}\Big)\geq 1-q$
+and the result of Theorem 1 follows from Lemma 1. ∎
+4.3 Proof of Theorem 2
+4.3.1 Preliminary result: A discrepancy bound for sampling from univariate mixtures
+The proof of the following lemma is inspired by that of Hlawka and Mück, (1972, “Statz 2”).
+Lemma 2.
+Let $x_{1},\dots,x_{M}$ be $M$ points in $\mathbb{R}$ for some $M\geq 1$, and for all $x\in\{x_{1},\dots,x_{M}\}$ let $\mu_{x}\in\mathcal{P}(\mathbb{R})$.
+In addition, let $\pi_{M}(\mathrm{d}x)=\sum_{m=1}^{M}W_{m}\delta_{\{x_{m}\}}\in\mathcal{P}(\mathbb{R})$ and $\mu=\sum_{m=1}^{M}(W_{m}\delta_{\{x_{m}\}}\otimes\mu_{x_{m}})\in\mathcal{P}(\mathbb{R}^{2})$.
+Lastly, let $\varphi_{\mu}:(0,1)^{2}\rightarrow\mathbb{R}^{2}$ be defined by
+$\varphi_{\mu}(u)=\Big(z_{1},F^{-1}_{\mu_{x_{1}}}(u_{2})\Big),\quad z_{1}=F^{-1}_{\pi_{M}}(u_{1}),\quad u=(u_{1},u_{2})\in(0,1)^{2}$
+and let $\{u^{n}\}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ for some $N\geq 1$.
+Assume that, for all $a\in\mathbb{R}$, the mapping $\{x_{1},\dots,x_{M}\}\ni x\mapsto F_{\mu_{x}}(a)$ is monotone and let $\mathcal{B}_{2}=\big\{\prod_{i=1}^{2}(-\infty,a_{i}],\,(a_{1},a_{2})\in\mathbb{R}^{2}\big\}$.
+Then,
+$\sup_{B\in\mathcal{B}_{2}}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big(\varphi_{\mu}(u^{n})\big)-\mu(B)\Big|\leq\sqrt{96}\,D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)^{1/2}+12D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big).$ (17)
+Proof.
+Let $B=(-\infty,a_{1}]\times(-\infty,a_{2}]$ for some $(a_{1},a_{2})\in\mathbb{R}^{2}$.
+Without loss of generality we assume below that the set $\{x_{m}\}_{m=1}^{M}$ is labelled in such a way that $x_{1}\leq x_{2}\leq\dots\leq x_{M}$.
+In addition, we assume that $a_{1}\geq x_{1}$ since otherwise the set $B$ is such that
+$\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big(\varphi_{\mu}(u^{n})\big)-\mu(B)\Big|=0.$
+To prove the lemma we let $G(B)=\big\{u\in(0,1)^{2}:\varphi_{\mu}(u)\in B\big\}$ so that
+$\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big(\varphi_{\mu}(u^{n})\big)-\mu(B)=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{G(B)}(u^{n})-\lambda_{2}(G(B)).$ (18)
+Letting $j_{B}=\max\{i\in\{1,\dots,M\}:x_{i}\leq a_{1}\}$, and using the convention that $W_{0}=0$, we note that
+$G(B)=\bigcup_{i=1}^{j_{B}}\Big(W_{i-1},W_{i}\Big]\times\big(0,F_{\mu_{x_{i}}}(b_{2})\big]\subseteq(0,1]^{2}.$ (19)
+To proceed further we let $L\in\mathbb{N}$ and $\mathsf{P}$ be the partition of $(0,1]^{2}$ in congruent squares of side $\frac{1}{L}$ such that $W\in\mathsf{P}$ if and only if $W=(\xi_{1}/L,(\xi_{1}+1)/L]\times(\xi_{2}/L,(\xi_{2}+1)/L]$ for some $(\xi_{1},\xi_{2})\in\{0,\dots,L-1\}^{2}$.
+Then, we let
+$\mathsf{U}_{1}=\Big\{W\in\mathsf{P}:\,\,W\subset\mathring{G(B)}\Big\},\quad\mathsf{U}_{2}=\Big\{W\in\mathsf{P}:\,\,W\cap\partial(G(B))\neq\emptyset\Big\}$
+and we let $U_{1}=\cup_{W\in\mathsf{U}_{1}}W$ and $U_{2}=\cup_{W\in\mathsf{U}_{2}}W$ with the convention that $\cup_{W\in\mathsf{S}}=\emptyset$ if $\mathsf{S}=\emptyset$.
+With this notation in place, and letting $U_{1}^{\prime}=G(B)\setminus U_{1}$, we have
+$\begin{split}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{G(B)}(u^{n})-\lambda_{2}(G(B))&=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})+\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\end{split}$ (20)
+and we start by studying the first term on the r.h.s. of (20).
+To do so assume first that $\mathsf{U}_{1}\neq\emptyset$ and, for all $\xi_{1}\in\{0,\dots,L-1\}$, let
+$S_{\xi_{1}}=\Big\{k\in\{0,\dots,L-1\}:\,\,(\xi_{1}/L,(\xi_{1}+1)/L]\times(k/L,(k+1)/L]\in\mathsf{U}_{1}\Big\}$
+and let $I=\{\xi_{1}\in\{0,\dots,L-1\}:\,S_{\xi_{1}}\neq\emptyset\}$. Then, using the expression (19) for $G(B)$, it is direct to see that for all $\xi_{1}\in I$ the set
+$Q(\xi_{1}):=\bigcup_{k\in S_{\xi_{1}}}(\xi_{1}/L,(\xi_{1}+1)/L]\times(k/L,(k+1)/L]$
+is such that $Q(\xi_{1})=(a,b]\subset(0,1]^{2}$ for some $a,b\in[0,1]^{2}$. Therefore, noting that $\cup_{\xi_{1}\in I}Q(\xi_{1})=U_{1}$, we obtain that
+$\begin{split}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})\Big|\leq\sum_{\xi_{1}\in I}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{Q(\xi_{1})}(u^{n})-\lambda_{2}(Q(\xi_{1}))\Big|&\leq 4|I|D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)\\ &\leq 4L\,D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)\end{split}$ (21)
+where the second inequality uses the fact that, since $\{u^{n}\}_{n=1}^{N}$ is assumed to be a point set in $(0,1)^{2}$, we have (see Niederreiter,, 1992, Remarks 2-3 and Proposition 2.4, pages 14-15)
+$\sup_{(a,b]\subset(0,1]^{2}}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{(a,b]}(u^{n})-\lambda_{2}\big((a,b]\big)\Big|\leq 4D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big).$ (22)
+If $\mathsf{U}_{1}=\emptyset$ then $\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})=0$ and thus (21) also holds in this case.
+We now study the second term on the r.h.s. of (20). To this end we note first that we can cover $U_{1}^{\prime}$ with sets in $\mathsf{U}_{2}$, and thus
+$-\lambda_{2}(U_{2})\leq-\lambda_{2}(U_{1}^{\prime})\leq\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\leq\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{2}}(u^{n})-\lambda_{2}(U_{2})+\lambda_{2}(U_{2}).$
+By using this latter result, we obtain that
+$\begin{split}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\Big|&\leq\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{2}}(u^{n})-\lambda_{2}(U_{2})\Big|+\lambda_{2}(U_{2})\leq|\mathsf{U}_{2}|\Big(4D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)+L^{-2}\Big)\end{split}$ (23)
+where the second inequality holds by (22).
+Using (19) and the fact that, by assumption, the function $\{x_{1},\dots,x_{M}\}\ni x\mapsto F_{\mu_{x}}(b_{2})$ is either non-increasing or non-decreasing, it is readily checked that $|\mathsf{U}_{2}|\leq 2L$.
+Together with (18) and (20)-(23), this implies that
+$\sup_{B\in\mathcal{B}_{2}}\Big|\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big(\varphi_{\mu}(u^{n})\big)-\mu(B)\Big|\leq 12LD^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)+2L^{-1}$ (24)
+and the result of the lemma follows by applying (24) with $L=\big\lceil\big\{6D^{*}_{2}\big(\{u^{n}\}_{n=1}^{N}\big)\big\}^{-1/2}\big\rceil$. ∎
+4.3.2 Proof of the theorem
+Proof of Theorem 2.
+By Lemma 2, for all $N\geq 1$ and using the convention that $\hat{\eta}^{N}_{0}M=\eta_{1}$, we have $\sup_{t\geq 1}\|\eta_{t}^{N}-\hat{\eta}_{t-1}^{N}M\|\leq\sqrt{238}D^{*}_{2}\big(\{U_{t}^{n}\}_{n=1}^{N}\big)$ and the result of Theorem 2 then follows from Lemma 1 and (13).
+∎
+4.4 Proof of Corollary 1
+Proof.
+In what follows we let $(\kappa,q)\in(0,1)$ be arbitrary and $\delta_{N,T,q}=N^{-1}\big(1+\log(T/q)\big)$ for all integers $N\geq 1$ and $T\geq 1$.
+In addition, we let $C_{\kappa}\in(\exp(2),\infty)$ so that the inequality $\delta_{N,T,q}^{-1/2}\geq C_{\kappa}^{1/2}\kappa^{-1}\geq\exp(1)$ holds for all $N\geq C_{\kappa}\kappa^{-2}\big(1+\log(T/q)\big)$.
+Then, using the fact that function $f(x)=x^{-1}\log(1+x)$ is non-increasing on $x\in(\exp(1),\infty)$, it follows that
+$N$ $\geq C_{\kappa}\kappa^{-2}\big(1+\log(T/q)\big)\implies\delta_{N,T,q}^{1/2}\log\big(1+\delta_{N,T,q}^{-1/2}\big)\leq C_{\kappa}^{-1/2}\kappa\log\big(1+C_{\kappa}^{1/2}\kappa^{-1}\big)$
+where, with the constant $\bar{C}_{1}\in(0,1)$ as in the statement of Theorem 1,
+$\bar{C}_{1}C_{\kappa}^{-1/2}\kappa\log\big(1+C_{\kappa}^{1/2}\kappa^{-1}\big)\leq\kappa\Leftrightarrow\frac{\log\big(1+C_{\kappa}^{1/2}\kappa^{-1}\big)}{C_{\kappa}^{1/2}}\leq\frac{1}{\bar{C}_{1}}.$ (25)
+To proceed further we let $D\in(1,\infty)$ and $f_{D}:(0,\infty)\rightarrow\mathbb{R}$ be defined by
+$f_{D}(x)=\frac{\log\big(1+xDv(x)\big)}{Dv(x)},\quad v(x):=\exp(1)+\log(1+x),\quad x\in(0,\infty).$
+Simple calculations show that the function $x\mapsto\log(1+x)/\log(x)$ is non-increasing on the interval $(1,\infty)$.
+Therefore, noting that, since $D\geq 1$, we have $xDv(x)\geq\exp(1)>1$ for all $x\geq 1$, it follows that
+$\begin{split}f_{D}(x)=\frac{\log\big(xDv(x)\big)}{Dv(x)}\frac{\log\big(1+xDv(x)\big)}{\log\big(xDv(x\big))}\leq\frac{\log\big(xDv(x)\big)}{Dv(x)}\log\big(1+\exp(1)\big),\quad\forall x\geq 1\end{split}$ (26)
+where
+$\frac{\log\big(xDv(x)\big)}{Dv(x)}\leq\frac{1}{D}\bigg(1+\frac{\log(D)}{\exp(1)}+\sup_{z\geq 1}\frac{\log\big(v(x))}{v(x)}\bigg),\quad\forall x\geq 1.$ (27)
+Consequently, by (26)-(27), for any $\epsilon\in(0,\infty)$ there exists a constant $D_{\epsilon}\in(1,\infty)$ such that we have $\sup_{x\in[1,\infty]}f_{D}(x)\leq\epsilon$ for all $D\in[D_{\epsilon},\infty)$.
+By using this latter result, and assuming without loss of generality that $\bar{C}_{1}>1$, it follows that for any constant $D\geq D_{1/\bar{C}_{1}}\geq 1$ the second inequality in (25) holds for $C_{\kappa}=D^{2}\big(\exp(1)+\log(1+\kappa^{-1})\big)^{2}>\exp(2)$.
+The result of the corollary then follows from Theorem 1.
+∎
+4.5 Proof of Corollary 2
+Proof.
+The result of Corollary 2 easily follows from Theorem 2 and the calculations done in the proof of Corollary 1. The details are therefore omitted to save space. ∎
+5 Proof of Proposition 1 and proof of Lemma 1
+5.1 Additional notation
+For a function $f\in\mathcal{C}^{1}(\mathbb{R})$ we let $V(f)=\int_{\mathbb{R}}|f^{\prime}(x)|\mathrm{d}x$ with the convention that $V(f)=0$ if $f^{\prime}(x)=0$ for all $x\in\mathbb{R}$.
+Abusing notation, for a function $f:\mathbb{R}\rightarrow\mathbb{R}$ continuously differentiable on some non-empty interval $(a,b)$ we let $V(\mathbbm{1}_{(a,b]}f)=V(\mathbbm{1}_{(a,b)}f)=\int_{a}^{b}|f^{\prime}(x)|\mathrm{d}x$.
+Next, for all $(z,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $M_{r,s^{2}}(z,\mathrm{d}x)=\mathcal{N}(rz,s^{2})$ and $G_{s}(z)=e^{-sz^{2}/2}$, and below $\varphi(\cdot)$ denote the probability density function of the $\mathcal{N}(0,1)$ distribution.
+Moreover, for all $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $h_{a,r,s}:\mathbb{R}\rightarrow[-1,1]$ be defined by
+$h_{a,r,s}(x)=\Phi\bigg(\frac{a-rx}{s}\bigg)-\Phi\bigg(\frac{a}{s}\bigg),\quad x\in\mathbb{R}$ (28)
+while, for any bounded function $H:\mathbb{R}\rightarrow(0,\infty)$, we let $\Psi_{H}:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be defined by
 $\Psi_{H}(\pi)=\frac{H(x)\pi(\mathrm{d}x)}{\pi(H)},\quad\pi\in\mathcal{P}(\mathbb{R}).$
-Finally, we let $\varphi(\cdot)$ and $\Phi(\cdot)$ denote the probability and cumulative density function of the $\mathcal{N}_{1}(0,1)$ distribution, respectively, and for all $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$ we let $h_{a,r,s}:\mathbb{R}\rightarrow[-1,1]$ be defined by
-$h_{a,r,s}(x)=\Phi\bigg{(}\frac{a-rx}{s}\bigg{)}-\Phi\bigg{(}\frac{a}{s}\bigg{)},\quad x\in\mathbb{R}.$ (11)
-A.2 Preliminary results
-A.2.1 A simple but useful technical lemma
+5.2 Preliminary results
+5.2.1 A simple but useful technical lemma
 Lemma 3.
@@
 Proof.
-Let $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$. If $r=0$ the result of the lemma is trivial and we therefore assume that $r\neq 0$ in what follows. Then,
+Let $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$. If $r=0$ the result of the lemma is trivial and we therefore assume that $r\neq 0$ in what follows. In this scenario, we have
 $V(h_{a,r,s})=\int_{\mathbb{R}}|h^{\prime}_{a,r,s}(x)|\mathrm{d}x$ $=\int_{\mathbb{R}}\frac{|r|}{s}\varphi\Big(\frac{a-rx}{s}\bigg)\mathrm{d}x=\int_{\mathbb{R}}\frac{1}{s}\varphi\Big(\frac{a-z}{s}\Big)\mathrm{d}z=1$
-showing the first part of the lemma. On the other hand, using the mean value theorem and noting that $|\Phi^{\prime}(x)|=|\varphi(x)|\leq 1/\sqrt{2\pi}$ for all $x\in\mathbb{R}$, we have
+showing the first part of the lemma. On the other hand, using the mean value theorem and noting that $|\Phi^{\prime}(x)|=|\varphi(x)|\leq 1/\sqrt{2\pi}$ for all $x\in\mathbb{R}$, we obtain that
 $|h_{a,r,s}(x)|$ $=\bigg|\Phi\bigg(\frac{a-rx}{s}\bigg)-\Phi\bigg(\frac{a}{s}\bigg)\bigg|\leq\frac{|rx|}{s\sqrt{2\pi}},\quad\forall x\in\mathbb{R}$
 showing the second part of the lemma. The proof is complete. ∎
-A.2.2 A first useful property of the model of Section 2.1
+5.2.2 A general 1-dimensional Koksma-Hlawka inequality
+The next result is a general $s=1$ dimensional Koksma-Hlawka inequality, where the integration is w.r.t.
+an arbitrary distribution $\mu\in\mathcal{P}(\mathbb{R})$.
+By contrast, in the original Koksma-Hlawka inequality, the dimension $s\geq 1$ is arbitrary but the integration is w.r.t.
+the $\mathcal{U}(0,1)^{s}$ distribution (see e.g.
+Dick and Pillichshammer,, 2010, Proposition 2.18, page 33).
+It is also worth mentioning that Aistleitner and Dick, (2015) derived a version of the Koksma-Hlawka inequality where the integration is w.r.t.
+an arbitrary probability distribution on $(0,1)^{s}$ for some arbitrary $s\geq 1$.
+Lemma 4.
+Let $\infty\leq a<b\leq\infty$ and let $f:\mathbb{R}\rightarrow\mathbb{R}$ be continuously differentiable on the interval $(a,b)$ and such that $V(f\mathbbm{1}_{(a,b)})<\infty$, and let $\pi,\mu\in\mathcal{P}(\mathbb{R})$.
+Then, with the convention that $f(a)=0$ when $a=$-$\infty$ and with the convention that $f(b)=0$ when $b=\infty$, we have
+$\begin{split}\Big|\int_{a}^{b}f(x)\big(\pi(\mathrm{d}x)-\mu(\mathrm{d}x)\big)\Big|\leq\Big(|f(a)|+|f(b)|+V(f\mathbbm{1}_{(a,b)})\Big)\|\pi-\mu\|.\end{split}$
+Proof.
+Assume first $(a,b)\in\mathbb{R}^{2}$ and note that, since the function $f$ is continuous, we have $\int_{a}^{b}f(x)\mathrm{d}\big(F_{\pi}(x)-F_{\mu}(x)\big)=\int_{a}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))$.
+Then, using the integration by part formula for Riemann-Stieltjes integrals (see Apostol,, 1973, Theorem 7.6, page 144), we have
+$\begin{split}\int_{a}^{b}\big(F_{\pi}(x)-F_{\mu}(x)\big)f^{\prime}(x)\mathrm{d}x&=\int_{a}^{b}\big(F_{\pi}(x)-F_{\mu}(x)\big)\mathrm{d}f(x)\\ &=f(b)\big(F_{\pi}(b)-F_{\mu}(b)\big)-f(a)\big(F_{\pi}(a)-F_{\mu}(a)\big)-\int_{a}^{b}f(x)\mathrm{d}\big(F_{\pi}(x)-F_{\mu}(x)\big)\end{split}$
+from which we deduce that
+$\begin{split}\Big|\int_{a}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))\Big|&\leq|f(b)|\,\big|F_{\pi}(b)-F_{\mu}(b)\big|+|f(a)|\,\big|F_{\pi}(a)-F_{\mu}(a)\big|\\ &+\Big|\int_{a}^{b}\big(F_{\pi}(x)-F_{\mu}(x)\big)f^{\prime}(x)\mathrm{d}x\Big|.\end{split}$ (29)
+The result of the lemma when $(a,b)\in\mathbb{R}^{2}$ directly follows from (29).
+Assume now that $a=$-$\infty$ and that $b\in\mathbb{R}$.
+Note that under the assumptions of the lemma we have $\sup_{x\in\mathbb{R}}|f(x)|<\infty$ while, by using the reverse Fatou’s lemma, we obtain that $\lim_{x\rightarrow-\infty}F_{\pi}(x)=\lim_{a\rightarrow-\infty}F_{\mu}(x)=0$.
+Therefore, using (29),
+$\Big|\int_{-\infty}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))\Big|$ $=\Big|\lim_{a\rightarrow-\infty}\int_{a}^{b}f(x)\mathrm{d}\big(F_{\pi}(x)-F_{\mu}(x)\big)\Big|$
+$\leq\big|f(b)|\,\big|F_{\pi}(b)-F_{\mu}(b)|$
+$+\limsup_{a\rightarrow-\infty}\Big|\int_{a}^{b}\big(F_{\pi}(x)-F_{\mu}(x)\big)f^{\prime}(x)\mathrm{d}x\Big|$
+$\leq\|\pi-\mu\|\Big(|f(b)|+V(\mathbbm{1}_{(-\infty,b)}f)\Big)$
+proving the result of the lemma when $a=$-$\infty$ and $b\in\mathbb{R}$.
+Using Fatou’s lemma we have $\lim_{x\rightarrow\infty}F_{\pi}(x)=\lim_{a\rightarrow\infty}F_{\mu}(x)=1$, and the result of the lemma for the two remaining case (i.e.
+for $b=\infty$ and $a\in\mathbb{R}\cup\{-\infty\}$) can proved using similar calculations as above, and is therefore omitted to save space.
+∎
+5.2.3 A useful property of the model of Section 2.1
 Lemma 5.
 Let $c_{0}=\sigma_{0}=0$, $\rho_{0}=1$, and let $(c_{k})_{k\geq 1}$, $(\rho_{k})_{k\geq 1}$ and $(\sigma_{k})_{k\geq 1}$ be such that
-$c_{k}=\frac{(c+c_{k-1})\rho^{2}}{1+(c+c_{k-1})\sigma^{2}},\quad\rho_{k}=\prod_{i=0}^{k-1}\frac{\rho}{1+(c+c_{i})\sigma^{2}},\quad\sigma_{k}^{2}=\sigma_{k-1}^{2}+\frac{\rho_{k-1}^{2}\sigma^{2}}{1+(c+c_{k-1})\sigma^{2}},\quad\forall k\geq 1.$
-Then, for all $k\geq 1$, we have
+$c_{k}=\frac{(c+c_{k-1})\rho^{2}}{1+(c+c_{k-1})\sigma^{2}},\quad\rho_{k}=\prod_{i=0}^{k-1}\frac{\rho}{1+(c+c_{i})\sigma^{2}},\quad\sigma_{k}^{2}=\sigma_{k-1}^{2}+\frac{\rho_{k-1}^{2}\sigma^{2}}{1+(c+c_{k-1})\sigma^{2}}$
+for all $k\geq 1$. Then, for all $k\geq 1$, we have
 $Q^{k}(x^{\prime},\mathrm{d}x)\propto G_{c_{k}}(x^{\prime})M_{\rho_{k},\sigma^{2}_{k}}(x^{\prime},\mathrm{d}x),\quad\forall x^{\prime}\in\mathbb{R}.$ (30)
@@
 Proof.
-The result in the first part of the lemma can be deduced from the calculations in Del Moral and Jacod, (2001) while the convergence result for the sequence $(\sigma_{k})_{k\geq 1}$ can be directly obtained from Del Moral and Horton, (2023, Theorem 3.5).
-However, for sake of completeness the whole lemma is proved in what follows.
+The result stated in (30) can be deduced from the calculations in Del Moral and Jacod, (2001) while the convergence result for the sequence $(\sigma_{k})_{k\geq 1}$ can be directly obtained from Del Moral and Horton, (2023, Theorem 3.5).
+However, for sake of completeness, a proof of the whole lemma is presented below.
 To prove the first part of the lemma remark first that, for all $a\in(0,\infty)$, we have
@@
 so that $c_{k}=f(c_{k-1})$ for all $k\geq 1$.
-Easy computations show that the function $f$ is non-decreasing on $[0,\infty)$, and thus the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and such that $c\rho^{2}/(1+c\sigma^{2})\leq c_{k}\leq\rho^{2}/\sigma^{2}$ for all $k\geq 1$.
+Easy computations show that the function $f$ is non-decreasing on $0,\infty)$, and thus the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and such that $c\rho^{2}/(1+c\sigma^{2})\leq c_{k}\leq\rho^{2}/\sigma^{2}$ for all $k\geq 1$.
 To study the sequence $(\rho_{k})_{k\geq 0}$ assume first that $|\rho|<1+c\sigma^{2}$.
@@
 $\frac{|\rho|}{1+(c+c_{k})\sigma^{2}}<1$ $\Leftrightarrow c_{k}>\frac{|\rho|-(1+c\sigma^{2})}{\sigma^{2}}.$ (33)
-To proceed further we remark that, since the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and bounded, the sequence converges in $[0,\infty)$ and we let $c_{\infty}=\lim_{k\rightarrow\infty}c_{k}$.
+To proceed further we remark that, since the sequence $(c_{k})_{k\geq 0}$ is non-decreasing and bounded, the sequence converges in $0,\infty)$ and we let $c_{\infty}=\lim_{k\rightarrow\infty}c_{k}$.
 In addition, simple calculations show that
 $c_{\star}:=\frac{\rho^{2}-(1+c\sigma^{2})+\sqrt{(1+c\sigma^{2}-\rho^{2})^{2}+4c\sigma^{2}\rho^{2}}}{2\sigma^{2}}$
-is the unique value of $x\in[0,\infty)$ such that $x=f(x)$, and since we are assuming that $|\rho|\geq 1+c\sigma^{2}>1$ the constant $c_{\star}$ is such that
-$c_{\star}\geq\frac{\rho^{2}-(1+c\sigma^{2})+\sqrt{(1+c\sigma^{2}-\rho^{2})^{2}}}{2\sigma^{2}}=\frac{\rho^{2}-(1+c\sigma^{2})}{\sigma^{2}}>\frac{\rho-(1+c\sigma^{2})}{\sigma^{2}}.$ (16)
-On the other hand, since $f$ is continuous on $[0,\infty)$ we have
+is the unique value of $x\in[0,\infty)$ such that $x=f(x)$.
+Moreover, since we are assuming that $|\rho|\geq 1+c\sigma^{2}>1$, it follows that $\rho^{2}-(1+c\sigma^{2})\geq 0$ and thus the constant $c_{\star}$ is such that
+$c_{\star}\geq\frac{\rho^{2}-(1+c\sigma^{2})+\sqrt{(1+c\sigma^{2}-\rho^{2})^{2}}}{2\sigma^{2}}=\frac{\rho^{2}-(1+c\sigma^{2})}{\sigma^{2}}>\frac{|\rho|-(1+c\sigma^{2})}{\sigma^{2}}.$ (34)
+On the other hand, since the function $f$ is continuous on $0,\infty)$ we have
 $c_{\infty}=\lim_{k\rightarrow\infty}c_{k+1}=\lim_{k\rightarrow\infty}f(c_{k})=f(\lim_{k\rightarrow\infty}c_{k})=f(c_{\infty})$
-showing that $c_{\infty}=c_{\star}$, and thus $\lim_{k\rightarrow\infty}c_{k}=c_{\star}$. Together with (16), this implies that there exists a $k_{\star}\in\mathbb{N}$ such that
-$c_{k_{\star}}>\frac{|\rho|-(1+c\sigma^{2})}{\sigma^{2}}$
-and thus, using (15) and the fact that the sequence $(c_{k})_{k\geq 0}$ is non-decreasing, we have
+showing that $c_{\infty}=c_{\star}$, and thus $\lim_{k\rightarrow\infty}c_{k}=c_{\star}$.
+Together with (34), this implies that there exists a $k_{\star}\in\mathbb{N}$ such that $c_{k_{\star}}>(|\rho|-(1+c\sigma^{2}))/\sigma^{2}$, and thus, using (33) and the fact that the sequence $(c_{k})_{k\geq 0}$ is non-decreasing, we have
 $\frac{|\rho|}{1+(c+c_{k})\sigma^{2}}\leq\epsilon_{\star}:=\frac{|\rho|}{1+(c+c_{k_{\star}})\sigma^{2}}<1,\quad\forall k\geq k_{\star}$
@@
 and thus, for all integers $1\leq k<m$ we have, with $C$ and $\epsilon_{\star}$ as above,
-$|\sigma_{m}^{2}-\sigma_{k}^{2}|=\Big{|}\sum_{s=k}^{m-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}\Big{|}\leq\frac{\sigma^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\rho_{s}^{2}\leq\frac{\sigma^{2}C^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\epsilon_{\star}^{2s}\leq\epsilon_{\star}^{2k}\frac{\sigma^{2}C^{2}}{(1-\epsilon_{\star})(1+c\sigma^{2})}.$ (17)
-Using (17) it is easily verified that the sequence $(\sigma^{2}_{k})_{k\geq 1}$ is Cauchy and therefore converges in $[0,\infty)$, and we let $\sigma^{2}_{\infty}=\lim_{k\rightarrow\infty}\sum_{s=0}^{k-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}$.
+$\begin{split}|\sigma_{m}^{2}-\sigma_{k}^{2}|=\Big|\sum_{s=k}^{m-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}\Big|\leq\frac{\sigma^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\rho_{s}^{2}&\leq\frac{\sigma^{2}C^{2}}{1+c\sigma^{2}}\sum_{s=k}^{m-1}\epsilon_{\star}^{2s}\leq\epsilon_{\star}^{2k}\frac{\sigma^{2}C^{2}}{(1-\epsilon_{\star})(1+c\sigma^{2})}.\end{split}$ (35)
+Using (35) it is easily verified that the sequence $(\sigma^{2}_{k})_{k\geq 1}$ is Cauchy and therefore converges in $0,\infty)$, and we let $\sigma^{2}_{\infty}=\lim_{k\rightarrow\infty}\sum_{s=0}^{k-1}\frac{\rho_{s}^{2}\sigma^{2}}{1+(c+c_{s})\sigma^{2}}$.
 Remark that since the sequence $(\sigma^{2}_{k})_{k\geq 1}$ is non-decreasing and such that $\sigma^{2}_{k}\geq\sigma^{2}/(1+c\sigma^{2})$ it follows that $\sigma^{2}_{\infty}>0$, and remark that, by (35), we have
@@
 The proof of the lemma is complete. ∎
-A.2.3 A second useful property of the model of Section 2.1
-The following lemma will be used to prove Lemma 2:
-Lemma 6.
-There exist constants $(\varrho_{1},\varrho_{2})\in(0,\infty)^{2}$ such that, for any function $f:\mathbb{R}\rightarrow\mathbb{R}$ verifying $\sup_{x\in\mathbb{R}}|f(x)|e^{-\varrho_{1}x^{2}}<\infty$, there is a constant $C_{f}\in(0,\infty)$ for which we have
-$\sup_{t\geq 1}\big{|}\hat{\eta}_{t}(f\mathbbm{1}_{[-b,b]^{c}})\big{|}\leq C_{f}e^{-\varrho_{2}b^{2}},\quad\sup_{t\geq 1}\big{|}\eta_{t}(f\mathbbm{1}_{[-b,b]^{c}})\leq C_{f}e^{-\varrho_{2}b^{2}},\quad\forall b\in(0,\infty).$
-Proof.
-We start with some preliminary computations. Let $b\in[2,\infty)$, $(a,r)\in\mathbb{R}^{2}$, $s\in(0,\infty)$,
-$\bar{\delta}_{a,r,s^{2}}=\min\bigg{\{}\frac{1}{s^{2}},\frac{a}{as^{2}+r^{2}}\bigg{\}}$ (18)
-and $\Psi(\cdot;a,s^{2})$ denote the cumulative density function of the $\mathcal{N}_{1}(a,s^{2})$ distribution.
-In addition, for all $\delta\in(0,\infty)$ and function $f:\mathbb{R}\rightarrow\mathbb{R}$ we let $C_{f,\delta}=\sup_{x\in\mathbb{R}}|f(x)|e^{-(\delta/2)x^{2}}\in(0,\infty]$.
-We now let $\delta\in(0,\bar{\delta}_{a,r,s^{2}})$ be fixed and $f:\mathbb{R}\rightarrow\mathbb{R}$ be such that $C_{f,\delta}<\infty$. Then, for all $x\in\mathbb{R}$ we have
-$\begin{split}\big{|}G_{a}(x)M_{r,s^{2}}\big{(}x,f\mathbbm{1}_{[-b,b]^{c}}\big{)}\big{|}&\leq\frac{C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{x^{2}}{2}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}M_{\frac{r}{1-\delta s^{2}},\frac{s^{2}}{1-\delta s^{2}}}\big{(}x,[-b,b]^{c}\big{)}\\ &\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{x^{2}}{2}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}\Phi\Big{(}(b,\infty);\frac{|rx|}{1-\delta s^{2}},\frac{s^{2}}{1-\delta s^{2}}\Big{)}.\end{split}$ (19)
-To proceed further let $x\in\mathbb{R}$ be such that $b\geq 1+3|rx|/(1-\delta s^{2})$. Then, recalling that for all $z>0$ we have $\Phi\big{(}(z,\infty);0,1\big{)}\leq e^{-z^{2}/2}/z$, it follows that
-$\Phi\Big{(}(b,\infty);\frac{|rx|}{1-\delta s^{2}},\frac{s^{2}}{1-\delta s^{2}}\Big{)}\leq\exp\bigg{(}-\frac{1-\delta s^{2}}{2s^{2}}\Big{(}b-\frac{|rx|}{1-\delta s^{2}}\Big{)}^{2}\bigg{)}\frac{s}{(1-\delta s^{2})^{1/2}}$
-which, together with (19), implies that
-$|G_{a}(x)M_{r,s^{2}}(x,f\mathbbm{1}_{[-b,b]^{c}})|$ $\leq\frac{2sC_{f,\delta}}{1-\delta s^{2}}\exp\Big{(}-\alpha b^{2}-\beta x^{2}+\gamma b|x|\Big{)}$ (20)
-with
-$\alpha=\frac{1-\delta s^{2}}{2s^{2}},\quad\beta=\frac{1}{2}\Big{(}\frac{r^{2}}{s^{2}}+a\Big{)},\quad\gamma=\frac{|r|}{s^{2}}.$
-If $r\neq 0$ we have, recalling that we are assuming that $x$ is such that $b\geq 1+3|rx|/(1-\delta s^{2})$,
-$-\alpha b^{2}+\gamma b|x|\leq-\alpha b^{2}+\gamma\frac{1-\delta s^{2}}{3|r|}b^{2}$ $=-\frac{1-\delta s^{2}}{2s^{2}}b^{2}+\frac{1-\delta s^{2}}{3s^{2}}b^{2}=-\frac{1-\delta s^{2}}{6s^{2}}b^{2}$
-while if $r=0$ we have $-\alpha b^{2}+\gamma b|x|=-\alpha b^{2}\leq-(1-\delta s^{2})/(6s^{2})$. Together with (20), this shows that
-$b\geq 1+3|rx|/(1-\delta s^{2})\implies\big{|}G_{a}(x)M_{r,s^{2}}\big{(}x,f\mathbbm{1}_{[-b,b]^{c}}\big{)}\big{|}$ $\leq\frac{2sC_{f,\delta}}{1-\delta s^{2}}\exp\Big{(}-\frac{1-\delta s^{2}}{6s^{2}}b^{2}\Big{)}.$ (21)
-Assume now that $x$ is such that $b<1+3|rx|/(1-\delta s^{2})$, so that $x^{2}>(1-\delta s^{2})^{2}(b-1)^{2}/(9r^{2})$.
-(Remark that since $b\geq 2$ we can have $b<1+3|rx|/(1-\delta s^{2})$ only if $r\neq 0$).
-Then, noting that since $\delta\in(0,\bar{\delta}_{a,r,s^{2}})$ we have $a-\delta r^{2}/(1-\delta s^{2})>0$, it directly follows from (19) that
-$\begin{split}b<1+3|rx|/(1-\delta s^{2})\implies\big{|}G_{a}(x)&M_{r,s^{2}}\big{(}x,f\mathbbm{1}_{[-b,b]^{c}}\big{)}\big{|}\\ &\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{x^{2}}{2}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}\\ &\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{(b-1)^{2}(1-\delta s^{2})^{2}}{18r^{2}}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}\\ &\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\exp\bigg{(}-\frac{b^{2}(1-\delta s^{2})^{2}}{36r^{2}}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{)}\end{split}$ (22)
-where the last inequality uses the fact that $(b-1)\geq b/2$ since $b\geq 2$.
-Then, by using (21) and (22), it follows that
-$\big{|}G_{a}(x)M_{r,s^{2}}\big{(}x,f\mathbbm{1}_{[-b,b]^{c}}\big{)}\big{|}\leq\frac{2C_{f,\delta}}{(1-\delta s^{2})^{1/2}}\bigg{(}1+\frac{s}{(1-\delta s^{2})^{1/2}}\bigg{)}e^{-b^{2}K(a,r,s^{2},\delta)},\quad\forall x\in\mathbb{R}$ (23)
-with
-$K(a,r,s^{2},\delta)=\min\bigg{\{}\frac{1-\delta s^{2}}{6s^{2}},\frac{(1-\delta s^{2})^{2}}{36r^{2}}\Big{(}a-\frac{\delta r^{2}}{1-\delta s^{2}}\Big{)}\bigg{\}}.$
-We are now in position to prove the lemma.
-To show the first part of the lemma let $t\geq 2$, $b\in[2,\infty)$, and let $(c_{k})_{k\geq 0}$, $(\rho_{k})_{k\geq 0}$ and $(\sigma_{k})_{k\geq 0}$ be as defined in Lemma 5.
-Then, by using (4) and Lemma 5, for any function $f:\mathbb{R}\rightarrow\mathbb{R}$ we have
-$\hat{\eta}_{t}\big{(}f\mathbbm{1}_{[-b,b]^{c}}\big{)}=\frac{\int_{\mathbb{R}}G_{c_{t-1}}(x)M_{\rho_{t-1},\sigma^{2}_{t-1}}(x,f\mathbbm{1}_{[-b,b]^{c}})\hat{\eta}_{1}(\mathrm{d}x)}{\hat{\eta}_{1}(G_{c_{t-1}})}.$ (24)
-By Lemma 5, we have $c_{t-1}\geq c_{1}$, $|\rho_{t-1}|\leq|\rho|$ and $\sigma_{t-1}^{2}\leq\sigma^{2}_{\infty}$, with $\sigma^{2}_{\infty}\in(0,\infty)$ as in Lemma 5. Therefore, letting
-$\bar{\delta}^{\prime}=\min\bigg{\{}\frac{1}{\sigma^{2}_{\infty}},\frac{c_{1}}{2\rho^{2}+c_{1}\sigma^{2}_{\infty}}\bigg{\}}$
-it follows that
-$c_{t-1}-\frac{\delta^{\prime}\rho_{t-1}^{2}}{1-\delta^{\prime}\sigma_{t-1}^{2}}\geq\frac{c_{t-1}}{2}\geq\frac{c_{1}}{2},\quad\forall\delta^{\prime}\in(0,\bar{\delta}^{\prime}).$ (25)
-We now let $\delta^{\prime}\in(0,\bar{\delta^{\prime}})$ be fixed and $f:\mathbb{R}\rightarrow\mathbb{R}$ be such that $C_{f,\delta^{\prime}}<\infty$.
-Then, noting that $\delta^{\prime}<\bar{\delta}^{\prime}\leq\bar{\delta}_{c_{t-1},\rho_{t-1},\sigma^{2}_{t-1}}$ (with $\bar{\delta}_{a,r,s^{2}}$ defined in (18) for any $(a,r)\in\mathbb{R}^{2}$ and $s\in(0,\infty)$), by applying (23) with $(a,r,s^{2},\delta)=(c_{t-1},\rho_{t-1},\sigma^{2}_{t-1},\delta^{\prime})$ we have
-$\begin{split}|G_{c_{t-1}}(x)M_{\rho_{t-1},\sigma^{2}_{t-1}}(x,f\mathbbm{1}_{[-b,b]^{c}})|&\leq\frac{2C_{f,\delta^{\prime}}}{(1-\delta^{\prime}\sigma_{t-1}^{2})^{1/2}}\bigg{(}1+\frac{\sigma_{t-1}}{(1-\delta^{\prime}\sigma_{t-1}^{2})^{1/2}}\bigg{)}e^{-K_{t-1}b^{2}}\\ &\leq\frac{2C_{f,\delta^{\prime}}}{(1-\delta^{\prime}\sigma_{\infty}^{2})^{1/2}}\bigg{(}1+\frac{\sigma_{\infty}}{(1-\delta^{\prime}\sigma_{\infty}^{2})^{1/2}}\bigg{)}e^{-K_{t-1}b^{2}}\quad\forall x\in\mathbb{R}\end{split}$ (26)
-with
-$\begin{split}K_{t-1}&=\min\bigg{\{}\frac{1-\delta^{\prime}\sigma_{t-1}^{2}}{6\sigma_{t-1}^{2}},\frac{(1-\delta^{\prime}\sigma_{t-1}^{2})^{2}}{36\rho_{t-1}^{2}}\Big{(}c_{t-1}-\frac{\delta^{\prime}\rho_{t-1}^{2}}{1-\delta^{\prime}\sigma_{t-1}^{2}}\Big{)}\bigg{\}}\\ &\geq\varrho:=\min\bigg{\{}\frac{1-\delta^{\prime}\sigma_{\infty}^{2}}{6\sigma_{\infty}^{2}},\frac{(1-\delta^{\prime}\sigma_{\infty}^{2})^{2}}{36\rho^{2}}\frac{c_{1}}{2}\bigg{\}}>0\end{split}$ (27)
-and where the first inequality holds by (25).
-On the other hand, by Lemma 5 we have $c_{t-1}\leq\rho^{2}/\sigma^{2}$ and thus
-$\hat{\eta}_{1}(G_{c_{t-1}})\geq\hat{\eta}_{1}(G_{\rho^{2}/\sigma^{2}})=\frac{\eta_{1}(G_{c+\rho^{2}/\sigma^{2}})}{\eta_{1}(G_{c})}.$
-By combining this latter result with (24), (26) and (27), we obtain that
-$|\hat{\eta}_{t}(f\mathbbm{1}_{[-b,b]^{c}})|$ $\leq e^{-\varrho b^{2}}\frac{2C_{f,\delta^{\prime}}}{\eta_{1}(G_{c+\rho^{2}/\sigma^{2}})(1-\delta^{\prime}\sigma_{\infty}^{2})^{1/2}}\bigg{(}1+\frac{\sigma_{\infty}}{(1-\delta^{\prime}\sigma_{\infty}^{2})^{1/2}}\bigg{)},\quad\forall t\geq 2,\quad\forall b\in[2,\infty)$
-showing the first part of the lemma.
-The prove the second part of the lemma let $f:\mathbb{R}\rightarrow\mathbb{R}$ and note that, for any $b\in(0,\infty)$ we have, using (4) and Lemma 5,
-$\eta_{t+1}(f\mathbbm{1}_{[-b,b]^{c}})=\frac{\int_{\mathbb{R}}G_{c_{t-1}}(x)M_{\rho\rho_{t-1},\sigma^{2}+\rho\sigma^{2}_{t-1}}(x,f\mathbbm{1}_{[-b,b]^{c}})\hat{\eta}_{1}(\mathrm{d}x)}{\hat{\eta}_{1}(G_{c_{t-1}})},\quad\forall t\geq 1.$
-Then, the second part of the lemma readily follows from the computations used to show the first part of the lemma, and the proof of the lemma is complete. ∎
-A.2.4 A general 1-dimensional Koksma-Hlawka inequality
-The next result is a general $s=1$ dimensional Koksma-Hlawka inequality, where the integration is w.r.t.
-an arbitrary distribution $\mu\in\mathcal{P}(\mathbb{R})$.
-By contrast, in the original Koksma-Hlawka inequality $s\in\mathbb{N}$ is arbitrary but the integration is w.r.t.
-the $\mathcal{U}(0,1)^{s}$ distribution (see e.g.
-Dick and Pillichshammer,, 2010, Proposition 2.18, page 33).
-It is also worth mentioning that Aistleitner and Dick, (2015) derived a version of the Koksma-Hlawka inequality where the integration is w.r.t.
-an arbitrary probability distribution on $(0,1)^{s}$ for some $s\in\mathbb{N}$.
-Lemma 7.
-Let $-\infty\leq a<b\leq\infty$ and let $f:\mathbb{R}\rightarrow\mathbb{R}$ be continuously differentiable on the interval $(a,b)$ and such that $V(f\mathbbm{1}_{(a,b)})<\infty$, and let $\pi,\mu\in\mathcal{P}(\mathbb{R})$.
-Then, with the convention that $f(a)=0$ when $a=-\infty$ and with the convention that $f(b)=0$ when $b=\infty$, we have
-$\begin{split}\Big{|}\int_{a}^{b}f(x)\big{(}\pi(\mathrm{d}x)-\mu(\mathrm{d}x)\big{)}\Big{|}\leq\Big{(}|f(a)|+|f(b)|+V(f\mathbbm{1}_{(a,b)})\Big{)}\|\pi-\mu\|.\end{split}$
-Proof.
-Assume first $(a,b)\in\mathbb{R}^{2}$. Then, using the integration by part formula for Riemann-Stieltjes integrals (see Apostol,, 1973, Theorem 7.6, page 144), we have
-$\begin{split}\int_{a}^{b}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}f^{\prime}(x)\mathrm{d}x&=\int_{a}^{b}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}\mathrm{d}f(x)\\ &=f(b)\big{(}F_{\pi}(b)-F_{\mu}(b)\big{)}-f(a)\big{(}F_{\pi}(a)-F_{\mu}(a)\big{)}\\ &-\int_{a}^{b}f(x)\mathrm{d}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}\end{split}$ (28)
-and the result of the lemma follows upon noting that $\int_{a}^{b}f(x)\mathrm{d}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}=\int_{a}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))$.
-Assume now that $a=-\infty$ and that $b\in\mathbb{R}$, and note that under the assumptions of the lemma we have $\sup_{x\in\mathbb{R}}|f(x)|<\infty$. Then, by (28),
-$\bigg{|}\int_{-\infty}^{b}f(x)(\pi(\mathrm{d}x)-\mu(\mathrm{d}x))\bigg{|}$ $=\bigg{|}\lim_{a\rightarrow-\infty}\int_{a}^{b}f(x)\mathrm{d}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}\bigg{|}$
-$=\bigg{|}f(b)\big{(}F_{\pi}(b)-F_{\mu}(b)\big{)}-\lim_{a\rightarrow-\infty}\int_{a}^{b}\big{(}F_{\pi}(x)-F_{\mu}(x)\big{)}f^{\prime}(x)\mathrm{d}x\bigg{|}$
-$\leq\|\pi-\mu\|\big{(}|f(b)|+V(\mathbbm{1}_{(-\infty,b)}f)\big{)}$
-proving the result of the lemma when $a=-\infty$ and $b\in\mathbb{R}$.
-The proof of the result of the lemma for the two remaining case (i.e.
-for $b=\infty$ and $a\in\mathbb{R}\cup\{-\infty\}$) is similar and is therefore omitted to save space.
-∎
-A.2.5 A discrepancy bound for 1-dimensional importance sampling
-From Lemma 7 we readily obtain the following corollary:
-Corollary 1.
-Let $H\in\mathcal{C}^{1}(\mathbb{R})$ be such that $V(H)<\infty$ and such that $H(x)\geq 0$ for all $x\in\mathbb{R}$, and let $\mu,\pi\in\mathcal{P}(\mathbb{R})$. Then,
-$\|\Psi_{H}(\mu)-\Psi_{H}(\pi)\|\leq\frac{\|H\|_{\infty}+2V(H)}{\pi(H)}\|\mu-\pi\|.$
-Proof.
-The result follows from Lemma 7 and the fact that, for all $B\in\mathcal{B}_{1}$, we have
-$|\Psi_{H}(\mu)(B)-\Psi_{H}(\pi)(B)|\leq\frac{|(\mu-\pi)(H\mathbbm{1}_{B})|+|(\mu-\pi)(H)|}{\pi(H)}.$
-A.2.6 A discrepancy bound for sampling from univariate mixtures
-The following lemma will be used to study jointly the resampling and mutation steps of Algorithm 1. Its proof is inspired by that of Hlawka and Mück, (1972, “Statz 2”).
-Lemma 8.
-Let $M\in\mathbb{N}$, $\pi_{M}(\mathrm{d}x)=\sum_{m=1}^{M}W_{m}\delta_{\{x_{m}\}}\in\mathcal{P}(\mathbb{R})$ and, for all $x\in\{x_{1},\dots,x_{M}\}$, let $\mu_{x}\in\mathcal{P}(\mathbb{R})$.
-In addition, let $\mu=\sum_{m=1}^{M}W_{m}\delta_{\{x_{m}\}}\otimes\mu_{x_{m}}\in\mathcal{P}(\mathbb{R}^{2})$ and $\varphi_{\mu}:(0,1)^{2}\rightarrow\mathbb{R}^{2}$ be defined by
-$\varphi_{\mu}(u)=\Big{(}x_{1},F^{-1}_{\mu_{x_{1}}}(u_{2})\Big{)},\quad x_{1}=F^{-1}_{\pi_{M}}(u_{1}),\quad u=(u_{1},u_{2})\in(0,1)^{2}.$
-Finally, let $N\geq 1$, $\{u^{n}\}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ and assume that for all $a\in\mathbb{R}$ the mapping $\{x_{1},\dots,x_{M}\}\ni x\mapsto F_{\mu_{x}}(a)$ is monotone. Then,
-$\sup_{B\in\mathcal{B}_{2}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big{(}\varphi_{\mu}(u^{n})\big{)}-\mu(B)\Big{|}\leq\sqrt{96}\,D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}^{1/2}+12D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}.$ (29)
-Proof.
-Let $B=(-\infty,a_{1}]\times(-\infty,a_{2}]$ for some $(a_{1},a_{2})\in\mathbb{R}^{2}$.
-Without loss of generality we assume below that the set $\{x_{m}\}_{m=1}^{M}$ is labelled in such a way that $x_{1}\leq x_{2}\leq\dots\leq x_{M}$.
-In addition, we assume that $a_{1}\geq x_{1}$ since otherwise the set $B$ is such that
-$\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big{(}\varphi_{\mu}(u^{n})\big{)}-\mu(B)\Big{|}=0.$
-To prove the lemma we let $G(B)=\big{\{}u\in(0,1)^{2}:\varphi_{\mu}(u)\in B\big{\}}$ so that
-$\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big{(}\varphi_{\mu}(u^{n})\big{)}-\mu(B)=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{G(B)}(u^{n})-\lambda_{2}(G(B)).$ (30)
-Letting $j_{B}=\max\{i\in\{1,\dots,M\}:x_{i}\leq a_{1}\}$, and using the convention that $W_{i-1}=0$ for $i=1$, we note that
-$G(B)=\bigcup_{i=1}^{j_{B}}\Big{(}W_{i-1},W_{i}\Big{]}\times\big{(}0,F_{\mu_{x_{i}}}(b_{2})\big{]}\subseteq(0,1]^{2}.$ (31)
-To proceed further we let $L\in\mathbb{N}$ and $\mathsf{P}$ be the partition of $(0,1]^{2}$ in congruent squares of side $\frac{1}{L}$ such that $W\in\mathsf{P}$ if and only if $W=(\xi_{1}/L,(\xi_{1}+1)/L]\times(\xi_{2}/L,(\xi_{2}+1)/L]$ for some $(\xi_{1},\xi_{2})\in\{0,\dots,L-1\}^{2}$.
-Then, we let
-$\mathsf{U}_{1}=\Big{\{}W\in\mathsf{P}:\,\,W\subset\mathring{G(B)}\Big{\}},\quad\mathsf{U}_{2}=\Big{\{}W\in\mathsf{P}:\,\,W\cap\partial(G(B))\neq\emptyset\Big{\}}$
-and we let $U_{1}=\cup_{W\in\mathsf{U}_{1}}W$ and $U_{2}=\cup_{W\in\mathsf{U}_{2}}W$ with the convention that $\cup_{W\in\mathsf{S}}=\emptyset$ if $\mathsf{S}=\emptyset$.
-With this notation in place, and letting $U_{1}^{\prime}=G(B)\setminus U_{1}$, we have
-$\begin{split}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{G(B)}(u^{n})-\lambda_{2}(G(B))&=\Big{(}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})\Big{)}+\Big{(}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\Big{)}\end{split}$ (32)
-and we start by studying the first term on the r.h.s. of (32).
-To do so assume first that $\mathsf{U}_{1}\neq\emptyset$ and, for all $\xi_{1}\in\{0,\dots,L-1\}$, let
-$S_{\xi_{1}}=\Big{\{}k\in\{0,\dots,L-1\}:\,\,(\xi_{1}/L,(\xi_{1}+1)/L]\times(k/L,(k+1)/L]\in\mathsf{U}_{1}\Big{\}}$
-and let $I=\{\xi_{1}\in\{0,\dots,L-1\}:\,S_{\xi_{1}}\neq\emptyset\}$. Then, using the expression (31) for $G(B)$, it is direct to see that for all $\xi_{1}\in I$ the set
-$Q(\xi_{1}):=\bigcup_{k\in S_{\xi_{1}}}(\xi_{1}/L,(\xi_{1}+1)/L]\times(k/L,(k+1)/L]$
-belongs to $\mathcal{B}_{2}$ and that we have $\cup_{\xi_{1}\in I}Q(\xi_{1})=U_{1}$. Therefore,
-[…]
-where the second inequality uses the fact that, since $\{u^{n}\}_{n=1}^{N}$ is assumed to be a point set in $(0,1)^{2}$, we have (see Niederreiter,, 1992, Remarks 2-3 and Proposition 2.4, pages 14-15)
-$\sup_{(a,b]\subset(0,1)^{2}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{(a,b]}(u^{n})-\lambda_{2}\big{(}(a,b]\big{)}\Big{|}\leq 4D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}.$ (34)
-If $\mathsf{U}_{1}=\emptyset$ we have $\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}}(u^{n})-\lambda_{2}(U_{1})=0$ and thus (33) also holds in this case.
-We now study the second term on the r.h.s. of (32). To this end note that we can cover $U_{1}^{\prime}$ with sets in $\mathsf{U}_{2}$, and thus
-$-\lambda_{2}(U_{2})\leq-\lambda_{2}(U_{1}^{\prime})\leq\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\leq\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{2}}(u^{n})-\lambda_{2}(U_{2})+\lambda_{2}(U_{2}).$
-By using this latter result, we obtain that
-$\begin{split}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{1}^{\prime}}(u^{n})-\lambda_{2}(U_{1}^{\prime})\Big{|}&\leq\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{U_{2}}(u^{n})-\lambda_{2}(U_{2})\Big{|}+\lambda_{2}(U_{2})\leq|\mathsf{U}_{2}|\Big{(}4D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}+L^{-2}\Big{)}\end{split}$ (35)
-where the second inequality holds by (34).
-Using (31) and the fact that, by assumption, the function $\{x_{1},\dots,x_{M}\}\ni x\mapsto F_{\mu_{x}}(b_{2})$ is either non-increasing or non-decreasing, it is readily checked that $|\mathsf{U}_{2}|\leq 2L$.
-Together with (30) and (32)-(35), this implies that
-$\sup_{B\in\mathcal{B}_{2}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{B}\big{(}\varphi_{\mu}(u^{n})\big{)}-\mu(B)\Big{|}\leq 12LD^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}+2L^{-1}$
-and the result of the lemma follows by applying this latter inequality with
-$L=\Big{\lceil}\big{\{}6D^{*}_{2}\big{(}\{u^{n}\}_{n=1}^{N}\big{)}\big{\}}^{-1/2}\Big{\rceil}.$
-A.2.7 A stability result for the particle system generated by Algorithm 1 based on low discrepancy point sets
-Lemma 9.
-Let $N\geq 1$ and, for all $t\geq 1$, let $\{u_{t}^{n}=(u_{t,1}^{n},u_{t,2}^{n})\}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ such that, for $i=1,2$, we have $u_{t,i}\neq a2^{-k}$ for all $k\in\{1,\dots,7\}$ and all $a\in\{0,\dots,2^{7}\}$, and such that $D^{*}_{2}(\{u_{t}^{n}\}_{n=1}^{N})<2^{-10}$.
-Consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}=\{u_{t}^{n}\}_{n=1}^{N}$ for all $t\geq 1$ and let $\xi\in(0,\infty)$ be such that
-$\begin{split}\xi>\max\bigg{\{}&|F_{\eta_{1}}^{-1}(1/4)|,|F_{\eta_{1}}^{-1}(3/4)|,\frac{c\sigma^{2}\Phi^{-1}(3/4)^{2}+2\log(256/3)}{2c\sigma\Phi^{-1}(3/4)},\sigma\Phi^{-1}(3/4)+\sigma\Phi^{-1}(7/8)\bigg{\}}.\end{split}$
-Then, $\inf_{t\geq 1}\eta_{t}^{N}\big{(}[-\xi,\xi]\big{)}\geq 2^{-8}$ and $\inf_{t\geq 1}\hat{\eta}_{t}^{N}\big{(}[-\xi,\xi]\big{)}\geq 1/4$.
-Remark 5.
-It can be shown that if, in the statement of the lemma, $\{U_{t}^{n}\}_{n=1}^{N}$ is a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$ then the conclusion of the lemma holds $\mathbb{P}$-a.s.
-for $N=64$ and all $N\geq 128$.
-Proof.
-To prove the lemma let $t\geq 1$ and remark first that since $\{u_{t}^{n}\}_{n=1}^{N}$ is a point set in $(0,1)^{2}$ we have (see Niederreiter,, 1992, Remarks 2-3 and Proposition 2.4, pages 14-15)
-$\begin{split}\sup_{[a,b]\subset[0,1)^{2}}\Big{|}\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}_{[a,b]}(u_{t}^{n})-\lambda_{2}\big{(}[a,b]\big{)}\Big{|}\leq 4D^{*}_{2}(\{u_{t}^{n}\}_{n=1}^{N}).\end{split}$ (36)
-Let $[a,b]\subset(0,1)^{2}$ be such that $\lambda_{2}([a,b])\geq 2^{-7}$. Then, using (36), it follows that under the assumptions on $\{u_{t}^{n}\}_{n=1}^{N}$ we have
-$\sum_{n=1}^{N}\mathbbm{1}_{[a,b]}(u_{t}^{n})\geq N\Big{(}\lambda_{2}([a,b])-4D^{*}_{2}(\{u_{t}^{n}\}_{n=1}^{N})\Big{)}\geq N\big{(}2^{-7}-2^{-8}\big{)}=N2^{-8}$
-showing that, for all $t\geq 1$, every closed interval $[a,b]\subset(0,1)^{2}$ of volume at least equal to $2^{-7}$ contains at least $N2^{-8}$ elements of $P_{t,N}:=\{u_{t}^{n}\}_{n=1}^{N}$.
-We now let
-[…]
-so that proving the lemma amounts to showing that
-$\frac{1}{N}\sum_{n=1}^{N}\delta_{[-\xi,\xi]}(X_{t}^{n})\geq 2^{-8},\quad\sum_{n\in I_{t}\cup J_{t}}W_{t}^{n}\geq\frac{1}{4},\quad\forall t\geq 1.$ (37)
-To show that (37) holds for $t=1$ remark that $X_{1}^{n}=F_{\eta_{1}}^{-1}(u_{1,1}^{n})$ for all $n\in\{1,\dots,N\}$ and remark that the set $Q:=\cup_{i=0}^{2^{6}-1}\big{[}(2^{5}+i)2^{-7},(2^{5}+i+1)2^{-7}\big{]}$ is such that $\sum_{i=1}^{N}\mathbbm{1}_{Q}(u_{1,1}^{n})=\sum_{i=1}^{N}\mathbbm{1}_{[1/4,3/4]}(u_{1,1}^{n})$ under the assumptions on $\{u_{1}^{n}\}_{n=1}^{N}$.
-Therefore, it follows from the above computations that the interval $[1/4,3/4]$ contains at least $N/4$ elements of $\{u_{1,1}^{n}\}_{n=1}^{N}$, and thus the set $\{X_{1}^{n}\}_{n=1}^{N}$ has at least $N/4$ elements in the interval $[F_{\eta_{1}}^{-1}(1/4),F_{\eta_{1}}^{-1}(3/4)]\subset[-\xi,\xi]$.
-Then, by using the definition of $G(\cdot)$, it follows that $\sum_{n\in I_{1}\cup J_{1}}W_{1}^{n}\geq 1/4$, and consequently (37) holds for $t=1$.
-Assume now that (37) holds for some $t\geq 1$.
-Then, we have $\sum_{n\in I_{t}\cup J_{t}}W_{t}^{n}\geq 1/4$ and thus we have $\sum_{n\in I_{t}}W_{t}^{n}\geq 1/8$ and/or $\sum_{n\in J_{t}}W_{t}^{n}\geq 1/8$.
-Assume first that $\sum_{n\in I_{t}}W_{t}^{n}\geq 1/8$.
-Then, using the definition of $G(\cdot)$, it is easily checked that there exists an $i\in\{0,\dots,15\}$ such that $0\leq F^{-1}_{\hat{\eta}_{t}^{N}}(u)\leq\xi$ for all $u\in[i/16,(1+i)/16]$.
-Therefore, since as proved above the set $[i/16,(1+i)/16]\times[1/8,2/8]$ contains at least $N2^{-8}$ elements of $P_{t+1,N}$, it follows that there exists a set $\hat{I}_{t}\subseteq\{1,\dots,N\}$ such that $|\hat{I}_{t}|\geq N2^{-8}$ and such that, for all $n\in\hat{I}_{n}$,
-$\hat{X}_{t}^{n}+\sigma\Phi^{-1}(1/8)\leq X_{t+1}^{n}=\hat{X}_{t}^{n}+\sigma\Phi^{-1}(u_{t+1,2}^{n})\leq\hat{X}_{t}^{n}+\sigma\Phi^{-1}(2/8),\quad 0\leq\hat{X}_{t}^{n}\leq\xi.$ (38)
-To proceed further remark that if $(\hat{X}_{t}^{n},X_{t+1}^{n})$ is as in (38) then
-$|X_{t+1}^{n}|$ $\leq\max\big{\{}|\xi+\sigma\Phi^{-1}(2/8)|,\sigma|\Phi^{-1}(1/8)|\big{\}}=\max\big{\{}|\xi-\sigma\Phi^{-1}(6/8)|,\sigma\Phi^{-1}(7/8)\big{\}}\leq\xi-\delta$
-where the inequality holds with $\delta=\sigma\Phi^{-1}(6/8)$ and uses the condition imposed on $\xi$ in the statement of the lemma.
-By combining the above calculations, it follows that if $\sum_{n\in I_{t}}W_{t}^{n}\geq 1/8$ then there exists a set $\hat{I}_{t}\subseteq\{1,\dots,N\}$ such that $|\hat{I}_{t}|\geq N2^{-8}$ and such that $|X_{t+1}^{n}|\leq\xi-\delta<\xi$ for all $n\in\hat{I}_{t}$.
-Using a similar argument as above, one can easily show that if $\sum_{n\in J_{t}}W_{t}^{n}\geq 1/8$ then there exists a set $\hat{J}_{t}\subseteq\{1,\dots,N\}$ such that $|\hat{J}_{t}|\geq N2^{-8}$ and such that, for all $n\in\hat{J}_{t}$,
-$|X_{t+1}^{n}|$ $\leq\max\big{\{}|-\xi+\sigma\Phi^{-1}(6/8)|,\sigma\Phi^{-1}(7/8)\big{\}}=\max\big{\{}|\xi-\sigma\Phi^{-1}(6/8)|,\sigma\Phi^{-1}(7/8)\big{\}}\leq\xi-\delta.$
-Therefore, under the inductive hypothesis, we have
-$\sum_{n=1}^{N}\delta_{[-\xi+\delta,\xi-\delta]}(X_{t+1}^{n})\geq N2^{-8}$ (39)
-and thus
-$\begin{split}\sum_{n\not\in(I_{t+1}\cup J_{t+1})}W_{t+1}^{n}&\leq\frac{Ne^{-\frac{c}{2}\xi^{2}}}{Ne^{-\frac{c}{2}\xi^{2}}+N2^{-8}e^{-\frac{c}{2}(\xi-\delta)^{2}}}=\frac{2^{8}}{2^{8}+e^{-\frac{c}{2}\delta^{2}+c\delta\xi}}<\frac{3}{4}\end{split}$ (40)
-where the last inequality holds under the condition on $\xi$ imposed in the statement of the lemma.
-Combining (39) and (40) show that (37) holds at time $t+1$ and the proof of (37), and thus of the lemma, is complete.
-∎
-A.2.8 A discrepancy bound for Algorithm 1 based on low discrepancy point sets
-Lemma 10.
-Let $(T,N)\in\mathbb{N}^{2}$ and $\{u_{t}^{n}\}_{n=1}^{N}$ be a point set in $(0,1)^{2}$ for all $t\in\{1,\dots,T\}$, and consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}=\{u_{t}^{n}\}_{n=1}^{N}$ for all $t\in\{1,\dots,T\}$.
-Assume that there exist a non-empty and bounded interval $I\subset\mathbb{R}$ and a constant $\iota\in(0,1]$ such that
-$\hat{\eta}_{t}^{N}(I)\geq\iota,\quad\eta_{t}^{N}(I)\geq\iota,\quad\forall t\in\{1,\dots,T\}.$
-Then, there exists a constant $C\in(0,\infty)$, depending only on $c$, $\sigma$, $\rho$, $\iota$, $I$ and $\eta_{1}$, such that, with $\epsilon_{\star}\in(0,1)$ as in Lemma 5,
-$\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\inf_{\gamma\in(0,1)}\Big{(}\sqrt{2400}\delta^{1/2}+60\delta\Big{)}^{1-\gamma}\frac{C^{\gamma}}{1-\epsilon_{\star}^{\gamma}},\quad\delta=\sup_{t\in\{1,\dots,T\}}D_{2}^{*}(\{u_{t}^{n}\}_{n=1}^{N}).$
-In addition, $\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\sqrt{96}\delta^{1/2}+12\delta+\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|$.
-Proof.
-In what follows we let $(c_{t})_{t\geq 0}$, $(\rho_{t})_{t\geq 0}$ and $(\sigma_{t})_{t\geq 0}$ be as defined in Lemma 5, and for all $a\in\mathbb{R}$ we let $f_{a}=\mathbbm{1}_{(-\infty,a]}$ and
-$g_{a,k}(x)=\Phi\Big{(}\frac{a-\rho_{k}x}{\sigma_{k}}\Big{)},\quad h_{a,k}(x)=\Phi\Big{(}\frac{a-\rho_{k}x}{\sigma_{k}}\Big{)}-\Phi\Big{(}\frac{a}{\sigma_{k}}\Big{)},\quad\forall x\in\mathbb{R},\quad\forall k\geq 1.$
-Next, we let $\phi:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be such that $\phi(\eta)=\Psi_{G}(\eta M)$ for all $\eta\in\mathcal{P}(\mathbb{R})$, and for all $t\geq 1$ we let $\phi^{t}=\phi^{t-1}\phi$ with the convention that $\phi^{t-1}(\eta)=\eta$ for all $\eta\in\mathcal{P}(\mathbb{R})$ when $t=1$.
-Using this notation and following the approach introduced by Del Moral, (2004, Chapter 7) for studying PFs we have, using the convention that $\hat{\eta}^{N}_{0}M=\eta_{1}$ (so that $\phi(\hat{\eta}^{N}_{0})=\Psi_{G}(\eta_{1})$),
-$\hat{\eta}_{t}^{N}-\hat{\eta}_{t}$ $=\sum_{p=1}^{t}\Big{(}\phi^{t-p}(\hat{\eta}^{N}_{p})-\phi^{t-p}\big{(}\phi(\hat{\eta}^{N}_{p-1})\big{)}\Big{)},\quad\forall t\geq 1$ (41)
-and we now study the terms inside the sums.
-To this aim assume for now that $T\geq 3$, and let $t\in\{2,\dots,T\}$ and $p\in\{1,\dots,t-1\}$ be fixed. Then, for all $a\in\mathbb{R}$ we have, by Lemma 5,
-$\begin{split}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})&=\frac{\hat{\eta}^{N}_{p}Q^{t-p}(f_{a})}{\hat{\eta}^{N}_{p}Q^{t-p}(\mathbb{R})}=\frac{\sum_{n=1}^{N}W_{p}^{n}G_{c_{t-p}}(X_{p}^{n})g_{a,t-p}(X_{p}^{n})}{\sum_{n=1}^{N}W_{p}^{n}G_{c_{t-p}}(X_{p}^{n})}=\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(g_{a,t-p})\end{split}$ (42)
-and, similarly,
-$\phi^{t-p}\big{(}\phi(\hat{\eta}^{N}_{p-1}))(f_{a})=\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(g_{a,t-p}).$ (43)
-Then, by combining (42) and (43), we obtain that
-$\begin{split}\sup_{a\in\mathbb{R}}\big{|}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})-&\phi^{t-p}\big{(}\phi_{p}(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\big{|}\\ &=\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(g_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(g_{a,t-p})\Big{|}\\ &=\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(h_{a,t-p})\Big{|}.\end{split}$ (44)
-To proceed further remark that, by Lemma 4, we have $\sup_{(a,k)\in\mathbb{R}\times\mathbb{N}}V(h_{a,k})\leq 1$ and thus, by Lemma 7,
-$\begin{split}\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}&\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(h_{a,t-p})\Big{|}\leq\Big{\|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}\Big{\|}.\end{split}$ (45)
-Noting that
-$\|G_{a}\|_{\infty}=1,\quad V(G_{a})=\int_{\mathbb{R}}|G^{\prime}_{a}(x)|\mathrm{d}x=a\int_{\mathbb{R}}|x|G_{a}(x)\mathrm{d}x=2,\quad\forall a\in(0,\infty)$ (46)
-it follows from Corollary 1 that
-$\big{\|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}\Big{\|}$ $\leq\frac{5\|\hat{\eta}^{N}_{p}-\hat{\eta}^{N}_{p-1}M\|}{\hat{\eta}^{N}_{p-1}M(G_{c+c_{t-p}})}\leq\frac{5\|\hat{\eta}^{N}_{p}-\hat{\eta}^{N}_{p-1}M\|}{\hat{\eta}^{N}_{p-1}M(G_{c+c_{\star}})}$ (47)
-where the second inequality uses the fact that, by Lemma 5, we have $\sup_{k\geq 1}c_{k}\leq c_{\star}:=\rho^{2}/\sigma^{2}$.
-Then, using (47) and noting that under the assumptions of the lemma there exists a constant $\iota^{\prime}\in(0,1]$ such that
-$\inf_{k\in\mathbb{N}_{0}}\inf_{x\in I}\hat{\eta}_{k}^{N}M(G_{c+c_{\star}})\geq\iota^{\prime},$ (48)
-it follows that
-$\begin{split}\big{\|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}\Big{\|}\leq\frac{5\|\eta^{N}_{p-1}-\hat{\eta}^{N}_{p-1}M\|}{\iota^{\prime}}.\end{split}$ (49)
-By Lemma 8 and with $\delta$ as in the statement of the lemma, we have
-$\|\eta^{N}_{p}-\hat{\eta}^{N}_{p-1}M\|\leq\sqrt{96}\delta^{1/2}+12\delta$ (50)
-and thus, by combining (45), (49) and (50), we obtain that
-$\begin{split}\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}&(h_{a,t-p})\Big{|}\leq\frac{\sqrt{2400}\delta^{1/2}+60\delta}{\iota^{\prime}}.\end{split}$ (51)
-On the other hand, by Lemma 4 we have
-$|h_{a,k}(x)|\leq\frac{|\rho_{k}x|}{\sigma_{k}\sqrt{2\pi}},\quad\forall(a,x)\in\mathbb{R}^{2},\quad\forall k\geq 1$
-and therefore, noting that $|x|G_{a}(x)\leq a^{-1/2}e^{-1/2}$ for all $x\in\mathbb{R}$ and all $a\in(0,\infty)$, it follows that
-$G_{c+c_{t-p}}(x)|h_{a,t-p}(x)|\leq\frac{|\rho_{t-p}|e^{-1/2}}{\sigma_{t-p}\sqrt{(c+c_{t-p})2\pi}}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}},\quad\forall a\in\mathbb{R}$
-where the second inequality uses the fact that $\inf_{k\geq 1}\sigma^{2}_{k}=\sigma^{2}_{1}=\sigma^{2}/(1+c\sigma^{2})>0$ while $\inf_{k\geq 1}c_{k}\geq 0$.
-Using this result with and recalling that $c_{\star}=\rho^{2}/\sigma^{2}\geq\sup_{k\geq 1}c_{k}$, it follows that we have both
-$\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})\Big{|}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\eta_{p}^{N}(G_{c+c_{t-p}})}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\eta_{p}^{N}(G_{c+c_{\star}})}$
-and
-$\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(h_{a,t-p})\Big{|}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\hat{\eta}^{N}_{p-1}M(G_{c+c_{t-p}})}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\hat{\eta}^{N}_{p-1}M(G_{c+c_{\star}})}$
-and thus
-$\begin{split}\sup_{a\in\mathbb{R}}\Big{|}\Psi_{G_{c+c_{t-p}}}(\eta^{N}_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big{(}\hat{\eta}^{N}_{p-1}M\big{)}(h_{a,t-p})\Big{|}\leq\frac{\bar{C}}{\iota^{\prime}}|\rho_{t-p}|\end{split}$ (52)
-with $\bar{C}=\max\{1,2/(c\sigma_{1}^{2})^{1/2}\}$ and assuming without generality that the constant $\iota^{\prime}>0$ such that (48) holds is sufficiently small to ensure that $\big{(}\iota\inf_{x\in I}G_{c+c_{\star}}(x)\big{)}\geq\iota^{\prime}$, with the constant $\iota$ and the interval $I$ as in the statement of the lemma.
-Let $\gamma\in(0,1)$ be fixed. Then, by combining (44), (51) and (52) it follows that
-$\begin{split}\sup_{a\in\mathbb{R}}\big{|}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})-\phi^{t-p}\big{(}\phi_{p}(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\big{|}&\leq\frac{|\rho_{t-p}|^{\gamma}\bar{C}^{\gamma}\big{(}\sqrt{2400}\delta^{1/2}+60\delta\big{)}^{1-\gamma}}{\iota^{\prime}}.\end{split}$ (53)
-The inequality in (53) holds for all $t\in\{2,\dots,T\}$ and all $p\in\{1,\dots,t-1\}$ (assuming that $T\geq 3$), and we now consider the remaining cases where $p=t$ for some $t\in\{1,\dots,T\}$.
-To do so let $t\in\{1,\dots,T\}$ and $p=t$ and note that
-$\sup_{a\in\mathbb{R}}\big{|}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})-\phi^{t-p}\big{(}\phi(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\big{|}$ $=\sup_{a\in\mathbb{R}}\big{|}\hat{\eta}^{N}_{t}(f_{a})-\phi(\hat{\eta}^{N}_{t-1})(f_{a})\big{|}$
-$=\sup_{a\in\mathbb{R}}\big{|}\Psi_{G}(\eta^{N}_{t})(f_{a})-\Psi_{G}(\hat{\eta}^{N}_{t-1}M)(f_{a})\big{|}$
-$\leq\frac{5\|\eta^{N}_{t}-\hat{\eta}^{N}_{t-1}M\|}{\hat{\eta}^{N}_{t-1}M(G_{c})}$
-$\leq\frac{5\|\eta^{N}_{t}-\hat{\eta}^{N}_{t-1}M\|}{\hat{\eta}^{N}_{t-1}M(G_{c+c_{\star}})}$
-where the first inequality holds by Corollary 1 and uses (46), and where the last inequality uses the fact that $c_{\star}=\rho^{2}/\sigma^{2}\geq 0$.
-Then, using (48) and the fact that, by Lemma 8, we have $\|\eta^{N}_{t}-\hat{\eta}^{N}_{t-1}M\|\leq\sqrt{96}\delta^{1/2}+12\delta$ for all $t\in\{1,\dots,T\}$, it follows that for all $t\in\{1,\dots,T\}$ we have, for $p=t$ and recalling that $\rho_{0}=1$,
-$\sup_{a\in\mathbb{R}}\big{|}\phi^{p-t}(\hat{\eta}^{N}_{p})(f_{a})-\phi^{p-t}\big{(}\phi(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\big{|}$ $\leq(\iota^{\prime})^{-1}\min\big{\{}1,\sqrt{2400}\delta^{1/2}+60\delta\big{\}}$
-$\leq(\iota^{\prime})^{-1}\min\big{\{}1,\big{(}\sqrt{2400}\delta^{1/2}+60\delta\big{)}^{1-\gamma}\big{\}}$
-$\leq(\iota^{\prime})^{-1}|\rho_{t-p}|^{\gamma}\bar{C}^{\gamma}\big{(}\sqrt{2400}\delta^{1/2}+60\delta\big{)}^{1-\gamma}$
-showing that (53) holds for any integers $1\leq p\leq t\leq T$.
-We now let $C_{\star}\in(0,\infty)$ and $\epsilon_{\star}\in(0,1)$ be as in Lemma 5, so that $|\rho_{k}|\leq C_{\star}\epsilon_{\star}^{k}$ for all $k\geq 0$.
-Then, by using (41) and (53), it follows that
-$\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|=\sup_{t\in\{1,\dots,T\}}\sup_{a\in\mathbb{R}}\big{|}\hat{\eta}_{t}^{N}(f_{a})-\hat{\eta}_{t}(f_{a})\big{|}$ $\leq\sum_{p=1}^{T}\sup_{a\in\mathbb{R}}\Big{|}\phi^{t-p}(\hat{\eta}^{N}_{p})(f_{a})-\phi^{t-p}\big{(}\phi(\hat{\eta}^{N}_{p-1})\big{)}(f_{a})\Big{|}$
-$\leq(\iota^{\prime})^{-1}\big{(}\sqrt{2400}\delta^{1/2}+60\delta\big{)}^{1-\gamma}\frac{(C_{\star}\bar{C})^{\gamma}}{1-\epsilon_{\star}^{\gamma}}$
-and since $\gamma\in(0,1)$ is arbitrary the first part of the lemma follows.
-To show the second part of the lemma let $t\in\{2,\dots,T\}$ (assuming that $T\geq 2$) and note that
-$\|\eta_{t}^{N}-\eta_{t}\|$ $\leq\|\eta_{t}^{N}-\hat{\eta}^{N}_{t-1}M\|+\|\hat{\eta}^{N}_{t-1}M-\hat{\eta}_{t-1}M\|\leq\sqrt{96}\delta^{1/2}+12\delta+\|\hat{\eta}^{N}_{t-1}M-\hat{\eta}_{t-1}M\|$ (54)
-where the second inequality holds by Lemma 8. Noting that for all $a\in\mathbb{R}$ we have
-$M(f_{a})(x)=\Phi\bigg{(}\frac{a-\rho x}{\sigma}\bigg{)}=:\tilde{g}_{a}(x),\quad\forall x\in\mathbb{R}$
-where, using Lemma 4, $V(\tilde{g}_{a})=V(h_{a,\rho,\sigma})\leq 1$, it follows from Lemma 7 that (assuming that $T\geq 2$)
-$\|\hat{\eta}^{N}_{t-1}M-\hat{\eta}_{t-1}M\|=\sup_{a\in\mathbb{R}}|\hat{\eta}^{N}_{t-1}(\tilde{g}_{a})-\hat{\eta}_{t-1}(\tilde{g}_{a})|\leq\|\hat{\eta}^{N}_{t-1}-\hat{\eta}_{t-1}\|,\quad\forall t\in\{2,\dots,T\}.$
-By combining this latter result with (54) we obtain that
-$\sup_{t\in\{2,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\sqrt{96}\delta^{1/2}+12\delta+\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}^{N}_{t}-\hat{\eta}_{t}\|.$
-Then, the second part of the lemma follows upon noting that $\|\eta_{1}^{N}-\eta_{1}\|\leq\sqrt{96}\delta^{1/2}+12\delta$ by Lemma 8. The proof of the lemma is complete. ∎
-A.3 Proof of Lemmas 1-3, Theorems 1-2 and Propositions 1-3
-Lemma 1 is a direct consequence of Lemmas 9-10.
-Lemma 2 is a direct consequence of Lemmas 1, 6 and 7.
-Theorem 2 directly follows from (9) and Lemma 1.
-Proposition 2 is a direct consequence of Lemma 1 and Theorem 1.
-Consequently, below we only prove Lemma 3, Theorem 1 and Propositions 1- 3.
-A.3.1 Proof of Proposition 1
+5.3 Proof of Proposition 1
 Proof.
 The result of the proposition trivially holds if $\rho=0$ and therefore below we assume that $\rho\neq 0$.
-Let $(c_{t})_{t\geq 0}$, $(\rho_{t})_{t\geq 0}$, $(\sigma_{t})_{t\geq 0}$, $C_{\star}\in(0,\infty)$ and $\epsilon_{\star}\in(0,1)$ be as in Lemma 5, and for all $a\in\mathbb{R}$ let $f_{a}=\mathbbm{1}_{(-\infty,a]}$ and, for all $r\in\mathbb{R}$ and $s\in(0,\infty)$, let $h_{a,r,s}$ be as defined in (11) and let $p_{a,s}=\Phi(a/s)$.
+Let $(c_{t})_{t\geq 0}$, $(\rho_{t})_{t\geq 0}$, $(\sigma_{t})_{t\geq 0}$, $C_{\star}\in(0,\infty)$ and $\epsilon_{\star}\in(0,1)$ be as in Lemma 5, and for all $a\in\mathbb{R}$ let $f_{a}=\mathbbm{1}_{(-\infty,a]}$ and, for all $r\in\mathbb{R}$ and $s\in(0,\infty)$, let $h_{a,r,s}$ be as defined in (28) and $p_{a,s}=\Phi(a/s)$.
 We now let $a\in\mathbb{R}$ be fixed. Then, using (6) and Lemma 5, we have
@@
 $|h_{a,\rho_{t},\sigma_{t}}(x)|$ $\leq\frac{|x\rho_{t}|}{\sigma_{1}\sqrt{2\pi}},\quad\forall x\in\mathbb{R},\quad\forall t\geq 1.$
-By combining this latter result and (55), and by using the fact that $\sup_{x\in\mathbb{R}}|x|G_{q}(x)\leq q^{-1/2}e^{-1/2}$ for all $q\in(0,\infty)$, it follows that
-$\big{|}\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{t}}\big{|}\leq\big{|}\Psi_{G_{c_{t}}}(\hat{\eta}_{1})(h_{a,\rho_{t},\sigma_{t}})\big{|}\leq\frac{|\rho_{t}|e^{-1/2}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{2\pi c_{t}}}\leq\frac{|\rho_{t}|\sqrt{1+c\sigma^{2}}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{c\rho^{2}}},\quad\forall t\geq 1$ (56)
+By combining this latter result and (36), and by using the fact that $\sup_{x\in\mathbb{R}}|x|G_{z}(x)\leq z^{-1/2}e^{-1/2}$ for all $z\in(0,\infty)$, it follows that, for all $t\geq 1$
+$\begin{split}\big|\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{t}}\big|\leq\big|\Psi_{G_{c_{t}}}(\hat{\eta}_{1})(h_{a,\rho_{t},\sigma_{t}})\big|\leq\frac{|\rho_{t}|e^{-1/2}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{2\pi c_{t}}}\leq\frac{|\rho_{t}|\sqrt{1+c\sigma^{2}}}{\hat{\eta}_{1}(G_{c_{t}})\sigma_{1}\sqrt{c\rho^{2}}}\end{split}$ (37)
 where the last inequality uses Lemma 5.
-By Lemma 5 we have $G_{c_{t}}(x)\geq G_{\rho^{2}/\sigma^{2}}(x)$ for all $x\in\mathbb{R}$ and $|\rho^{t}|\leq C_{\star}\epsilon_{\star}^{t}$ for all $t\geq 1$, and thus, by (56),
+By Lemma 5 we have $G_{c_{t}}(x)\geq G_{\rho^{2}/\sigma^{2}}(x)$ for all $x\in\mathbb{R}$ while $|\rho^{t}|\leq C_{\star}\epsilon_{\star}^{t}$ for all $t\geq 1$, and thus, by (37),
 $\big|\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{t}}\big|\leq C_{1}|\rho_{t}|\leq C_{1}\epsilon_{\star}^{t},\quad\forall t\geq 1$ (38)
@@
 $C_{1}=\frac{C_{\star}\sqrt{1+c\sigma^{2}}}{\hat{\eta}_{1}(G_{\rho^{2}/\sigma^{2}})\sigma_{1}\sqrt{c\rho^{2}}}.$
-On the other hand, using the mean value theorem, for all $t\geq 1$ there exists a $v_{t,a}\in([-1,1])$ such that
+On the other hand, using the mean value theorem, for all $t\geq 1$ there exists a constant $v_{t,a}\in[-1,1]$ such that
 $|p_{a,\sigma_{t}}-p_{a,\sigma_{\infty}}|$ $\leq\frac{|\sigma_{t}-\sigma_{\infty}|}{\sigma_{t}+v_{t,a}(\sigma_{t}-\sigma_{\infty})}\,\,\frac{|a|}{\sigma_{t}+v_{t,a}(\sigma_{t}-\sigma_{\infty})}\varphi\Big(\frac{a}{\sigma_{t}+v_{t,a}(\sigma_{t}-\sigma_{\infty})}\Big).$
@@
 $|\sigma_{t}-\sigma_{\infty}|^{2}\leq|\sigma^{2}_{t}-\sigma^{2}_{\infty}|\leq C^{2}_{\star}\epsilon_{\star}^{2t},\quad\forall t\geq 1.$ (40)
-Noting that in the above calculations $a\in\mathbb{R}$ is arbitrary, by combining (57), (58) and (59) we obtain that
-$\|\hat{\eta}_{t+1}-\mathcal{N}_{1}(0,\sigma^{2}_{\infty})\|=\sup_{a\in\mathbb{R}}\big{|}\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{\infty}}\big{|}\leq\big{(}C_{1}+C_{\star}/\sigma_{1}\big{)}\epsilon_{\star}^{t},\quad\forall t\geq 1.$
+Noting that in the above calculations the constant $a\in\mathbb{R}$ is arbitrary, by combining (38), (39) and (40) we obtain that
+$\|\hat{\eta}_{t+1}-\mathcal{N}(0,\sigma^{2}_{\infty})\|=\sup_{a\in\mathbb{R}}\big|\hat{\eta}_{t+1}(f_{a})-p_{a,\sigma_{\infty}}\big|\leq\big(C_{1}+C_{\star}/\sigma_{1}\big)\epsilon_{\star}^{t},\quad\forall t\geq 1.$
 To complete the proof of the proposition it remains to show that
-$\|\eta_{t+1}-\mathcal{N}_{1}(0,\rho\sigma^{2}_{\infty}+\sigma^{2})\|\leq\|\eta_{t}-\mathcal{N}_{1}(0,\sigma^{2}_{\infty})\|,\quad\forall t\geq 1.$ (60)
-To this aim we let $\hat{\eta}_{\infty}=\mathcal{N}_{1}(0,\sigma^{2}_{\infty})$ and note that, for all $t\geq 2$, we have
-$\|\eta_{t}-\mathcal{N}_{1}(0,\rho\sigma^{2}_{\infty}+\sigma^{2})\|$ $=\|\hat{\eta}_{t-1}M-\hat{\eta}_{\infty}M\|=\sup_{a\in\mathbb{R}}|\hat{\eta}_{t-1}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|.$
+$\|\eta_{t+1}-\mathcal{N}(0,\rho^{2}\sigma^{2}_{\infty}+\sigma^{2})\|\leq\|\hat{\eta}_{t}-\mathcal{N}(0,\sigma^{2}_{\infty})\|,\quad\forall t\geq 1.$ (41)
+To this aim we let $\hat{\eta}_{\infty}=\mathcal{N}(0,\sigma^{2}_{\infty})$ and note that, for all $t\geq 1$, we have
+$\|\eta_{t+1}-\mathcal{N}(0,\rho^{2}\sigma^{2}_{\infty}+\sigma^{2})\|$ $=\|\hat{\eta}_{t}M-\hat{\eta}_{\infty}M\|=\sup_{a\in\mathbb{R}}|\hat{\eta}_{t}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|.$ (42)
 Since $\sup_{a\in\mathbb{R}}V(h_{a,\rho,\sigma})\leq 1$ by Lemma 3, it follows from Lemma 4 that
-$\sup_{a\in\mathbb{R}}|\hat{\eta}_{t-1}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|\leq\|\hat{\eta}_{t-1}-\hat{\eta}_{\infty}\|$
-and (60) follows. The proof of the proposition is complete. ∎
-A.3.2 Proof of Lemma 3
+$\sup_{a\in\mathbb{R}}|\hat{\eta}_{t}(h_{a,\rho,\sigma})-\hat{\eta}_{\infty}(h_{a,\rho,\sigma})|\leq\|\hat{\eta}_{t}-\hat{\eta}_{\infty}\|$
+and (41) then follows from (42). The proof of the proposition is complete. ∎
+5.4 Proof of Lemma 1
+5.4.1 Two preliminary results
+From Lemma 4 we readily obtain the following discrepancy bound for 1-dimensional importance sampling:
+Corollary 3.
+Let $H\in\mathcal{C}^{1}(\mathbb{R})$ be such that $V(H)<\infty$ and such that $H(x)\geq 0$ for all $x\in\mathbb{R}$, and let $\mu,\pi\in\mathcal{P}(\mathbb{R})$. Then,
+$\|\Psi_{H}(\mu)-\Psi_{H}(\pi)\|\leq\frac{\|H\|_{\infty}+2V(H)}{\pi(H)}\|\mu-\pi\|.$
 Proof.
-Let $\delta\in(0,1)$ and remark that, since $\{U_{1}^{n}\}_{n=1}^{N}$ is a set if $N$ independent $\mathcal{U}(0,1)^{s}$ random numbers, it follows from Aistleitner and Hofer, (2014, Theorem 1) that for all $\gamma\in(0,1)$ we have
-$\mathbb{P}\Big{(}D^{*}_{s}\big{(}\{U_{1}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}\geq 1-\gamma,\quad\forall N\geq N_{s,\delta,\gamma}:=\frac{160s+33\log(1/\gamma)}{\delta^{2}}.$
-Therefore, since ($(\{U_{t}^{n}\}_{n=1}^{N})_{t\geq 1}$ is assumed to be a sequence of independent sets, it follows that for all $\gamma\in(0,1)$ we have
-$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}D^{*}_{s}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}$ $=\mathbb{P}\Big{(}D^{*}_{s}\big{(}\{U_{1}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}^{T}\geq(1-\gamma)^{T},\quad\forall N\geq N_{s,\delta,\gamma}.$
-We now let $q\in(0,1)$ so that, by applying the latter result with $\gamma=1-(1-q)^{1/T}$, we have
-$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}D^{*}_{s}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}\geq 1-q,\quad\forall N\geq N_{s,\delta,1-(1-q)^{1/T}}.$
-Using Bernoulli’s inequality (see e.g. Li and Yeh,, 2013), we have $1-(1-q)^{1/T}\geq T^{-1}q$ and thus
-$N_{s,\delta,1-(1-q)^{1/T}}\leq\frac{160s-33\log(T^{-1}q)}{\delta^{2}}=\frac{160s+33\log(T)+33\log(1/q)}{\delta^{2}}$
-and the proof of the lemma is complete. ∎
-A.3.3 Proof of Theorem 1
+The result follows from Lemma 4 and the fact that, for all interval $B=(-\infty,a)\subset\mathbb{R}$, we have
+$|\Psi_{H}(\mu)(B)-\Psi_{H}(\pi)(B)|\leq\frac{|(\mu-\pi)(H\mathbbm{1}_{B})|+|(\mu-\pi)(H)|}{\pi(H)}.$
+The next lemma is the key intermediate result used to prove Lemma 1:
+Lemma 6.
+Let $I\subseteq\mathbb{R}$ be a compact set.
+Then, there exists a constant $C_{I}\in(1,\infty)$ such that, for any sequence $(\mu_{t})_{t\geq 1}$ and letting $\hat{\mu}_{t}=\Psi_{G}(\mu_{t})$ for all $t\geq 1$, for any $T\geq 1$ we have
+$\ \max\Big\{\sup_{t\in\{1,\dots,T\}}\|\hat{\mu}_{t}-\hat{\eta}_{t}\|,\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\eta_{t}\|\Big\}\leq\frac{C_{I}}{\iota_{T-1}}\Delta_{T}\log\big(1+\Delta_{T}^{-1}\big)$
+where $\iota_{T-1}=\inf_{t\in\{1,\dots,T-1\}}\mu_{t}(I)$ and $\Delta_{T}=\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\hat{\mu}_{t-1}M)\|$, with the convention that $\iota_{0}=1$ and that $\hat{\mu}_{0}M=\eta_{1}$.
 Proof.
-We start with two preliminary results.
-To state the first one let $\pi_{M}(\mathrm{d}x)=\sum_{m=1}^{M}W_{m}\delta_{\{x_{m}\}}\in\mathcal{P}(\mathbb{R})$ for some $M\in\mathbb{N}$, for all $x\in\{x_{1},\dots,x_{M}\}$ let $\mu_{x}\in\mathcal{P}(\mathbb{R})$, let $\mu=\sum_{m=1}^{M}W_{m}\mu_{x_{m}}\in\mathcal{P}(\mathbb{R})$, $\{u^{n}\}_{n=1}^{N}$ be a point set in $(0,1)$ and $\mu^{N}=N^{-1}\sum_{n=1}^{N}\delta_{F^{-1}_{\mu}(u^{n})}$.
-Then, it is direct to see that $\|\mu^{N}-\mu\|=D^{*}_{1}(\{u^{n}\}_{n=1}^{N})$.
-To state the second preliminary result let ($(\{U_{t}^{n}\}_{n=1}^{N})_{t\geq 1}$ be as in the statement of the theorem, and let $\mu_{1}^{N}=\eta_{1}^{N}$ and
-$\mu_{t+1}^{N}=\frac{1}{N}\sum_{n=1}^{N}\delta_{F^{-1}_{\hat{\mu}_{t}^{N}M}(U_{t+1,1}^{n})},\quad\hat{\mu}^{N}_{t}=\Psi_{G}(\mu_{t}^{N}),\quad\forall t\geq 1.$
-Then, it is direct to see that we have $(\hat{\eta}_{t}^{N})_{t\geq 1}\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{dist}}}}}{{=}}(\hat{\mu}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}\stackrel{{\scriptstyle\mathclap{\mbox{\tiny{dist}}}}}{{=}}(\mu_{t}^{N})_{t\geq 1}$.
-Using these two results, it follows that to prove the theorem we can use the version of Lemma 1 we obtain if, in Lemma 8, we replace the right-hand size of (29) by $2D_{1}^{*}(\{u_{1}^{n}\}_{n=1}^{N})$, that is, to prove the lemma we can do as if the conclusion of Lemma 1 holds with $\delta_{N,T}=\sup_{t\in\{1,\dots,T\}}D_{1}^{*}(\{U_{t,1}^{n}\}_{n=1}^{N})^{2}$.
-Using this observation the result of the theorem then follows from Lemma 3.
+Let $(\mu_{t})_{t\geq 1}$ and $(\hat{\mu}_{t})_{t\geq 1}$ be as in the statement of the lemma, $(c_{t})_{t\geq 0}$, $(\rho_{t})_{t\geq 0}$ and $(\sigma_{t})_{t\geq 0}$ be as defined in Lemma 5, and for all $a\in\mathbb{R}$ let $f_{a}=\mathbbm{1}_{(-\infty,a]}$ and
+$g_{a,k}(x)=\Phi\Big(\frac{a-\rho_{k}x}{\sigma_{k}}\Big),\quad h_{a,k}(x)=\Phi\Big(\frac{a-\rho_{k}x}{\sigma_{k}}\Big)-\Phi\Big(\frac{a}{\sigma_{k}}\Big),\quad\forall x\in\mathbb{R},\quad\forall k\geq 1.$
+Next, we let $\phi:\mathcal{P}(\mathbb{R})\rightarrow\mathcal{P}(\mathbb{R})$ be such that $\phi(\eta)=\Psi_{G}(\eta M)$ for all $\eta\in\mathcal{P}(\mathbb{R})$, and for all $t\geq 1$ we let $\phi^{t}=\phi^{t-1}\phi$ with the convention that $\phi^{0}(\eta)=\eta$ for all $\eta\in\mathcal{P}(\mathbb{R})$.
+Note that $\hat{\eta}_{t+1}=\phi^{t}(\hat{\eta}_{1})$ for all $t\geq 1$.
+Using this notation, and following the approach introduced by Del Moral, (2004, Chapter 7) for studying PFs, we have, using the convention that $\phi(\hat{\mu}_{0})=\hat{\eta}_{1}$,
+$\hat{\mu}_{t}-\hat{\eta}_{t}$ $=\sum_{p=1}^{t}\Big(\phi^{t-p}(\hat{\mu}_{p})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)\Big),\quad\forall t\geq 1$ (43)
+and we now study the terms inside the sums.
+To this aim, we assume for now that $T\geq 3$ , and we let $t\in\{2,\dots,T\}$ and $p\in\{1,\dots,t-1\}$ be fixed. Then, for all $a\in\mathbb{R}$ we have, by Lemma 5,
+$\begin{split}\phi^{t-p}(\hat{\mu}_{p})(f_{a})&=\frac{\hat{\mu}_{p}Q^{t-p}(f_{a})}{\hat{\mu}_{p}Q^{t-p}(\mathbb{R})}=\Psi_{G_{c+c_{t-p}}}(\mu_{p})(g_{a,t-p})\end{split}$ (44)
+and
+$\phi^{t-p}\big(\phi(\hat{\mu}_{p-1}))(f_{a})=\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(g_{a,t-p}).$ (45)
+By combining (44) and (45), we obtain that
+$\begin{split}\sup_{a\in\mathbb{R}}\big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-&\phi^{t-p}\big(\phi_{p}(\hat{\mu}_{p-1})\big)(f_{a})\big|\\ &=\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(g_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(g_{a,t-p})\Big|\\ &=\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(h_{a,t-p})\Big|.\end{split}$ (46)
+To proceed further we note that, by Lemma 3, we have $\sup_{(a,k)\in\mathbb{R}\times\mathbb{N}}V(h_{a,k})\leq 1$ and thus, by Lemma 4,
+$\begin{split}\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}&\big(\hat{\mu}_{p-1}M\big)(h_{a,t-p})\Big|\leq\Big\|\Psi_{G_{c+c_{t-p}}}(\mu_{p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)\Big\|.\end{split}$ (47)
+Noting that
+$\|G_{a}\|_{\infty}=1,\quad V(G_{a})=\int_{\mathbb{R}}|G^{\prime}_{a}(x)|\mathrm{d}x=a\int_{\mathbb{R}}|x|G_{a}(x)\mathrm{d}x=2,\quad\forall a\in(0,\infty)$ (48)
+it follows from Corollary 3 that
+$\begin{split}\big\|\Psi_{G_{c+c_{t-p}}}(\mu_{p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)\Big\|\leq\frac{5\|\mu_{p}-\hat{\mu}_{p-1}M\|}{\mu_{p}(G_{c+c_{t-p}})}&\leq\frac{5\|\mu_{p}-\hat{\mu}_{p-1}M\|}{\mu_{p}(G_{c+c_{\star}})}\leq\frac{5\Delta_{T}}{\mu_{p}(G_{c+c_{\star}})}\end{split}$ (49)
+where the second inequality uses the fact that, by Lemma 5, we have $\sup_{k\geq 1}c_{k}\leq c_{\star}:=\rho^{2}/\sigma^{2}$.
+Noting that since $p<T$ we have $\mu_{p}(G_{c+c_{\star}})\geq\iota^{\prime}\iota_{T-1}$, with $\iota^{\prime}=\inf_{x\in I}G_{c+c_{\star}}(x)>0$, it follows from (49) that
+$\begin{split}\big\|\Psi_{G_{c+c_{t-p}}}(\mu_{p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)\Big\|\leq\frac{5\Delta_{T}}{\iota^{\prime}\iota_{T-1}}\end{split}$
+which, together with (47), shows that
+$\begin{split}\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)&(h_{a,t-p})\Big|\leq\frac{5\Delta_{T}}{\iota^{\prime}\iota_{T-1}}.\end{split}$ (50)
+On the other hand, by Lemma 3 we have
+$|h_{a,k}(x)|\leq\frac{|\rho_{k}x|}{\sigma_{k}\sqrt{2\pi}},\quad\forall(a,x)\in\mathbb{R}^{2},\quad\forall k\geq 1$
+and therefore, noting that $|x|G_{a}(x)\leq a^{-1/2}e^{-1/2}$ for all $x\in\mathbb{R}$ and all $a\in(0,\infty)$,
+$G_{c+c_{t-p}}(x)|h_{a,t-p}(x)|\leq\frac{|\rho_{t-p}|e^{-1/2}}{\sigma_{t-p}\sqrt{(c+c_{t-p})2\pi}}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}},\quad\forall a\in\mathbb{R}$
+where the second inequality uses the fact that $\inf_{k\geq 1}\sigma^{2}_{k}=\sigma^{2}_{1}=\sigma^{2}/(1+c\sigma^{2})>0$ while $\inf_{k\geq 1}c_{k}\geq 0$ by Lemma 5.
+Using the latter result and recalling that $c_{\star}=\rho^{2}/\sigma^{2}\geq\sup_{k\geq 1}c_{k}$, it follows that we have both
+$\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})\Big|\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\mu_{p}(G_{c+c_{t-p}})}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\mu_{p}(G_{c+c_{\star}})}$
+and
+$\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(h_{a,t-p})\Big|\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\hat{\mu}_{p-1}M(G_{c+c_{t-p}})}\leq\frac{|\rho_{t-p}|}{\sigma_{1}\,\sqrt{c}\,\hat{\mu}_{p-1}M(G_{c+c_{\star}})}$
+where $\hat{\mu}_{p-1}M(G_{c+c_{\star}})\geq\iota^{\prime\prime}\iota_{T-1}$ with $\iota^{\prime\prime}=\eta_{1}(G_{c+c_{\star}})\inf_{x\in I}G(x)M(x,G_{c+c_{\star}})$.
+(This definition of $\iota^{\prime\prime}$ allows to cover both the case where $p=1$ and the case where $p>1$.) Consequently, if we define the constant $\bar{C}=\max\{5,2/(c\sigma_{1}^{2})^{1/2}\}$, then
+$\begin{split}\sup_{a\in\mathbb{R}}\Big|\Psi_{G_{c+c_{t-p}}}(\mu_{p})(h_{a,t-p})-\Psi_{G_{c+c_{t-p}}}\big(\hat{\mu}_{p-1}M\big)(h_{a,t-p})\Big|\leq\frac{\bar{C}}{\iota^{\prime\prime}\iota_{T-1}}|\rho_{t-p}|.\end{split}$ (51)
+By combining (46), (50) and (51), and recalling that $\bar{C}\geq 5$, we obtain that
+$\begin{split}\sup_{a\in\mathbb{R}}\big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi_{p}(\hat{\mu}_{p-1})\big)(f_{a})\big|&\leq\frac{\bar{C}}{\iota^{\prime\prime\prime}\iota_{T-1}}\,|\rho_{t-p}|^{\gamma}\Delta_{T}^{1-\gamma},\quad\forall\gamma\in[0,1]\end{split}$ (52)
+with $\iota^{\prime\prime}=\min\{\iota^{\prime},\iota^{\prime\prime}\}>0$.
+The result in (52) holds for all $t\in\{2,\dots,T\}$ and all $p\in\{1,\dots,t-1\}$ (assuming that $T\geq 3$), and we now show that it also holds when $p=t$ for some $t\in\{1,\dots,T\}$, and with $T\geq 1$ arbitrary.
+To do so we let $\gamma\in[0,1]$, $t\in\{1,\dots,T\}$ and $p=t$, and we note that
+$\sup_{a\in\mathbb{R}}\big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\big|$ $=\sup_{a\in\mathbb{R}}\big|\hat{\mu}_{t}(f_{a})-\phi(\hat{\mu}_{t-1})(f_{a})\big|$
+$=\sup_{a\in\mathbb{R}}\big|\Psi_{G}(\mu_{t})(f_{a})-\Psi_{G}(\hat{\mu}_{t-1}M)(f_{a})\big|$
+$\leq\frac{5\|\mu_{t}-\hat{\mu}_{t-1}M\|}{\hat{\mu}_{t-1}M(G_{c})}$
+$\leq\frac{5\Delta_{T}}{\hat{\mu}_{t-1}M(G_{c+c_{\star}})}$
+where the first inequality holds by Corollary 3 and uses (48), and where the last inequality uses the fact that $c_{\star}=\rho^{2}/\sigma^{2}\geq 0$.
+Recalling that $\hat{\mu}_{k}M(G_{c+c_{\star}})\geq\iota^{\prime\prime}\iota_{T-1}\geq\iota^{\prime\prime\prime}\iota_{T-1}$ for all $k\in\{0,\dots,T-1\}$, and noting that since we are assuming that $p=t$ we have $\rho_{t-p}=\rho_{0}=1$, it follows that
+$\sup_{a\in\mathbb{R}}\big|\phi^{p-t}(\hat{\mu}_{p})(f_{a})-\phi^{p-t}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\big|$ $\leq\bar{C}(\iota^{\prime\prime\prime}\iota_{T-1})^{-1}\min\big\{1,\Delta_{T}\big\}$
+$\leq\bar{C}(\iota^{\prime\prime\prime}\iota_{T-1})^{-1}\min\big\{1,\Delta_{T}^{1-\gamma}\big\}$
+$\leq\bar{C}(\iota^{\prime\prime\prime}\iota_{T-1})^{-1}|\rho_{t-p}|^{\gamma}\Delta_{T}^{1-\gamma}$
+showing that (52) holds for any integers $1\leq p\leq t\leq T$.
+We now let $C_{\star}\in(0,\infty)$ and $\epsilon_{\star}\in(0,1)$ be as in Lemma 5, so that $|\rho_{k}|\leq C_{\star}\epsilon_{\star}^{k}$ for all $k\geq 0$.
+In addition, we let $t\in\{1,\dots,T\}$ and $k\in\mathbb{N}$.
+Then, by using (43) and (52), and using the convention that empty sums equal zero, we have
+$\begin{split}\|\hat{\mu}_{t}-\hat{\eta}_{t}\|&=\sup_{a\in\mathbb{R}}\big|\hat{\mu}_{t}(f_{a})-\hat{\eta}_{t}(f_{a})\big|\\ &\leq\sum_{p=1}^{t}\sup_{a\in\mathbb{R}}\Big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\Big|\\ &=\sum_{p=1}^{t-k}\sup_{a\in\mathbb{R}}\Big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\Big|\\ &+\sum_{p=t-k+1}^{t}\sup_{a\in\mathbb{R}}\Big|\phi^{t-p}(\hat{\mu}_{p})(f_{a})-\phi^{t-p}\big(\phi(\hat{\mu}_{p-1})\big)(f_{a})\Big|\\ &\leq\frac{\bar{C}}{\iota^{\prime\prime\prime}\iota_{T-1}}\bigg(C_{\star}\sum_{p=1}^{t-k}\epsilon_{\star}^{t-p}+k\Delta_{T}\bigg)\\ &\leq\frac{\bar{C}}{\iota^{\prime\prime\prime}\iota_{T-1}}\bigg(\epsilon_{\star}^{k}C_{\star}(1-\epsilon_{\star})^{-1}+k\Delta_{T}\bigg).\end{split}$ (53)
+By applying (53) with
+$k=\bigg\lceil\frac{1}{\log(1/\epsilon_{\star})}\log\bigg(\frac{C^{\prime\prime}\log(1/\epsilon_{\star})}{\Delta_{T}}\bigg)\bigg\rceil,\quad C^{\prime\prime}:=C_{\star}(1-\epsilon_{\star})^{-1},$
+we obtain that
+$\|\hat{\mu}_{t}-\hat{\eta}_{t}\|$ $\leq\Delta_{T}\,\bigg(\frac{\bar{C}}{\iota^{\prime\prime\prime}\iota_{T-1}\log(1/\epsilon_{\star})}\bigg)\bigg(2+\log(C^{\prime\prime})+\log\big(\log(1/\epsilon_{\star})\big)-\log(\epsilon_{\star}\Delta_{T})\bigg)$
+$\leq C^{\prime\prime\prime}\Delta_{T}\log(1+\Delta_{T}^{-1})$
+where the second inequality holds for some constant $C^{\prime\prime\prime}\in(0,\infty)$ depending only on $C^{\prime\prime}$ and $\epsilon_{\star}$, and uses the fact that $\Delta_{T}\in[0,1]$.
+The first part of the lemma follows.
+To show the second part of the lemma assume first that $T\geq 2$, let $t\in\{2,\dots,T\}$, and note that
+$\|\mu_{t}-\eta_{t}\|$ $\leq\|\mu_{t}-\hat{\mu}_{t-1}M\|+\|\hat{\mu}_{t-1}M-\hat{\eta}_{t-1}M\|\leq\Delta_{T}+\|\hat{\mu}_{t-1}M-\hat{\eta}_{t-1}M\|.$ (54)
+Noting that for all $a\in\mathbb{R}$ we have
+$M(f_{a})(x)=\Phi\bigg(\frac{a-\rho x}{\sigma}\bigg)=:\tilde{g}_{a}(x),\quad\forall x\in\mathbb{R}$
+where, using Lemma 3, $V(\tilde{g}_{a})=V(h_{a,\rho,\sigma})\leq 1$, it follows from Lemma 4 that
+$\|\hat{\mu}_{t-1}M-\hat{\eta}_{t-1}M\|=\sup_{a\in\mathbb{R}}|\hat{\mu}_{t-1}(\tilde{g}_{a})-\hat{\eta}_{t-1}(\tilde{g}_{a})|\leq\|\hat{\mu}_{t-1}-\hat{\eta}_{t-1}\|.$
+By combining this latter result with (54) we obtain that
+$\sup_{t\in\{2,\dots,T\}}\|\mu_{t}-\eta_{t}\|\leq\Delta_{T}+\sup_{t\in\{1,\dots,T\}}\|\hat{\mu}_{t}-\hat{\eta}_{t}\|.$
+Then, the second part of the lemma follows from the first part of the lemma, upon noting that $\|\hat{\mu}_{1}-\eta_{1}\|=\|\hat{\mu}_{1}-\hat{\mu}_{0}^{N}M\|\leq\Delta_{T}$.
+This ends the proof of the lemma.
 ∎
-A.3.4 Proof of Proposition 3
-Proof.
-Let $\kappa\in(0,1)$ and let $T_{\kappa}\in\mathbb{N}$ and $a_{\kappa}\in(0,\infty)$ be such that we have both $\eta_{t}([-a_{\kappa},a_{\kappa}])\leq\kappa$ and $\hat{\eta}_{t}([-a_{\kappa},a_{\kappa}])\leq\kappa$ for all $T\geq T_{\kappa}$.
-Remark that such constants $T_{\kappa}$ and $a_{\kappa}$ exist by Proposition 1.
-Next, let $\varrho_{\kappa}=\min\big{\{}1-\Phi(4a_{\kappa}/\sigma),1/2\big{\}}$ so that, using (5) and with $(\mathcal{F}^{N}_{t})_{t\geq 1}$ as defined in Section 3.1, we have
-$\mathbb{P}\Big{(}\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}^{N}_{t-1}\Big{)}\geq\varrho^{N}_{\kappa},\quad\forall t\geq 1.$ (61)
-Noting that for all $t\geq T_{\kappa}$ we have
-$1-\mathbb{P}\Big{(}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa|\mathcal{F}_{t-1}^{N}\Big{)}\geq\mathbb{P}\Big{(}\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}_{t-1}^{N}\Big{)}$
-it follows from (61) that
-$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa\Big{)}\leq(1-\varrho^{N}_{\kappa})^{T-T_{\kappa}+1}.$
-To complete the proof of the proposition let $q\in(0,1)$ and note that
-$(1-\varrho^{N}_{\kappa})^{T-T_{\kappa}+1}<1-q,\quad\forall N<\frac{\log\big{(}(1-(1-q)^{\frac{1}{T-T_{\kappa}+1}})^{-1}\big{)}}{\log(1/\varrho_{\kappa})}\leq\frac{\log(T-T_{\kappa}+1)-\log(q)}{\log(1/\varrho_{\kappa})}$
-where the inequality uses the fact that, by Bernoulli’s inequality (see e.g. Li and Yeh,, 2013), we have $(1-q)^{1/(T-T_{\kappa}+1)}\leq 1-q/(T-T_{\kappa}+1)$. The result of the proposition follows. ∎
+5.4.2 Proof of the lemma
+Proof of Lemma 1.
+Let $(\mu_{t})_{t\geq 1}$ and $(\hat{\mu}_{t})_{t\geq 1}$ be as in the statement of the lemma, and let $I=[-\xi,\xi]$ for some $\xi\in(0,\infty)$ such that $\eta_{1}(I)>0$.
+Then, by using Lemma 5, we can easily verify verify that $\iota_{1}:=\inf_{t\geq 1}\hat{\eta}_{t}(I)>0$, which implies that $\iota_{2}:=\inf_{t\geq 2}\eta_{t}(I)\geq\iota_{1}\inf_{x\in I}M(x,I)>0$.
+Consequently, there exists a constant $\iota\in(0,1)$ such that $\inf_{t\geq 1}\eta_{t}(I)\geq\iota$.
+To prove the lemma we let $\tilde{\Delta}_{T}=\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\hat{\mu}_{t-1}M\|_{\mathrm{D}}$ for all $T\geq 1$, with the convention that $\hat{\mu}_{0}M=\eta_{1}$, and we remark first that, for all $t\geq 1$,
+$\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq\iota/2\implies\mu_{t}(I)=\eta_{t}(I)+\mu_{t}(I)-\eta_{t}(I)\geq\eta_{t}(I)-\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\geq\iota/2.$ (55)
+Then, by using (2), (55) and Lemma 6, to prove the lemma it suffices to show that there exists an $\epsilon\in(0,1)$ such that
+$\tilde{\Delta}_{T}\leq\epsilon\implies\sup_{t\in\{1,\dots,T\}}\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq\iota/2,\quad\forall T\geq 1.$ (56)
+To do so we let $C_{I}\in[1,\infty)$ be as in the statement of Lemma 6 (when $I=[-\xi,\xi]$), $\delta\in(0,1)$ be arbitrary, and $C^{\prime}=\max\big\{1,\sup_{x\in(0,1]}x^{1-\delta}\log(1+1/x)\big\}<\infty$.
+With this notation in place, we show by induction on $T\geq 1$ that (56) holds for $\epsilon=(\iota^{2}/(8C_{I}\,C^{\prime}))^{1/\delta}$.
+Since $\epsilon\leq\iota/2$ and $\tilde{\Delta}_{1}=\|\mu_{1}-\eta_{1}\|_{\mathrm{D}}$, the implication in (56) holds when $T=1$, and we now assume that it holds when $T=s$ for some $s\geq 1$.
+Then, by using (55), it follows that $\mu_{t}(I)\geq\iota/2$ for all $t\in\{1,\dots,s\}$ and thus, by Lemma 6 and using (2), we have
+$\sup_{t\in\{1,\dots,s+1\}}\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq 2\sup_{t\in\{1,\dots,s+1\}}\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq\frac{4C_{I}}{\iota}\Delta_{s+1}\log(1+\Delta_{s+1}^{-1})$ (57)
+where
+$\frac{4C_{I}}{\iota}\Delta_{s+1}\log(1+\Delta_{s+1}^{-1})\leq\frac{\iota}{2}\Leftrightarrow\Delta_{s+1}\log(1+\Delta_{s+1}^{-1})\leq\frac{\iota^{2}}{8C_{I}}.$ (58)
+By the definition of the constant $C^{\prime}$, we have $\Delta_{s+1}\log(1+\Delta_{s+1}^{-1})\leq C^{\prime}\,\Delta_{s+1}^{\delta}\leq C^{\prime}\,\tilde{\Delta}_{s+1}^{\delta}$ which, together with (57) and (58), shows that $\tilde{\Delta}_{s+1}\leq\epsilon\implies\sup_{t\in\{1,\dots,s+1\}}\|\mu_{t}-\eta_{t}\|_{\mathrm{D}}\leq\iota/2$, showing that (56) holds for $T=s+1$.
+This concludes the proof of (56) and thus of the lemma.